### v2

# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

Abstract.
Conformally symplectic diffeomorphisms $f:M\rightarrow M$ transform a symplectic form $\omega$ on a manifold $M$ into a multiple of itself, $f^{*}\omega=\eta\omega$. They appear naturally in applications. We assume $\omega$ is bounded, as some of the results in this paper may fail otherwise.
We show that there are deep interactions between the topological properties of the manifold, the dynamical properties of the map, and the geometry of invariant manifolds. When $\eta\neq 1$ the manifold $M$ cannot be closed, as the volume grows or decays under iteration.
We show that, when the symplectic form is not exact, the possible conformal factors $\eta$ are related to topological properties of the manifold. For some manifolds the conformal factors are restricted to be algebraic numbers.
We also find relations between dynamical properties (relations between growth rate of vectors and $\eta$) and symplectic properties (whether $\omega$ vanishes or is non-degenerate on certain subspaces).
Normally hyperbolic invariant manifolds (NHIMs) and their (un)stable manifolds are important landmarks that organize long-term dynamical behaviour. We prove that a NHIM is symplectic if and only if the rates satisfy certain pairing rules and if and only if the rates and the conformal factor satisfy certain (natural) inequalities.
Homoclinic excursions to NHIMs, which are crucial for long-term dynamics – particularly for Arnold diffusion – are quantitatively described by scattering maps. These maps give the trajectory asymptotic in the future as a function of the trajectory asymptotic in the past. We prove that the scattering maps are symplectic even if the dynamics is dissipative. We also show that if the symplectic form is exact, then the scattering maps are exact, even if the dynamics is not exact. We give a variational interpretation of scattering maps in the conformally symplectic setting.
We also show that similar properties of NHIMs and scattering maps hold in the case when $\omega$ is presymplectic. In dynamical systems with many rates (e.g., quasi-integrable systems near multiple resonances), pre-symplectic geometries appear naturally.
Research of M.G. was partially supported by the NSF grant DMS-2307718.
Research of T.M-Seara is supported by the grants PID-2021-122954NB-100 and PID2024-158570NB-I00, both financed by MCIN/AEI/ 10.13039/501100011033/ and by “ERDF A way of making Europe” and by the Spanish State Research Agency, through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020- 001084-M)
1. Introduction
1.1. Overview
We study interactions between topology, geometry, and dynamics in the context of conformally symplectic maps and their invariant manifolds.
Given a symplectic manifold $(M,\omega)$ of dimension $d$, a diffeomorphism on $M$ is conformally symplectic if there exists a conformal factor $\eta>0$ such that
$f^{*}\omega=\eta\omega.$
Conformally symplectic maps and their invariant manifolds present intrinsic mathematical interest, e.g., [AllaisA24, Banyaga02, Lee46, Vaisman75, Vaisman85]. They also arise in applications, such as mechanical systems with friction proportional to velocity, e.g., [CallejaCL22, MargheriOR14, BiascoCh09], as Euler-Lagrange equations of exponentially discounted systems – which appear in financial models and control theory – e.g., [Bensoussan88], or thermostat systems, e.g. [wojtkowski1998conformally]
The goal of this paper is to study the interactions between the topology of the manifold $M$, the dynamical properties of the map $f$, and the geometric structures that organize the dynamics.
The case $\eta=1$ corresponds to symplectic maps. The case $\eta\neq 1$ has some similarities with the symplectic case but also important differences [^1] The limit $\eta\to 1$ is a singular limit, as the properties of the system change dramatically.. Notably, the symplectic volume changes under iteration by a factor, $f^{*}\left(\omega^{d/2}\right)=\eta^{d/2}\left(\omega^{d/2}\right)$. So, the only invariant objects are of volume 0 or $\infty$. The volume 0 case includes interesting objects such as Birkhoff attractors [arnaud2024higherdimensionalbirkhoffattractors], however, in this paper we will concentrate on invariant manifolds of infinite volume. In the infinite-volume setting, uniform boundedness properties of differentiable objects are not straightforward, and they must be carefully formulated and handled.
We will assume that $\|\omega\|$ is bounded (a non-trivial assumption in non-compact manifolds) Indeed, we will show that some of the results in this paper fail for unbounded $\omega$ (see example 5.5).
On the other hand, we do not need that the non-degeneracy properties of $\omega$ are uniform (that is, we can allow that $|\omega^{d/2}|$ is not equivalent to the Riemannian volume). Many of the results of the main Theorems 3.1 and 3.3 (not the pairing rules obtained in Theorem 3.1) work even when $\omega$ is a presymplectic form (see Theorem 3.9).
Many of the remarkable properties of symplectic dynamics extend, often with suitable modifications, to the conformally symplectic setting. Examples include KAM [CallejaCL13, CallejaCL22], and, under convexity assumptions, Aubry-Mather theory [maro2017aubry] and Hamilton-Jacobi theory [Gomes08]. On the other hand, new phenomena such as attractors appear for conformally symplectic systems. In conformally symplectic (but not symplectic) systems it is impossible to have invariant manifolds finite non-zero volume.
A large part of this paper is devoted to the study of Normally Hyperbolic Invariant Manifolds (NHIMs) and their stable/unstable manifolds, which – together with KAM theory – are among the principal sources of invariant objects in symplectic dynamics. We study the properties of NHIMs, their stable and unstable manifolds and their homoclinic intersections, in the conformally symplectic setting.
In this overview, we will present informally the main ideas and results in this paper. Precise formulations will require preliminary definitions undertaken in Section 2.
1.2. Summary of the results
We provide a road map for the results in this paper.
1.2.1. Topology and conformally symplectic dynamics
  • We examine the role of exactness in conformally symplectic dynamics.
    In symplectic geometry, it makes a significant difference when the symplectic form is exact, $\omega=d\alpha$, and the map preserves the action form $\alpha$ up to an exact differential. In conformally symplectic systems the analogue is:
    $f^{*}\alpha=\eta\alpha+dP^{f},$
    for some primitive function $P^{f}$.
    Notice that, even if we assume that $\omega$ is bounded, it is not natural – and often impossible – to assume $\alpha$ is bounded. See Section 9.4. In the conformally symplectic case, it happens often that a map is exact for some action forms but not for others (see Section 5.1).
    This has relations with the de Rham cohomology of the manifold. For two action forms $\alpha,\tilde{\alpha}$ for $\omega$, we have $d(\alpha-\tilde{\alpha})=0$ but not necessarily $\tilde{\alpha}-\alpha=dG$. For any action form $\alpha$, $\alpha+dG$ is an action form too, and if $f$ is exact for $\alpha$, then $f$ is also exact for $\tilde{\alpha}$. We call the addition of $dG$ to an action form a gauge transformation. We study systematically the effect of gauge transformations on primitive functions.
  • In some manifolds, there are relations between the algebraic topology and the conformal factors $\eta$. Consequently, on these manifolds, the possible values of $\eta$ for non-exact $\omega$ are algebraic numbers. See Section 4.2. Using these relations, in Section 4.2, we present an answer to a question raised in [ArnaudF24].
1.2.2. Vanishing lemmas
The interaction between dynamics and geometry arises when we consider rates of growth of vectors under iteration of the map $f$.
The following is a description of the results in Section 6:
  • Since
    $\omega(x)(u,v)=\eta^{-n}\omega(f^{n}(x))(Df^{n}(x)u,Df^{n}(x)v),$
    we see that
    $|\omega(x)(u,v)|\leq\eta^{-n}\|\omega\|\|Df^{n}(x)u\|\|Df^{n}(x)v\|.$
    From this, we can derive vanishing lemmas stating that $\omega$ vanishes on subbundles of tangent vectors whose growth rates satisfy certain relations with respect to $\eta$.
  • Conversely, if the restriction of the form is non-degenerate, the rates in different subbundles must be related in a way that prevents the vanishing of $\omega$.
    These allows us to prove global versions for NHIM of the widely studied pairing rules [Dessler88, dettmann1996proof, wojtkowski1998conformally] for periodic orbits of Lyapunov exponents.
  • One consequence of the vanishing lemmas is that, in systems with many rates (e.g., quasi-integrable systems near multiple resonances), presymplectic geometry naturally arises (i.e., $\omega$ is closed but may be degenerate).
1.2.3. NHIMs for conformally symplectic systems
We recall that a NHIM is an invariant manifold such that there are gaps between the rates of growth of vectors tangent to the manifold and the rates of growth in the stable and unstable bundles that span the normal bundle. See [Fenichel71, Fenichel74, Fenichel77, Pesin, BatesLZ08], and also Appendix A.
NHIMs enjoy regularity properties, are persistent under perturbations and, more importantly for us, they have (un)stable manifolds foliated by strong (un)stable manifolds.
The fact that the invariant manifolds for conformally symplectic maps cannot be compact creates some technical subtleties in the analysis of these objects. The results in this paper depend only on a few regularity properties that we have identified explicitly (H1)-(H4). Some details on the theory of NHIMs that puts these properties in a broader context are in Appendix A.
It should be noted that the unboundedness of manifolds is not merely a technical inconvenience; it can give rise to new geometric phenomena, such as an NHIM folding into itself. See [Eldering12, Example 3.8], reproduced here in Figure 2. To avoid such pathological examples, we have introduced an explicit assumption (U2) that the NHIM has a uniform tubular neighborhood.
  • The main result on NHIM is the following:
    A NHIM is symplectic if and only if either of the two holds:
    • the rates of vectors in the tangent space satisfy some pairing rules and the rates along the stable and unstable manifold also satisfy other pairing rules. See (P).
    • The rates and the conformal factor satisfy certain (natural) inequalities (see (S))
    See Theorem 3.1 and Corollary 3.2.
  • A consequence of the NHIM being symplectic is that the (un)stable manifold is co-isotropic (hence presymplectic), and the kernel of $\omega_{\mid W^{s}_{\Lambda}}$ integrates to give the strong (un)stable foliation.
    This relation between dynamics and presymplectic geometry turns out to be crucial for other subsequent results.
The proof of the results in this paragraph show that the non-degeneracy of the symplectic form on the NHIM implies pairing rules on the rates, and that pairing rules on the rates imply vanishing lemmas for the strong stable and unstable foliations. Interestingly, starting from geometric assumptions, we obtain relations between rates that, in turn, yield new geometric conclusions.
We also obtain results that start with relations between rates and pass through geometric structures to produce new relations between rates. See Section 6.6.
1.2.4. Scattering maps for conformally symplectic systems
It is well known that homoclinic excursions resulting from intersections of stable and unstable manifolds of a NHIM give rise to large-scale motions, symbolic dynamics, and other phenomena. One of our main goals is to understand the symplectic geometry associated to these homoclinic intersections.
An important tool in the study of homoclinic excursions to a NHIM is the scattering map, introduced in [DLS00].
We consider a class of homoclinic excursions given by a transversal intersection of stable/unstable manifolds (the precise conditions are given in Definition 2.22 in Section 2.10). Given a homoclinic excursion $\{f^{n}(x)\}_{n\in\mathbb{Z}}$ to a NHIM $\Lambda$, there are unique points $x_{+},x_{-}\in\Lambda$ such that
$d(f^{n}(x),f^{n}(x_{\pm}))\leq Ce^{-\lambda|n|}\quad\textrm{as}\quad n\to\pm\infty,$
where $\lambda$ is strictly bigger than the rates in the tangent space to the NHIM.
The scattering map $S$ is defined by $S(x_{-})=x_{+}$. As the homoclinic orbit moves along a transversal intersection, the scattering map is defined in an open set. See Figure 3 for a pictorial representation.
The scattering map is a very useful tool to understand long range motions and instability. Most of the applications of the scattering map in the instability problem have been for symplectic dynamics, since it gives a global way to connect invariant objects such as whiskered tori of (possibly) different topology or dimension [DelshamsLS06a, zbMATH06567884]. There are also results using the scattering map in systems with dissipation [Granados17, GideaLM22, AkingbadeGS23].
The scattering map is not itself part of the dynamics; rather, it should be regarded as a comparison between the dynamics restricted to the NHIM and the dynamics along the homoclinic excursion. Nevertheless, a surprising result [GideaLS20a, GideaLS20] shows that iterations of the scattering map interspersed with iterations of the dynamics restricted to the NHIM corresponds to true orbits of the map. The basic results of the previous two papers apply to general systems, including conformally symplectic systems. See also [DelshamsGR12].
Geometric properties of the scattering map in the symplectic case were systematically studied in [DLS08]. In this paper we develop similar results for conformally symplectic systems, but we encounter significant differences.
  • The anchor result in this paper is:
    The scattering map for conformally symplectic maps is symplectic.
    See Theorem 3.3 (A version of this result in the presymplectic case is Theorem 3.9).
    This result is surprising because the dynamics itself could be dissipative; nevertheless, the fact is that the scattering map is a comparison among different dynamics leads to the cancellation of the effects of the dissipation, which are of geometric nature.
    We present seven different proofs of the symplecticity of the scattering map.
    The variety of arguments allows to extend the result to other models that have appeared in the literature. See Appendix D.
  • Another result owed to cancellations is:
    If the symplectic form is exact, $\omega=d\alpha$, then the scattering map is exact.
    Notice that this result does not use that the map is exact. Again, this shows that the scattering map may enjoy properties that the original dynamics does not have.
    We present two different proofs. They depend crucially on the presymplectic nature of $W^{s}_{\Lambda}$.
  • If moreover, the map $f$ is exact, we derive formulas for the primitive function of the scattering map. See Section 9.
    The formulas consist of a finite series plus an integral term as the remainder. Similar formulas were known for symplectic twist maps, and used for numerical computations [Tabacman95].
    The theory behind these formulas, in the case of unbounded manifolds, presents some notable differences from that of bounded manifolds, which are commonly considered in symplectic systems. As we pointed out, there are different action forms corresponding to gauge transformations. The gauges chosen can affect the convergence of the series.
    When either the orbit asymptotic in the future or the orbit asymptotic in the past escape to $\infty$ (i.e., the orbit leaves any compact set after a finite number of steps), we show that there are (rather explicit) gauges where the series converges, and there are also gauges that make the series divergent.
  • We provide a variational interpretation of the primitive function of the scattering map. We remark that, under certain convexity assumptions and provided the Legendre transform can be implemented, the primitive function of the scattering map coincides with the renormalized variational principle used to construct connecting orbits. [Rabinowitz08, Bessi96].
1.3. Organization of the paper
In Section 2, we give careful presentations of several standard concepts. We expect that most readers will be familiar with some versions of these concepts. However, since we cross category boundaries and deal with technicalities such as finite differentiability on unbounded manifolds, certain subtleties must be explicitly addressed. This section could be skipped and used as a reference.
The precise formulation of the results are in Sections 3 and  4.
Section 5 provides several examples that illustrate the phenomena and serve as motivation for the results. It also contain examples which show that some hypotheses are necessary in the main theorems of sections  3 and  4. The examples do not enter into the formulation of results or the proofs.
The proofs of the main results are contained in the following sections. Notably, Section 6 includes statements and proofs of vanishing lemmas which will be a basic tool for other results. The proofs of the main results Theorem 3.1 and Theorem 3.3 are in Section 7 and in Section 8.
Section 9 studies the primitive function of the scattering map and provides formulas for it when the map $f$ is exact conformally symplectic.
The short Section 10, which provides the proof of theorem 3.9, checks that the proofs for some of the results for conformally symplectic maps also apply to presymplectic maps.
In Appendix A, we collect without proofs some of the results on invariant objects for NHIMs. In Appendix  B, we formulate and prove several results on hyperbolic bundles. Two are elementary results, Lemma B.1, B.2, about forward and backward rates. Lemma B.3 is a delicate perturbative result about rates of cocycles. In Appendix C, we present a detailed proof with explicit constants of a result sometimes called the Fiber Contraction Theorem or the Inclination Lemma. Some versions of this result are known in the theory of NHIMs (without explicit constants, or with compactness assumptions). In Appendix D we describe other relevant models that appeared in the recent literature, and sketch how several of the results here can be adapted. We hope that new results can be obtained for these and other models.
2. Preliminaries
In this section, we recall some standard definitions and tools to set the notation. This section can be skipped and referred to as needed. This paper involves connections between different categories, so some precise definitions are needed.
2.1. Some standard notions in differential geometry
2.1.1. Riemannian manifolds, differentiable maps and forms
Throughout this paper, we assume that $M$ is a $d$-dimensional, orientable, connected, $r$ times differentiable Riemannian manifold, with $r\geq r_{0}$ for some $r_{0}$ sufficiently large. We also assume that the Riemannian metric is $r$ times differentiable with the derivatives uniformly bounded (it may be more convenient assuming that the metric is $r+1$ times differentiable, see (U1’)).
Since the goal is dealing with conformally symplectic systems, this will require that the manifold $M$ is unbounded, hence, we will need to make explicit uniformity assumptions for the size and regularity of objects. Later, in Section 2.7, when we consider normally hyperbolic invariant manifolds (NHIMs) contained in $M$, we will need to deal with finite differentiability. (Even if the system is infinitely differentiable or analytic, the NHIM could be finitely differentiable even if it is compact. See Example A.7. The regularity of the NHIM and several other objects associated to it is limited by formulas involving the hyperbolicity rates).
Definition 2.1.
Given an open set $\mathcal{O}\subset M$, we say that map $f:\mathcal{O}\to M$ is of class $\mathcal{C}^{r}$ if it has continuous and uniformly bounded derivatives of order $1,\ldots,r$.
Similarly, we say that a form, or a vector field, or other object is $\mathcal{C}^{r}$ if the object and its derivatives up to order $r$ are continuous and uniformly bounded.
The Definition 2.1 is different from other notions of $C^{r}$-differentiable maps on non-compact manifolds. For example, it is different from differentiability in the sense of the Whitney. In this paper, Definition 2.1 is the only notion of differentiability used.
The sets of $\mathcal{C}^{r}$ vector fields, forms, and other objects with a linear structure, have a $\mathcal{C}^{r}$ norm given by the supremum of the size of the objects and their derivatives, which makes them into a Banach space. Under suitable uniformity assumptions (see (U1)), the set of $\mathcal{C}^{r}$ diffeomorphisms is a Banach manifold [Banyaga97].
2.1.2. Distances and regularity of manifolds
A Riemannian manifold $M$ is said to be bounded if it has finite diameter, i.e., $\sup_{x,y\in M}d(x,y)<\infty$, where $d$ is the Riemannian distance function. As mentioned earlier, in this paper we will consider unbounded manifolds.
The following definitions are standard, but we record them here since there are subtle. See also [PughSW97].
Definition 2.2.
We define the $\mathcal{C}^{0}$ distance among two submanifolds of a common Riemannian manifold, $N_{1},N_{2}\subset M$, as the Hausdorff distance:
$d_{\mathcal{C}^{0}}(N_{1},N_{2})=\sup_{x_{1}\in N_{1}}\inf_{x_{2}\in N_{2}}d(x_{1},x_{2})+\sup_{x_{2}\in N_{2}}\inf_{x_{1}\in N_{1}}d(x_{1},x_{2}).$
For $\mathcal{C}^{1}$ manifolds, we define: $d_{\mathcal{C}^{1}}(N_{1},N_{2})=d_{C^{0}}(TN_{1},TN_{2})$ where $TN$ is the tangent bundle of $N$. We recall that, if $M$ is a Riemannian manifold, we can define a natural metric in $TM$.
Analogously, we define higher differentiable distances by $d_{\mathcal{C}^{r}}(N_{1},N_{2})=d_{C^{r-1}}(TN_{1},TN_{2})$.
We also define the $\mathcal{C}^{r}$ distance among maps as the $\mathcal{C}^{r}$ distance among their graphs. For diffeomorphisms, the distance among diffeomorphisms is the maximum of the distance between the maps themselves and the distance among the inverses.
In this paper, we will need to study some foliations of a manifold by submanifolds. Foliations have regularity characterized by two numbers. One of them is the regularity of the leaves and another one is the dependence of the leaves on the base point along a transversal. We use the definition of [HPS, PughSW97].
Definition 2.3.
We say that a foliation is of class $\mathcal{C}^{\tilde{m},m}$ when the leaves are uniformly $\mathcal{C}^{m}$ manifolds and the dependence of the leaves in the $\mathcal{C}^{m}$ topology on the base point is locally $\mathcal{C}^{\tilde{m}}$.
Some formulations in terms of coordinates appear in (2.18) and (2.19).
2.1.3. Exponentials and connectors
We think of the geodesic flow as a second order equation on the manifold $M$.
Given $x\in M$, $v\in T_{x}M$, the exponential mapping $\exp_{x}(v)$ is defined as the point in the manifold which is the solution at time $1$ of the geodesic flow with initial point $x$ and initial velocity $v$. So, we can think of the exponential mapping at $x$ as a mapping from a neighborhood of $0\in T_{x}M$ to a neighborhood of $x$ in $M$,
It is well known [BurnsG05, Prop 4.5.2] that on a ball $B_{\rho_{x}}(0)\subseteq T_{x}M$ of radius $0<\rho_{x}\ll 1$, the exponential mapping is a diffeomorphism and induces a system of smooth coordinates on the neighborhood $\exp_{x}(B_{\rho_{x}}(0))$ of $x$ in $M$, referred to as geodesic coordinates. For any $x\in M$, the supremum of all such radii $\rho_{x}$ is called the injectivity radius.
If the metric is $\mathcal{C}^{2}$, it follows $\exp_{x}(B_{\rho_{x}/2}(0))$ contains a ball in $M$ centered in $x$ and with radius bounded from below.
If $N\subset M$ is a $\mathcal{C}^{2}$ submanifold of $M$, we can restrict the metric of $M$ to $N$ and use the exponential mapping in $N$ using the geodesic flow in $N$. The geodesic flow for the metric on $M$ with initial velocity in $TN$, in general, does not map to $N$.
The exponential mapping can also be used to identify neighboring tangent spaces. If $\exp_{x}(v)=y$, consider the linear map
(2.1) $S_{x}^{y}:=D\exp_{x}(v):T_{x}M\to T_{y}M.$
If $x,y$ are sufficiently close (depending only on derivatives of the metric) then $S_{x}^{y}$ is invertible, and we think of it as providing an identification of the two tangent spaces. The maps $S_{x}^{y}$ are called connectors in [HirschPPS69], where they are used in applications to hyperbolic systems.
2.1.4. Pullback operator and cells
We recall the following standard definition of pullback.
If $f:M\to M$ is a $\mathcal{C}^{r}$-diffeomorphism and $\beta$ is a $\mathcal{C}^{r}$-differentiable $k$-form on $M$, the pullback of $\beta$ is given by
(2.2) $(f^{*}\beta)(z)(v_{1},\ldots,v_{k})=\beta(f(z))(Df_{z}(v_{1}),\ldots,Df_{z}(v_{k}))$
for $z\in M$ and $v_{1},\ldots,v_{k}\in T_{z}M$.
A $k$-cell $\sigma$ on $M$ is the image of $0,1]^{k}$ through an orientation preserving, $\mathcal{C}^{1}$-embedding $\sigma:[0,1]^{k}\to M$. (We use the same notation for the mapping and its image. ) A choice of orientation on $0,1]^{k}$ induces an orientation on the $k$-cell via $\sigma$.
We will often use an equivalent definition of the pullback of $\beta$ in integral form:
$\int_{\sigma}f^{*}\beta=\int_{f(\sigma)}\beta,\quad\forall\ k-\text{cell}\ \sigma.$
2.2. A standing assumption on the manifold
Throughout the paper we will assume:
(U1) On the manifold $M$, there exists a uniform $\mathcal{C}^{r}$ system of
coordinates.
That is, there exist $\rho>0$ such that for every $x\in M$, there is a $\mathcal{C}^{r}$-coordinate system on the ball $B_{\rho}(x)\subseteq M$. The system of coordinates is uniform in the sense that the coordinate map that takes the balls $B_{\rho}(x)\subset M$ onto the ball $B_{1}(0)\subset{\mathbb{R}}^{d}$, as well as the inverse coordinate maps, have $\mathcal{C}^{r}$ norms that are bounded uniformly.
Remark 2.4.
Assumption (U1), under the previous hypothesis that the metric is uniformly $\mathcal{C}^{r}$ with bounded derivatives, implies that the Riemannian metric on $M$ is complete.
[…]
Using the exponential mapping we can give a concrete construction of the coordinate systems assumed in the standing hypothesis (U1).
The following assumption implies (U1):
(U1’) Assume
  • The metric $g$ on the manifold $M$ is $\mathcal{C}^{r+1}$;
  • The metric $g$ on the manifold $M$ has an injectivity radius bounded from below away from 0.
Remark 2.5.
The fact that (U1’) implies (U1) is standard. Since the injectivity radius is bounded away from zero implies that the size of the neighborhoods covered by the exponential mapping is bounded uniformly away from zero. The assumption on the metric $g$ to be $\mathcal{C}^{r+1}$ is made to obtain that the system of geodesic coordinates is $\mathcal{C}^{r}$.
Hypothesis (U1’) is sufficient but not necessary to obtain (U1). To obtain that the geodesic flow (and the exponential mapping) are $\mathcal{C}^{r}$, one does not need to assume that all the derivatives of $g$ are bounded. It suffices to assume only that the derivatives of the Riemann tensor are bounded. This property of a manifold is referred to as having bounded geometry [Cheeger1982finite, Eldering12].
2.3. Symplectic and presymplectic forms
Definition 2.6.
A $2$-form on a $\mathcal{C}^{r}$-manifold $M$ of even dimension $d$
$\omega:T\,M\times TM\to\mathbb{R}$
is a symplectic form if it is closed, i.e., $d\omega=0$, and non-degenerate, i.e., $\iota_{v}\omega=\omega(v,\cdot)=0\implies v=0$.
The form $\omega$ is exact if $\omega=d\alpha$ for some $1$-form [^2] We refer to $\alpha$ as an action form for $\omega$ following [Haro00], but other names are used in the literature such as symplectic potential, Liouville form, or primitive form. $\alpha$ (not necessarily unique).
Throughout the paper, for a submanifold $N\subset M$ we denote the form $\omega_{\mid N}$ as acting in $TN$, that is:
$\omega_{\mid N}:TN\times TN\to\mathbb{R}.$
Definition 2.7.
A submanifold $L\subset M$ is said to be isotropic if $\omega_{\mid L}=0$, that is, for each $y\in L$, $T_{y}L\subseteq T_{y}L^{\omega}$, where $T_{y}L^{\omega}=\{v\in T_{y}M\,|\,\omega(y)(v,u)=0,\,\forall u\in T_{y}L\}$ is the symplectic orthogonal of $T_{y}L$. A submanifold $L\subset M$ is said to be coisotropic if for each $y\in L$, $T_{y}L^{\omega}\subseteq T_{y}L$. A submanifold $L\subset M$ is Lagrangian if it is both isotropic and coisotropic, or, equivalently, $\omega_{\mid L}=0$ and $\textrm{dim}(L)=\textrm{dim}(M)/2$. (See, e.g., [Weinstein71, Weinstein73].)
We will also pay attention to forms that are degenerate and have constant rank (dimension of the kernel) [Souriau97].
Definition 2.8.
A $2$-form $\omega$ on a $\mathcal{C}^{r}$-manifold $M$ (not necessarily even dimensional) is presymplectic when $d\omega=0$ (but $\omega$ may be degenerate).
The degeneracy of the form can be characterized by its kernel
$K_{x}(\omega)=\{v\in T_{x}M\,|\,\iota_{v}(\omega)=0\}.$
The form $\omega$ is non-degenerate at $x$ iff $K_{x}(\omega)=\{0\}$, in which case the manifold must be even dimensional.
In some works [Souriau97, LibermannM87], the definition of presymplectic forms also includes the constant rank condition that $\dim K_{x}(\omega)=$ const. on open sets. In general, the constant rank is not an open condition, since a small perturbation of the form may decrease the dimension of the kernel.
Remark 2.9.
In this paper, presymplectic forms appear as restrictions of symplectic forms to invariant manifolds with some rates. In this case, the constant rank of the kernel is a consequence of the rate conditions. Since the kernels are related to the rates, they appear naturally and are stable under perturbations. In the opposite direction, the presymplectic structure shows that the distributions of vectors having these rates integrate to a foliation, which is not expected for general distributions.
2.4. Conformally symplectic maps
Definition 2.10.
A $\mathcal{C}^{r}$-map $f:M\to M$ is conformally symplectic if
(2.3) $(f^{*}\omega)(x)=\eta\cdot\omega(x),\textrm{ for all }x\in M,$
for some number $\eta>0$.
The definition of a conformally symplectic map between two different symplectic manifolds is analogous.
We will refer to $\eta$ as the conformal factor.
The condition (2.3) is equivalent to:
[…]
Remark 2.11.
It is well known([LibermannM87, Banyaga02]) that if $\omega$ is a two form of rank greater or equal than $4$, then, if, for a $1$-form $\beta$, we have $\beta\wedge\omega=0$, we conclude $\beta=0$.
A consequence is that, if ${\rm dim}(M)\geq 4$, $\omega$ is symplectic, $f^{*}\omega=\eta\omega$ for some function $\eta$, implies that $\eta$ is a constant.
Applying $d$ to (2.3), $0=f^{*}(d\omega)=d(f^{*}\omega)=d\eta\wedge\omega$. Hence $d\eta=0$.
When $\eta=1$ a system satisfying (2.3) is symplectic, so the results in this paper (that do not include explicitly $\eta\neq 1$) imply the results in [DLS08]. When $\eta\neq 1$, the symplectic volume contracts or expands, and, therefore, invariant manifolds have volume zero or infinite. The interesting case is when the symplectic volume is infinite. In many mechanical applications, the physical friction satisfies $\eta<1$, so we will take this as the default case.
Remark 2.12.
If $\|\omega\|$ is bounded (which we will assume in (U5)), the fact that a manifold $M$ has infinite symplectic volume implies that it is unbounded. Indeed, the symplectic volume and Riemannian volume satisfy $\int_{\sigma}\omega^{d/2}\leq C\|\omega\|^{d/2}\int_{\sigma}d\textrm{Vol}$ for any $d$-dimensional cell $\sigma\subseteq M$, where Vol is the Riemannian volume and $C>0$. Therefore, the fact that the symplectic volume is infinite implies that so is the Riemannian volume. Then the diameter has to be infinite too.
Dealing with unbounded manifolds, we will make explicit assumptions on the uniformity of the objects considered. Particularly, in Section 2.11, we will assume that $\omega$ is uniformly bounded in $\mathcal{C}^{0}$, and in sections 8.2.6 and 8.2.7 we will assume that $\omega$ is uniformly bounded in $\mathcal{C}^{1}$. We will also assume uniform bounds on the derivatives of the map, etc.
2.5. Exact conformally symplectic maps
If the symplectic form is exact, i.e., $\omega=d\alpha$, then (2.3) is equivalent to
$d(f^{*}\alpha-\eta\alpha)=0,$
Therefore,
$f^{*}\alpha=\eta\alpha+\beta\textrm{ for some $1$-form $\beta$ with }d\beta=0.$
In many cases of interest, there are topological reasons why the symplectic form $\omega$ has to be exact; see [ArnaudF24] and Section 4.1.
We say that a map is exact conformally symplectic for the action form $\alpha$ when the form $\beta$ above is not only closed but also exact. That is:
(2.4) $f^{*}\alpha=\eta\alpha+dP^{f}_{\alpha}$
for some function $P^{f}_{\alpha}:M\to\mathbb{R}$, called the primitive function of $f$.
If a conformally symplectic map has a primitive function, it has practical consequences. For instance, under twist conditions, a primitive function leads to generating functions and, therefore, variational principles for orbits. The conformally symplectic systems have variational principles very similar to those of the symplectic systems but involving discounts. These discounted variational principles have a direct interpretation in finance. See Section 9.3.
Lemma 2.13.
Let $(M_{i},\omega_{i}=d\alpha_{i})$, $i=1,2,3$ be exact symplectic manifolds:
Let $g:M_{1}\rightarrow M_{2}$, $f:M_{2}\rightarrow M_{3}$ be exact conformally symplectic with respect to the corresponding forms:
[…]
Then, $f\circ g$ is exact conformally symplectic and the primitive of $f\circ g$ is given by:
(2.5) $P^{f\circ g}=\eta_{f}P^{g}+P^{f}\circ g.$
Proof.
The proof is an straightforward calculation.
(2.6) $\begin{split}(f\circ g)^{*}\alpha_{3}&=g^{*}f^{*}\alpha_{3}=g^{*}(\eta_{f}\alpha_{2}+dP^{f})=\eta_{f}(\eta_{g}\alpha_{1}+dP^{g})+(dg)^{*}P^{f}\\ &=\eta_{f}\cdot\eta_{g}\alpha_{1}+d(\eta_{f}P^{g}+P^{f}\circ g).\end{split}$
The same arguments also show that the inverse of a conformally symplectic exact map $f$ is also exact with primitive:
(2.7) $P^{f^{-1}}=-\frac{1}{\eta_{f}}P^{f}\circ f^{-1}.$
2.5.1. Gauge transformations
The action form $\alpha$ for $\omega$ is non-unique and the notion of exactness may depend on the form $\alpha$ considered. See Example 5.1, which shows that a map may be exact for some $\alpha$ but not for others.
In general, if a map is exact for $\alpha$, it will also be exact under the gauge transformation of the action form  [^3] In electromagnetism, where the electromagnetic field is an exact 2-form exterior derivative of an electromagnetic 1-form (called vector potential), one often uses adding gradient to the vector potential [Thirring, Zangwill]. The term “gauge” was introduced in [Weyl51] for electromagnetic vector potentials
(2.8) $\tilde{\alpha}=\alpha+dG$
for any globally defined function $G$. Gauge transformations are changes to the action form that do not change its cohomology class.
The gauge transformation (2.8) induces a change of the primitive function of $f$. For the form $\tilde{\alpha}$ in (2.8), we have
(2.9) $P^{f}_{\tilde{\alpha}}=P^{f}_{\alpha}+G\circ f-\eta G.$
Remark 2.14.
The theory of exactness and gauge transformations for maps from one manifold to another has some differences from the theory for maps from a manifold to itself. Since the action forms in the domain and the range are different, the theory of maps to different manifolds has more flexibility that the theory of self-maps.
Let $(M_{i},\omega_{i})$, $i=1,2$, be symplectic manifolds and $f:M_{1}\to M_{2}$ be a conformally symplectic map, that is, $f^{*}\omega_{2}=\eta\omega_{1}$.
If $M_{2}$ is exact symplectic with $\omega_{2}=d\alpha_{2}$, then $d(f^{*}\alpha_{2})=\eta\omega_{1}$, so $M_{1}$ is also exact symplectic for the action form $\alpha_{1}=\frac{1}{\eta}f^{*}\alpha_{2}$, and $f$ is exact with primitive $P_{\alpha_{1},\alpha_{2}}^{f}=0$.
The condition that $f$ is exact for $\alpha_{1}$ and $\alpha_{2}$ is that $f^{*}\alpha_{2}=\eta\alpha_{1}+dP^{f}_{\alpha_{1},\alpha_{2}}$.
If we change $\alpha_{i}$ to $\alpha_{i}+dG_{i}$, $i=1,2$, we see that $P^{f}_{\alpha_{1}+dG_{1},\alpha_{2}+dG_{2}}=P^{f}_{\alpha_{1},\alpha_{2}}+G_{2}\circ f-\eta G_{1}$. Hence, we can make any primitive to be zero, by gauge transformations either in the domain or in the range.
In the case that $M_{2}\subset M_{1}$, it is natural to consider that the forms in $M_{2}$ are restrictions of those in $M_{1}$ and that $G_{2}$ is the restriction of $G_{1}$ to $M_{2}$.
As mentioned before, for unbounded manifolds, boundedness properties of maps and forms are important. In most of the results of this paper, we assume $\omega$ is bounded (see (U5)). On the other hand, we will not impose any boundedness assumption on $\alpha$ or on $G$. In some cases, there are lower bounds for any action form in terms of the Riemannian geometry of the manifold. These lower bounds grow to infinity as the point goes to infinity. See Section 9.4, Lemma  9.7. The boundedness of $\omega$ happens in many cases of interest, but boundedness of $\alpha$ may be impossible in some manifolds.
2.6. Presymplectic maps
Presymplectic maps appear naturally when we consider NHIMs with rates that are incompatible with having a symplectic structure (see Part (A) of Theorem 3.1 and Section 6.6).
Definition 2.15.
Let $\omega$ be a presymplectic form (see Definition 2.8) on $M$. We say that a $\mathcal{C}^{r}$-map $f$ is conformally presymplectic if there exists a real valued function $\eta:M\to\mathbb{R}$ and constants $\eta_{\pm}>0$ such that
(2.10) $(f^{*}\omega)(x)=\eta(x)\omega(x),\quad 0<\eta_{-}\leq\eta(x)\leq\eta_{+}<\infty,\textrm{ for all }x\in M.$
While, in general, the conformal factor $\eta(x)$ of a conformally presymplectic map is a function, if the kernel of the presymplectic form has codimension at least 44, the conformal factor has to be a constant (see Proposition 3.4), similarly to the case of conformally symplectic maps (see Remark 2.11).
2.7. Normally hyperbolic invariant manifolds
In this section, we recall the definition of a normally hyperbolic invariant manifold (NHIM).
Definition 2.16.
Let $M$ be a manifold endowed with a smooth Riemannian metric and $f:M\to M$ be a $\mathcal{C}^{1}$-diffeomorphism.
Let $\Lambda\subset M$ be a unbounded, boundaryless, and connected submanifold invariant by $f$.
We say that $\Lambda$ is a NHIM for $f$ if there exists a splitting
(B) $T_{x}M=T_{x}\Lambda\oplus E_{x}^{s}\oplus E_{x}^{u},\textrm{ for all }x\in\Lambda$
that is invariant under $Df$, and, furthermore, there exist rates
(R) $0<\lambda_{\pm}<1,\quad 0\leq\mu_{\pm},\quad\lambda_{+}\mu_{-}<1,\quad\lambda_{-}\mu_{+}<1$
and constants $C_{\pm},D_{\pm}>0$ such that, for all $x\in M$ we have:
[…]
(See Fig. 1.)
[…]
Figure 1. Hyperbolic rates
We denote the dimension of $\Lambda$ by $d_{c}$, and the dimensions of $E_{x}^{s}$, $E_{x}^{u}$ by $d_{s}$, $d_{u}$, respectively, where $d_{c}+d_{s}+d_{u}=d$ with $d=\dim(M)$.
Condition (R) implies that there are gaps between the rates along the stable/unstable bundles and the rates along the NHIM, i.e.,
[…]
As a consequence of the rate conditions (H), by Lemma B.2 we have that
(2.12) $\frac{1}{\mu_{-}}\leq\mu_{+}.$
Since invariant manifolds for conformally symplectic systems are unbounded (see Remark 2.12) we need to make explicit uniformity assumptions on the properties of the objects considered.
We have already made the assumption (U1) (which is implied by (U1’)) that on the manifold $M$, there exists a uniform $\mathcal{C}^{r}$ system of coordinates.
Given a NHIM $\Lambda$, we fix a uniform $\rho$-neighborhood of $\Lambda$ in $M$
(2.13) $\mathcal{O}_{\rho}=\{y\in M\,|\,d(y,\Lambda)<\rho\},\textrm{ for }\rho>0.$
and we assume:
(U2) We assume that the manifold $\Lambda$ has a uniform tubular neigh-
borhood: there is a $\mathcal{C}^{r}$ diffeomorphism, with a $\mathcal{C}^{r}$ inverse,
from a uniform neighborhood of the zero section of $E^{s}\oplus E^{u}$
to the uniform neighborhood $\mathcal{O}_{\rho}$ of $\Lambda$ (see (2.13)).
Remark 2.17.
An assumption that implies assumption (U2) is that the exponential mapping
(2.14) $\mathcal{C}(x,s,u)=\exp_{x}(s+u),\quad x\in\Lambda,s\in E^{s}_{x},u\in E^{u}_{x},$
defines a $\mathcal{C}^{r}$-diffeomorphism from a uniform neighborhood of the zero section of $E^{s}\oplus E^{u}$ to the uniform neighborhood $\mathcal{O}_{\rho}$ of $\Lambda$.
Definition 2.16 says that in a small neighborhood of $\Lambda\subset M$ of any point $x\in\Lambda$, the manifold $M$ is the product of the manifold $\Lambda$ and of the fibers of the bundles $E^{s}$ and $E^{u}$. For compact manifolds, this gives a tubular neighborhood of $\Lambda$. See [HPS].
For unbounded manifolds, it can happen that the manifold $\Lambda$ folds back and comes close to itself. A concrete example appears in [Eldering12, Example 3.8], which we illustrate in Fig. 2. The content of assumption (U2) is that this folding back of the manifold $\Lambda$ does not happen.
Figure 2. A NHIM without a uniform tubular neighborhood. After [Eldering12, Example 3.8]
It is important to note that the assumptions (U1) and (U2) are of a very different nature, namely:
  • (U1) describes a local property of $M$;
  • (U2) describes a global property of $\Lambda$.
In particular, as shown by the previous example, the assumption (U2) does not follow from (U1) or the other assumptions for a NHIM.
We also assume that $f$, $f^{-1}$ are $r$ times differentiable with uniformly bounded derivatives in $\mathcal{O}_{\rho}$:
(U3) $f\in\mathcal{C}^{r}(\mathcal{O}_{\rho})$ and $f^{-1}\in\mathcal{C}^{r}(f(\mathcal{O}_{\rho}))$.
A consequence of (U2) and (U3) is that there exist local stable and unstable manifolds $W^{s,\textrm{loc}}_{\Lambda}$, $W^{u,\textrm{loc}}_{\Lambda}$ in $\mathcal{O}_{\rho}$, as well as stable and unstable fibers $W^{s,\textrm{loc}}_{x}$, $W^{u,\textrm{loc}}_{x}$, for $x\in\Lambda$:
(2.15) $W^{u,s,\textrm{loc}}_{\Lambda}=\bigcup_{x\in\Lambda}W^{s,u,\textrm{loc}}_{x}.$
The union on the right-hand side of (2.15) is a disjoint union. Therefore, the decomposition of $W^{u,s,\textrm{loc}}_{\Lambda}$ into strong stable leaves is a foliation and, moreover, the leaves are identified uniquely by their intersection with $\Lambda$.
These (un)stable and strong (un)stable manifolds and their regularity are described in Appendix A (see Theorem A.1). We will take advantage of the remarkable fact that $W^{u,s,\textrm{loc}}_{\Lambda}$ admit topological characterizations, but has consequences for rates of convergence and regularity.
We will assume that the rates on the manifolds $\lambda_{\pm}$, $\mu_{\pm}$ (see (R)) are such that:
  • (H1)
    The manifold $\Lambda$ is $\mathcal{C}^{1}$.
  • (H2)
    The manifolds $W^{s,\textrm{loc}}_{\Lambda}$, $W^{u,\textrm{loc}}_{\Lambda}$ are $\mathcal{C}^{1}$.
  • (H3)
    The manifolds $W^{s,\textrm{loc}}_{x}$, $W^{u,\textrm{loc}}_{x}$ are $\mathcal{C}^{1}$ uniformly in $x$.
  • (H4)
    The foliations of the (un)stable manifold by strong (un)stable manifolds
    (2.16) $W^{s,\textrm{loc}}_{\Lambda}=\bigcup_{x\in\Lambda}W^{s,\textrm{loc}}_{x},\quad W^{u,\textrm{loc}}_{\Lambda}=\bigcup_{x\in\Lambda}W^{u,\textrm{loc}}_{x},$
    are of type $\mathcal{C}^{1,1}$. (See Definition 2.3.)
The hypotheses (H1), (H2) (H3) are sufficient to use standard differential geometry tools and define forms, etc. The hypothesis (H4) is the natural one to ensure that the wave maps $\Omega_{\pm}$ (projections along the foliation, formally defined in Definition 2.20), are $\mathcal{C}^{1}$.
In some proofs we will require slightly stronger regularity properties, which we will make explicit. These amount to assumptions on the rates.
In this paper, we will assume that
(N) $\mu_{+}\geq 1,\quad\mu_{-}\geq 1$
The assumption (N) is to avoid complicated statements. If $\mu_{+}<1$, there is a unique fixed point $p\in\Lambda$, and $\Lambda$ is a submanifold of the stable manifold of $p$. Many of the regularity statements obtained in the general theory of NHIMs remain true, but they are far from optimal. Also, in some estimates (derived from [DLS08, Proposition 15] or similar) we use bounds like $C_{1}\mu_{+}^{n}+C_{2}\mu_{+}^{2n}\leq C\mu_{+}^{2n}$, for $n\geq 0$. If we do not assume (N), different algebraic expressions for the bounds would be required for $\mu_{+}\geq 1$ and for $\mu_{+}<1$, and therefore many statements would need to distinguish between the two different cases.
In some examples, we do not use (N) but we mention this explicitly.
We will not use the persistence of NHIMs and their dependence on parameters in this paper, but we will use the existence and properties of (un)stable manifolds foliated by strong (un)stable manifolds. Of course, persistence and dependence on parameters of NHIMs is likely to become useful in future work.
2.8. A system of coordinates
In $\mathcal{O}_{\rho}$ (see (2.13)), by taking the system of coordinates assumed in (U1) and restricting it to $W^{s,u,\textrm{loc}}_{\Lambda}$ we can define a new system of coordinates on $W^{s,u,\textrm{loc}}_{\Lambda}$ as below. We use that $W^{s,u,\textrm{loc}}_{\Lambda}$ is foliated by $W^{s,u,\textrm{loc}}_{x}$, and that the foliation is $\mathcal{C}^{1,1}$. In a small enough neighborhood, we can consider the foliation by the $W^{s,u,\textrm{loc}}_{x}$ as a Cartesian product. Any point in $W^{s,u,\textrm{loc}}_{x}$ is given by the coordinate $y$ on the strong (un)stable manifold. Thus, we obtain a new system of coordinates $\varphi$ around any $x\in\Lambda$ such that
(2.17) $\{\varphi(x,y)|\ y\in B_{\tilde{\rho}}(0)\}=W^{s,u,\textrm{loc}}_{x},$
for $B_{\tilde{\rho}}(0)\subseteq E^{s,u}_{x}$. Condition (H4) implies that there exists a constant $C$ so that
(2.18) $\|\partial_{x}\partial_{y}\varphi(x,y)\|\leq C.$
If the foliation is $\mathcal{C}^{\tilde{m},m}$, then there exist a constant $C$ so that for all $0\leq i\leq\tilde{m}$, $0\leq j\leq m$,
(2.19) $\|\partial_{x}^{i}\partial_{y}^{j}\varphi(x,y)\|\leq C.$
A more geometric version of the system of coordinates $\varphi$ is obtained using the exponential mapping. Given $(x,y)$, $x\in\Lambda$, $y\in E^{s,u}_{x}$, we define $(x,y)\mapsto\exp_{x}(y)$, where $\exp$ now denotes the exponential mapping on $W^{s,u,\textrm{loc}}_{x}$.
Remark 2.18.
The system of coordinates $\varphi$ constructed using the exponential mapping on $W^{s}_{x}$ has remarkable regularity properties. Given the smoothness of the $W^{s,\textrm{loc}}_{x}$ the dependence of $\varphi(x,y)$ is basically as smooth as the map $f$. On the other hand, the dependence on $x$ is not as differentiable and is limited by the hyperbolicity rates. See Appendix  A.
Even if the coordinate system $\varphi$ has been constructed without reference to symplectic forms, the vanishing lemmas (e.g. Remark 6.6, Lemma 6.8) show that the coordinate system $\varphi$ enjoys rather remarkable symplectic properties. We anticipate that these properties will be crucial in the proofs of Sections 8.2.3,8.2.4, 8.3.2. It is also a possibility of one step in Section 8.2.7.
2.9. Optimal rates
The way we have formulated the rate conditions in (H), $\mu_{+}$, $\mu_{-}$, $\lambda_{+}$, $\lambda_{-}$ are only bounds on the growth of vectors and can be replaced by other rates. Hence, the only way that one can find relations among them is for the ‘optimal’ bounds, which we denote by $\mu_{+}^{*}$, $\mu_{-}^{*}$, $\lambda_{+}^{*}$, $\lambda_{-}^{*}$, respectively.
More precisely, we define:
(2.20) $\begin{split}\lambda_{+}^{*}=&\inf\{\lambda_{+}\,|\,\|Df^{n}(x)v\|\leq C_{+}\lambda_{+}^{n}\|v\|,\forall n\geq 0,\,\forall x\in\Lambda,\,\forall v\in E^{s}_{x}\},\\ \lambda_{-}^{*}=&\inf\{\lambda_{-}\,|\,\|Df^{n}(x)v\|\leq C_{-}\lambda_{-}^{n}\|v\|,\forall n\leq 0,\,\forall x\in\Lambda,\,\forall v\in E^{u}_{x}\},\\ \mu_{+}^{*}=&\inf\{\mu_{+}\,|\,\|Df^{n}(x)v\|\leq D_{+}\mu_{+}^{n}\|v\|,\forall n\geq 0,\,\forall x\in\Lambda,\,\forall v\in T_{x}\Lambda\},\\ \mu_{-}^{*}=&\inf\{\mu_{-}\,|\,\|Df^{n}(x)v\|\leq D_{-}\mu_{-}^{n}\|v\|,\forall n\leq 0,\,\forall x\in\Lambda,\,\forall v\in T_{x}\Lambda\}.\end{split}$
Notice that, in general, we do not have that
$\exists C_{+}^{*}>0\,\textrm{ s.t. }\|Df^{n}(x)v\|\leq C_{+}^{*}(\lambda_{+}^{*})^{n}\|v\|,\,\forall n\geq 0,\,\forall x\in\Lambda,\,\forall v\in E^{s}_{x},$
but only that
$\forall\varepsilon>0\,\exists C_{+}(\varepsilon)\,\textrm{ s.t. }\|Df^{n}(x)v\|\leq C_{+}(\varepsilon)(\lambda_{+}^{*}+\varepsilon)^{n}\|v\|,\,\forall n\geq 0,\,\forall x\in\Lambda,\,\forall v\in E^{s}_{x}.$
Similar statements hold for the other optimal rates.
Remark 2.19.
From (2.11) and (2.12), we obtain the relations:
(2.21) $\lambda_{+}^{*}\lambda_{-}^{*}<$ $1,$
(2.22) $\mu_{+}^{*}\mu_{-}^{*}\geq$ $1.$
Note that (2.12) (and so (2.22)) is trivial if we assume (N).
2.10. The scattering map
We recall here the scattering map [DLS00, DLS08].
Assume that $\Lambda$ is a NHIM for $f$ and the conditions (H1-H4) are satisfied.
2.10.1. The wave maps
Definition 2.20.
For a point $x\in W_{\Lambda}^{s,\textrm{loc}}$ (resp. $x\in W_{\Lambda}^{u,\textrm{loc}}$), we denote by $x_{+}$ (resp. $x_{-}$) the unique point in $\Lambda$ which satisfies $x\in W_{x_{+}}^{s,\textrm{loc}}$ (resp. $x\in W_{x_{-}}^{u,\textrm{loc}}$).
Consequently the wave maps:
(2.23) $\begin{split}\Omega_{\pm}:W_{\Lambda}^{s,u,\textrm{loc}}&\longrightarrow\Lambda,\\ x&\mapsto x_{\pm},\end{split}$
are well defined.
The standing assumption (H4) implies that the regularity of the wave maps is at least $\mathcal{C}^{1}$. If the foliations of the stable (unstable) manifolds were $\mathcal{C}^{\tilde{m},m}$ as in Theorem A.1, the wave maps would be $\mathcal{C}^{\tilde{m}}$.
From the definition of $\Omega_{\pm}$, it follows that the wave maps satisfy the following equivariance relations:
(2.24) $\begin{split}&\Omega_{\pm}\circ f^{n}_{\mid W^{s,u}_{\Lambda}}=f_{\mid\Lambda}^{n}\circ\Omega_{\pm},\textrm{ for }n\in\mathbb{Z},\end{split}$
where we denote by $f_{\mid N}$ the restriction of the map to a submanifold $N\subseteq M$.
Notice that (2.24) allows to define the wave maps in the global (un)stable manifolds. They will be differentiable as many times as the map $f$ and the foliation by the strong (un)stable maps. In particular, we can define the pullback by the wave maps. Nevertheless, it could happen that the derivatives of $\Omega_{\pm}$ are not uniformly bounded along the global (un)stable manifolds (e.g. if the (un)stable manifold oscillates).
2.10.2. Homoclinic channels
The goal of this section is to define homoclinic intersections between $W^{u,\textrm{loc}}_{\Lambda}$ and $W^{s,\textrm{loc}}_{\Lambda}$ that give rise to a smooth family of homoclinic orbits to $\Lambda$.
We assume that there is a homoclinic manifold $\Gamma\subset W_{\Lambda}^{s}\cap W_{\Lambda}^{u}$ (we require more conditions on $\Gamma$ below).
More concretely, assume there exist $_{-}, N+N_{+}$
$\Gamma\subseteq f^{N_{-}}(W^{u,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho})\cap f^{-N_{+}}(W^{s,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}),$
where $\mathcal{O}_{\rho}$ is defined in (2.13), and, abusing notation, we write:
$W^{s,\textrm{loc}}_{\Lambda}=\bigcup_{0\leq n\leq N_{+}}f^{-n}(W^{s,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho})\textrm{ and }W^{u,\textrm{loc}}_{\Lambda}=\bigcup_{0\leq n\leq N_{-}}f^{n}(W^{u,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}).$
consequently
(2.25) $\Gamma\subseteq W^{u,\textrm{loc}}_{\Lambda}\cap W^{s,\textrm{loc}}_{\Lambda}.$
Since only a finite number of iterates are involved, the regularity of $W^{s,u,\textrm{loc}}_{\Lambda}$ (as well as of its foliation) is the same as that for $W^{s,u,\textrm{loc}}\cap\mathcal{O}_{\rho}$.
Now, we consider some neighborhoods of the stable and unstable manifolds:
(2.26) $\mathcal{O}^{+}_{\rho_{+}}=\mathcal{O}_{\rho_{+}}(W^{s,\textrm{loc}}_{\Lambda}),\,\mathcal{O}^{-}_{\rho_{-}}=\mathcal{O}_{\rho_{-}}(W^{u,\textrm{loc}}_{\Lambda}),$
and let
(2.27) $\mathcal{O}:=\mathcal{O}_{\rho}\cup\mathcal{O}^{+}_{\rho_{+}}\cup\mathcal{O}^{-}_{\rho_{-}},$
We assume that:
(U4) $f\in\mathcal{C}^{r}(\mathcal{O})\textrm{ and }f^{-1}\in\mathcal{C}^{r}(f(\mathcal{O})).$
Observe that hypothesis U4 implies hypothesis (U3), but we keep them separated because some local results require only (U3) while more global ones require U4.
Remark 2.21.
Assumptions (U3) and (U4) are satisfied if we make the simpler assumption that $f$ and $f^{-1}$ are $\mathcal{C}^{r}$ on $M$. However, there are examples, for instance in Celestial Mechanics, where $f$ is unbounded (due to singularities of the vector field) but (U3) and (U4) hold.
2.10.3. Definition of the scattering map
In this section we define the scattering map. By assumption (2.25), the stable and unstable manifolds of $\Lambda$, $W_{\Lambda}^{s,\textrm{loc}}$ and $W_{\Lambda}^{u,\textrm{loc}}$, intersect along the homoclinic manifold $\Gamma$.
We furthermore assume that the intersection between $W_{\Lambda}^{s}$ and $W_{\Lambda}^{u}$ along the homoclinic manifold $\Gamma$ is transversal (see condition (2.28)) and that $\Gamma$ is transversal to the strong (un)stable foliations (2.16) (see condition (2.29)). More concretely:
(2.28) $\begin{split}&\forall\ x\in\Gamma,\textrm{we have:}\\ &T_{x}M=T_{x}W_{\Lambda}^{s}+T_{x}W_{\Lambda}^{u},\\ &T_{x}W_{\Lambda}^{s}\cap T_{x}W_{\Lambda}^{u}=T_{x}\Gamma.\end{split}$
(2.29) $\begin{split}&\forall\ x\in\Gamma,\textrm{we have:}\\ &T_{x}\Gamma\oplus T_{x}W_{x_{+}}^{s,\textrm{loc}}=T_{x}W_{\Lambda}^{s,\textrm{loc}},\\ &T_{x}\Gamma\oplus T_{x}W_{x_{-}}^{u,\textrm{loc}}=T_{x}W_{\Lambda}^{u,\textrm{loc}}.\end{split}$
Given a manifold $\Gamma$ verifying (2.28) and (2.29), we can consider the wave maps $\Omega_{\pm}$ of (2.23) restricted to $\Gamma$.
Under the assumptions (2.28), (2.29) and (H4), we have that $\Gamma$ is $\mathcal{C}^{1}$ and that $\Omega_{\pm}$ are $\mathcal{C}^{1}$ local diffeomorphisms from $\Gamma$ to $\Lambda$.
Definition 2.22.
We say that $\Gamma$ is a homoclinic channel if:
  1. (1)
    $\Gamma\subset W_{\Lambda}^{s,\textrm{loc}}\cap W_{\Lambda}^{u,\textrm{loc}}$ verifies (2.28) and (2.29).
  2. (2)
    The wave map ${\Omega_{-}}_{\mid\Gamma}:\Gamma\to\Omega_{-}(\Gamma)\subset\Lambda$ is a $\mathcal{C}^{1}$-diffeomorphism.
The last hypothesis in Definition  2.22, that ${\Omega_{-}}_{\mid\Gamma}$ is a diffeomorphism from its domain to its range, can always be arranged by restricting $\Gamma$ to a smaller neighborhood where the implicit function theorem applies.
Remark 2.23.
If $\Gamma$ verifies the definition of a homoclinic channel, so do subsets of $\Gamma$. Therefore, there is no loss of generality in considering small enough channels. One can assume without loss of generality that they are bounded.
We denote by $\Omega_{\pm}^{\Gamma}=({\Omega_{\pm}})_{\mid\Gamma}$, and $H_{\pm}^{\Gamma}=\Omega_{\pm}^{\Gamma}(\Gamma)\subset\Lambda$, so that
$\Omega_{\pm}^{\Gamma}:\Gamma\longrightarrow H_{\pm}^{\Gamma}$
are $\mathcal{C}^{1}$-diffeomorphisms.
We define the scattering map associated to $\Gamma$ as follows:
Definition 2.24.
Given a homoclinic channel $\Gamma$ and $\Omega_{\pm}^{\Gamma}:\Gamma\to H_{\pm}^{\Gamma}$ the associated wave maps, we define the scattering map associated to $\Gamma$ to be the $\mathcal{C}^{1}$-diffeomorphism
$S:H_{-}^{\Gamma}\subset\Lambda\to H_{+}^{\Gamma}\subset\Lambda$
given by
(2.30) $S=S^{\Gamma}=\Omega_{+}^{\Gamma}\circ(\Omega_{-}^{\Gamma})^{-1}.$
See Fig. 3
Figure 3. The scattering map.
Since the projections $\Omega_{\pm}$ satisfy (2.24), we have the $\Omega_{\pm}^{\Gamma}$ and $S^{\Gamma}$ satisfy the equivariance relations
(2.31) $\begin{split}\Omega_{\pm}^{\Gamma}=&f_{\mid\Lambda}^{-n}\circ\Omega_{\pm}^{f^{n}(\Gamma)}\circ f^{n},\\ S^{\Gamma}=&f_{\mid\Lambda}^{-n}\circ S^{f^{n}(\Gamma)}\circ f^{n}.\end{split}$
Remark 2.25.
The fact that the scattering map is $\mathcal{C}^{1}$ is a consequence of (H4). If the foliation of the stable (unstable) manifolds are $\mathcal{C}^{\tilde{m},m}$ then the scattering map is $\mathcal{C}^{\tilde{m}}$.
In general, the scattering map depends on $\Gamma$ and is only locally defined. In [DLS00] there are examples where the local domain of the scattering map cannot be extended to a global one (moving along a cycle in $\Lambda$ leads to lack of monodromy).
The scattering map provides an efficient way to quantify the effect of homoclinic trajectories on the NHIM $\Lambda$. In [DLS00, DelshamsLS06a] it is shown that it can be used to study the heteroclinic intersections between invariant objects in $\Lambda$. In [GideaLS20, GideaLS20a] it is shown that iterations of the map restricted to the NHIM $\Lambda$ combined with iterates of the scattering map $S$ are closely followed by true orbits.
In typical situations, we have many scattering maps (due to the existence of multiple intersections of the stable and unstable manifolds). All these scattering maps can be used to generate rich dynamical behavior.
2.11. Uniformity assumptions on the symplectic form
In Section 3 we will assume that $M$ is endowed with a symplectic (presymplectic) form $\omega$. We assume the $\omega$ is $\mathcal{C}^{0}$ on $\mathcal{O}$ (see (2.27)), i.e.:
(U5) $\|\omega_{\mid\mathcal{O}}\|=\sup_{x\in\mathcal{O}}\|\omega(x)\|\leq M_{\omega}<\infty.$
In Section 8.2.6 and Section 8.2.7 we will require the stronger condition that $\omega$ is $\mathcal{C}^{1}$ on $\mathcal{O}$:
(U5) $\|\omega_{\mid\mathcal{O}}\|_{\mathcal{C}^{1}}=\sup_{x\in\mathcal{O}}\|\omega(x)\|+\sup_{x\in\mathcal{O}}\|D\omega(x)\|\leq M_{\omega}<\infty.$
Note that in Section 6.8 we will also consider the case when $\omega$ is unbounded.
3. Main results on NHIMs and scattering maps
In this section we state results on NHIMs. The proofs appear in Sections 7 and 8. Other results which do not involve involve NHIMs appear in Section 4.
3.1. Standing assumptions
Let us summarize the standing assumptions introduced so far:
  • (i)
    $(M,\omega)$ is an orientable, non-compact, connected, symplectic, Riemannian manifold satisfying condition (U1),
  • (ii)
    $f:M\to M$ is a conformally symplectic diffeomorphism of factor $\eta>0$ (see Definition 2.10),
  • (iii)
    $\Lambda$ is a NHIM for $f$ satisfying (B) and the rate conditions (R), (H) and (N), the regularity conditions (H1), (H2), (H3), (H4), and the uniformity condition (U2),
  • (iv)
    $\Gamma$ is a homoclinic channel (see Definition 2.22),
  • (v)
    $f$ satisfies the uniformity conditions (U3) and (U4),
  • (vi)
    The symplectic form $\omega$ satisfies the boundedness condition (U5).
We note that the condition (U4) implies (U3), but some of the results only use (U3). Similarly the condition (H4) implies (H2) and (H3), but some of the results only use (H2) or (H3). This is why we list all of these conditions separately.
3.2. Symplectic properties of NHIMs and pairing rules
The first main result of this paper is:
Theorem 3.1.
Under the standing assumptions from Section 3.1 we have:
(A) Symplecticity of the NHIM:
If the conformal factor $\eta$ and the hyperbolic rates $\lambda_{\pm}$, $\mu_{\pm}$ in (R) satisfy the inequalities
[…]
then the manifold $\Lambda$ is symplectic and $f_{\mid\Lambda}$ is conformally symplectic of conformal factor $\eta$.
(B) Pairing Rules:
The manifold $\Lambda$ is symplectic if and only if the optimal hyperbolicity rates $\mu_{\pm}^{*}$, $\lambda_{\pm}^{*}$ defined in (2.20) satisfy
(P) $\begin{split}\frac{\lambda_{+}^{*}}{\lambda_{-}^{*}}=&\eta,\textrm{ and}\\ \frac{\mu_{+}^{*}}{\mu_{-}^{*}}=&\eta.\end{split}$
The proof of Theorem 3.1 part (A) is given in Section 7.1; the main ingredient is a vanishing lemma (Lemma 6.5). Part (B) is proved in Section 7.2 using another vanishing lemma (Lemma 6.3).
Corollary 3.2.
Under the standing assumptions from Section 3.1 we have:
If $\Lambda$ is symplectic, then it has rates $\lambda_{\pm}$, $\mu_{\pm}$ satisfying (R) that moreover satisfy (S).
Proof.
Since $\Lambda$ is symplectic, by Theorem 3.1 part (B) the optimal rates satisfy the pairing rules (P). By the Definition  2.16 they also satisfy the rate conditions (R). Then we have:
$\begin{split}&\lambda^{*}_{+}<\frac{1}{\mu^{*}_{-}}=\frac{\eta}{\mu^{*}_{+}},\\ &\frac{1}{\lambda^{*}_{-}}>\mu^{*}_{+}={\eta}{\mu^{*}_{-}}.\end{split}$
With algebraic manipulations, this is precisely (S). Now, if we have that $\lambda_{+}^{*}\mu_{+}^{*}\eta^{-1}<1$, there exist $\lambda_{+}>\lambda_{+}^{*}$ and $\mu_{+}>\mu_{+}^{*}$ still satisfying $\lambda_{+}\mu_{+}\eta^{-1}<1$. An analogous reasoning gives the existence of $\lambda_{-}>\lambda_{-}^{*}$ and $\mu_{-}>\mu_{-}^{*}$ still satisfying $\lambda_{-}\mu_{-}\eta<1$. This concludes the proof. ∎
Theorem 3.1 and Corollary 3.2 show that condition (S) is necessary and sufficient for the manifold $\Lambda$ to be symplectic.
Note that in Theorem 3.1, the hypothesis is a condition on the rates S, and the conclusion is another condition on the rates (P). However, to arrive at this conclusion, we must go through the geometry, by showing that $\Lambda$ is symplectic.
In Section 6.6, after developing some tools, we show that some conditions on the rates of an isotropic invariant manifold obstruct normal hyperbolicity.
3.3. Symplectic properties of scattering maps
Theorem 3.3.
Under the standing assumptions from Section 3.1, assume that the conformal factor $\eta$ and the hyperbolic rates $\lambda_{\pm}$, $\mu_{\pm}$ in (R) satisfy the inequalities (S).
Then we have:
(A) Symplecticity of the homoclinic channel:
The manifold $\Gamma$ is symplectic.
(B) Symplecticity of the scattering map:
The wave maps $\Omega_{\pm}:W^{s,u,\mathrm{loc}}_{\Lambda}\to\Lambda$ defined in (2.23) and (2.24) satisfy:
(3.1) $\begin{split}(\Omega_{+})^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\mathrm{loc}}_{\Lambda}},\\ (\Omega_{-})^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{u,\mathrm{loc}}_{\Lambda}}.\end{split}$
As a consequence, since $\Gamma$ is symplectic $\Omega_{\pm}^{\Gamma}=(\Omega_{\pm})_{\mid\Gamma}$ are symplectic maps, and the scattering map $S=S^{\Gamma}=\Omega_{+}^{\Gamma}\circ(\Omega_{-}^{\Gamma})^{-1}$ is symplectic:
$S^{*}(\omega_{\mid\Lambda})=\omega_{\mid\Lambda}.$
(C) Exact symplecticity of the scattering map:
Assume further that the symplectic form is exact $\omega=d\alpha$.
Then,
(3.2) $\begin{split}(\Omega_{+})^{*}(\alpha_{\mid\Lambda})-\alpha_{\mid W^{s,\mathrm{loc}}_{\Lambda}}=dP^{+}_{\alpha},\\ (\Omega_{-})^{*}(\alpha_{\mid\Lambda})-\alpha_{\mid W^{u,\mathrm{loc}}_{\Lambda}}=dP^{-}_{\alpha},\end{split}$
where $P^{\pm}_{\alpha}$ are functions on $W^{s,u,\textrm{loc}}_{\Lambda}$, respectively.
Hence, the scattering map $S$ is exact with respect to $\alpha$, that is
$S^{*}(\alpha_{\mid\Lambda})=\alpha_{\mid\Lambda}+dP_{\alpha}^{S},$
where $P_{\alpha}^{S}$ is a function on $\Lambda$.
Explicit formulas for $P_{\alpha}^{\pm}$ and $P_{\alpha}^{S}$ in the case when $f$ is also exact are provided in Lemma 9.1.
A remarkable aspect of part (C) of Theorem 3.3 is that $\Omega_{\pm}^{\Gamma}$ and $S$ are exact symplectic for all action forms $\alpha$. For conformally symplectic systems, one expects that a map could be exact for some action form but not for others. See Example 5.1.
In Section 2.5.1 we have studied the effect of gauge transformations (changing $\alpha$ into $\alpha+dG$ for some function $G$) on the primitive functions of exact (conformally) symplectic maps. A remarkable result (see (2.9)) is that the primitive of the scattering map is invariant under normalized gauge changes, that is, gauge functions $G$ that vanish on $\Lambda$.
The proof of Theorem 3.3 is given in Section 8.
In Section 8.2 we give seven different proofs of part (B) of Theorem 3.3. Some of them require slightly different hypotheses. For instance, some of the proofs do not use that $\omega$ is closed or non-degenerate, other proofs assume that $\omega$ is $\mathcal{C}^{1}$-bounded, and other ones assume different conditions among the hyperbolic rates and the conformal factor. Thus, they are also applicable in non-symplectic contexts, as in the applications described in the Appendix D.
In Section 8.3 we give two different proofs of part (C) of Theorem 3.3. One is based on Stokes theorem, and the second one on Cartan’s magic formula.
In Section D, we present several problems that have appeared in the literature, for which the methods developed here apply even if the forms playing a role are not symplectic.
To avoid developing complicated language, when $\Omega_{+}^{*}\omega=\omega$, we will say that $\Omega_{+}$ is symplectic even if $\omega$ is not assumed to be closed or non-degenerate.
The variety of proofs shows that the remarkable cancellations leading to the symplectic properties of the scattering maps are at the crossroads of several ideas in symplectic geometry. It seems that this paper has only started to explore the possibilities.
3.4. Results for presymplectic systems
In this section we present some results analogous to Theorem 3.1 and Theorem 3.3 for presymplectic systems. We assume that $\omega$ is a presymplectic form on $M$ (see Definition 2.8) and $f$ is a conformally presymplectic map (see Section 2.6). A motivation for us is to study some NHIMs that appear in quasi-integrable systems near multiple resonances.
Similarly to the symplectic case (see Remark 2.11), under conditions on dimensionality, the conformal presymplectic factor needs to be a constant.
Proposition 3.4.
If for any $x\in M$ we have that $\textrm{codim}(K_{x}(\omega))\geq 4$, then the conformal presymplectic factor $\eta(x)$ is a constant.
Proof.
We will use the following
Lemma 3.5.
Assume $\omega(x)\neq 0$ for any $x\in M$. Then $d\eta=0$ on $K(\omega)$.
Proof.
We have
$0=f^{*}(d\omega)=df^{*}\omega=d(\eta\omega)=d\eta\wedge\omega.$
Then for all $u\in K_{x}(\omega)$, and all $v,w\in T_{x}M$ we have
$0=d\eta(x)(u)\omega(x)(v,w)+\textrm{ other terms with }\omega(x)(u,\cdot).$
Since $u\in K_{x}(\omega)$, the other terms are 0. Since $\omega(x)\neq 0$, there exist $v,w$ such that $\omega(x)(v,w)\neq 0$. Therefore $d\eta(x)(u)=0$. As the result is true for any $x\in M$ the lemma is proved. ∎
Now we proceed with the proof of Proposition 3.4. There exist $A\subset TM$ such that
$TM=K(\omega)\oplus A\textrm{ and }\omega_{\mid A}\textrm{ is non-degenerate.}$
Then using the same argument as in the symplectic case (see Remark 2.11) we obtain that
$\dim A\geq 4\Rightarrow d\eta_{\mid A}=0,$
therefore, using the result of the previous lemma we obtain $d\eta\equiv 0$ and consequently $\eta=\textrm{const.}$
Now we examine the integrability of the kernel of a presymplectic form. For a fixed $x$, the kernel $K_{x}(\omega)$ is a linear subspace of $T_{x}M$, and the family $K_{x}(\omega)$, $x\in M$, determines a distribution on $M$. Rank-11 kernels are just multiples of a vector field and can be integrated by solving ODE’s. For higher rank kernels, the integrability is non-trivial (see [LibermannM87, Souriau97, AlishahL12]).
We say that a presymplectic $\mathcal{C}^{r}$-form $\omega$ has the constant rank property on an open set $\mathcal{U}\subset M$ if the dimension of $K_{x}$, for $x\in\mathcal{U}$, is constant.
Lemma 3.6.
Let $\omega$ be a $\mathcal{C}^{r}$ ($r\geq 1$) presymplectic form with constant rank on $M$.
Then, $K(\omega)$ is an integrable distribution. That is, there exists a foliation $\mathcal{F}$ whose leaves are $\mathcal{C}^{r}$ isotropic manifolds.
The form induced by $\omega$ on the quotient of the manifold $M$ by the foliation $\mathcal{F}$ is a symplectic form.
Proof.
For any $\mathcal{C}^{r}$-vector fields ($r\geq 1$) $u,v,w$, applying the standard formula for the derivative of $\omega$ yields
$\begin{split}0=(d\omega)(u,v,w)&=u\omega(v,w)-v\omega(u,w)+w\omega(u,v)\\ &+\omega([u,v],w)-\omega([v,w],u)+\omega([w,u],v)\end{split}$
If we now assume that $u,v\in K$, we obtain that for any $w$ one of the terms survives. Hence, for any $w$, we have $\omega([u,v],w)=0$, i.e. $u,v]\in K$.
That is, the distribution $K$ is closed under taking commutators. This is the hypothesis of Frobenius theorem. A version of Frobenius theorem with low regularity appears in [Hartman, pp. 123-124]. See also [Yao23].
Applying now Frobenius theorem, we obtain the existence of a foliation $\mathcal{F}$ integrating the distribution given by the kernel $K(\omega)$.
The fact that the form is non-degenerate follows, quotienting by the kernel we obtain a non-degenerate form. ∎
Assume that $f$ is conformally presymplectic. If $u\in K_{x}(\omega)$ then, for any $v\in T_{x}M$, $\omega(x)(u,v)=0$ and, consequently:
$\omega(f(x))(Df(x)u,Df(x)v)=f^{*}\omega(x)(u,v)=\eta\omega(x)(u,v)=0$
therefore $Df(x)u\in K_{f(x)}(\omega)$. That is, a conformally presymplectic map $f$ transforms $K_{x}(\omega)$ into $K_{f(x)}(\omega)$. Consequently, when the rank is constant and the foliation $\mathcal{F}$ exists, the leaves of the foliation $\mathcal{F}$ are preserved by $f$. It is then natural to define an induced map $\tilde{f}$ in the space of leaves.
We can obtain a concrete representation of the space of leaves and the dynamics on it by taking transversal sections $T_{x},T_{f(x)}$ to the foliation $\mathcal{F}$ at $x$ and $f(x)$, respectively. Each transversal can be endowed with the restriction of $\omega$, which is non-degenerate since the transversal excludes the kernel of the form. Note that $\omega_{\mid T_{x}}$ and $\omega_{\mid T_{f(x)}}$ are closed because the exterior derivative commutes with the restriction. Then, given $y\in T_{x}$, associate to it $\tilde{y}=\tilde{f}(y)\in T_{f(x)}$ defined by $\tilde{y}=H\circ f(y)$ where $H$ is the holonomy map sending $x$ to $f(x)$. If $f$ is conformally presymplectic, then $\tilde{f}$ is conformally symplectic from $(T_{x},\omega_{\mid T_{x}})$ to $(T_{f(x)},\omega_{\mid T_{f(x)}})$. Similar constructions appear in [LibermannM87, p. 106 ff.]. We will apply a similar construction to the scattering map.
A useful consequence is the following:
Proposition 3.7.
Let $\Lambda$ be a NHIM for a conformal presymplectic map $f$.
Assume that $\omega_{\mid\Lambda}$ is constant rank presymplectic.
Let $\{\mathcal{F}_{x}\}_{x\in U}$ be the leaves of the foliation in $\Lambda$ integrating $K(\omega_{\mid\Lambda})$.
Then, $\{W^{s}_{\mathcal{F}_{x}}\}_{x\in U}$ is a foliation of $W^{s}_{\Lambda}$ integrating $K(\omega|_{W^{s}_{\Lambda}})$.
Proof of Proposition 3.7.
Since $\omega_{\mid\Lambda}$ is presymplectic, we can use Lemma 3.6 to integrate the kernel $K(\omega_{\mid\Lambda})$, yielding a foliation $\mathcal{F}$ of $\Lambda$. Let $\{\mathcal{F}_{x}\}_{x\in\Lambda}$ be the leaves of the foliation.
We observe that $\mathcal{F}_{x}$ is an isotropic submanifold in $\Lambda$. By applying Proposition 6.9, we obtain that $W^{s}_{\mathcal{F}_{x}}$ is an isotropic submanifold in $W^{s}_{\Lambda}$. It is also a foliation of $W^{s}_{\Lambda}$. ∎
Remark 3.8.
The constant rank property of presymplectic forms is taken as part of the definition in some treatments [Souriau97, LibermannM87]. As we noted earlier, the constant rank assumption is not an open condition, since adding an arbitrary small perturbation to the form may decrease the dimension of the kernel.
In this paper, the kernel of $\omega$ is obtained by applying vanishing lemmas (see Section 6) assuming conditions on the rates. Since the rates in bundles are continuous under small perturbation, we conclude that, in such a case, perturbations do not change the dimension of the kernel. Hence, in such cases, the constant rank assumption is very natural. More concretely, the symplectic form $\omega$ restricted to $W^{s,u}_{\Lambda}$ is presymplectic and has constant rank; see Proposition 6.9. The foliation integrating the kernel of $\omega_{\mid W^{s,u}_{\Lambda}}$ is $\{W^{s,u}_{x}\}_{x\in\Lambda}$. The directions complementary to the kernel integrate to give symplectic manifolds transverse to the leaves (such an example is a homoclinic channel as given in Definition 2.22).
Another example when the constant rank property is implied by the hyperbolic rates is shown in Example 5.4.
3.4.1. The scattering map for conformally presymplectic systems
The main result for conformally presymplectic systems is:
Theorem 3.9.
Assume that $\omega$ is a presymplectic form on $M$, $f$ is a conformally presymplectic map, and the standing assumptions (i), (iii)-(vi) from Section 3.1 hold for $f$ and $\omega$. Assume the conformal factor (2.10) satisfies the rate conditions:
[…]
Then:
(A) Presymplecticity of NHIM and of homoclinic channel:
$\Lambda$ and $\Gamma$ are presymplectic.
(B) Presymplecticity of scattering map:
The wave maps
$\Omega_{\pm}:W^{s,u,\mathrm{loc}}_{\Lambda}\to\Lambda$
preserve the presymplectic form $\omega$ in the sense of (3.1).
As a consequence, the scattering map $S$ associated to $\Gamma$ is presymplectic.
(C) Exact presymplecticity of scattering map:
Assume further that the presymplectic form is exact, i.e., $\omega=d\alpha$. Then, the scattering map $S$ is exact, that is
$S^{*}\alpha=\alpha+dP^{S}.$
(D) Dynamics in the kernel of the presymplectic form:
The foliation $\mathcal{F}$ described in Lemma 3.6 is preserved by both the dynamics $f$ and the scattering map $S$.
When these dynamics are projected onto any transversal section to the foliation, they give conformally symplectic and symplectic maps respectively.
The proof of Theorem 3.9 is given in Section 10.
4. Results on topology of manifolds with conformally symplectic dynamics
In this section, we show that there are interactions between the (co)homology of the manifold and the set of conformally symplectic factors. In particular, we present an answer to a question posed in [ArnaudF24, p. 160].
We do not need many of the analytical assumptions in Section 3.1, particularly those concerning NHIMs.
We assume that we have a well defined cohomology theory and that the $1$ and $2$-cohomology considered are finite dimensional (hence, we can define pull-back operators and they are finite dimensional). Then, we obtain results for maps which are conformally symplectic with respect to forms in this class.
For unbounded manifolds there are several possibilities of cohomology theories and they may give different obstructions to conformal factors.
When we discuss applications to concrete examples, we will make explicit the cohomology theory we are using.
Remark 4.1.
For unbounded manifolds, it is very natural to have infinite dimensional cohomology (for example, an unbounded cylinder with infinitely many handles attached). We do not explore these cases in this paper.
4.1. Topological obstructions to exactness
For a diffeomorphism $f$, we denote by $f^{\#}$ the induced map on cohomology and by $f_{\#}$ the induced map on homology. We will only consider the action on $1$- and 22-(co)homology, and when we need to make explicit the order of the cohomology, we will add a number to the symbol #\#.
We reserve the notation $f^{*}$ for pull-back Denoting by $\beta]$ the cohomology class of a closed form, we have $f^{\#}[\beta]=[f^{*}\beta]$.
Lemma 4.2.
Let $f$ be a conformally symplectic map for a non-exact form $\omega$, and $f^{\#2}:H^{2}(M)\to H^{2}(M)$ be the homomorphism induced by $f$ on the cohomology group of order $2$. Then the conformal factor $\eta$ is an eigenvalue for $f^{\#2}$.
Proof.
Since $f^{*}\omega=\eta\omega$, we have $f^{\#2}[\omega]=\eta[\omega]$. Then the conformal factor $\eta$ is an eigenvalue for $f^{\#2}$ because for a non-exact form, $\omega]\neq 0$. ∎
The result below is a converse of Lemma 4.2.
Lemma 4.3.
Assume $\eta$ is not an eigenvalue of $f^{\#2}$.
Then, the symplectic form $\omega$ is exact, $\omega=d\alpha$, for some $1$-form $\alpha$.
Proof.
Taking $2$-cohomology on the definition of conformally symplectic map (2.3), we obtain:
$f^{\#2}[\omega]-\eta[\omega]=0.$
Hence, if $\eta$ is not an eigenvalue of $f^{\#2}$, then $\omega]=0$. Therefore, there is a $1$-form $\alpha$ so that $\omega=d\alpha$. ∎
Lemma 4.4.
Assume that $\omega=d\alpha$ for some $1$-form $\alpha$.
  • (i)
    If $\eta$ is not an eigenvalue of $f^{\#1}$, then there exists a closed $1$-form $\beta$ so that $\tilde{\alpha}=\alpha+\beta$ satisfies $\omega=d\tilde{\alpha}$ and
    (4.1) $f^{*}{\tilde{\alpha}}-\eta{\tilde{\alpha}}=dP$
    for some primitive function $P$, and therefore $f$ is exact with respect to $\tilde{\alpha}$. The $1$-form $\beta$ is unique up to the addition of an exact form $dG$, for some function $G:M\to\mathbb{R}$. Hence, the function $P$ is unique up to the addition of $G\circ f-\eta G$.
  • (ii)
    If $\eta$ is an eigenvalue of $f^{\#1}$ we have the following dichotomy:
    • (ii.a)
      If $f^{*}\alpha-\eta\alpha]\not\in\textrm{Range}(f^{\#1}-\eta\textrm{Id})$ then there is no closed $1$-form $\beta$ such that $f$ is exact for $\tilde{\alpha}$.
    • (ii.b)
      If $f^{*}\alpha-\eta\alpha]\in\textrm{Range}(f^{\#1}-\eta\textrm{Id})$ then there exists a closed $1$-form $\beta$ such that $f$ is exact for $\tilde{\alpha}$. Moreover, such $\beta$ is unique up to the addition of an exact form $dG$ and of a linear combination of $\beta_{1},\ldots,\beta_{l}$, where $\beta_{i}]$, $i=1,\ldots,l$, are a basis of the space of eigenvectors of $f^{\#1}$ corresponding to $\eta$ and $l$ is the algebraic multiplicity of $\eta$ as an eigenvalue of $f^{\#1}$.
  • (iii)
    If $f$ is symplectic, i.e., $\eta=1$, and $f^{\#1}=\textrm{Id}$, then either $f$ is exact for all actions $\alpha$ or $f$ is not exact for any $\alpha$.
Proof.
Let $\tilde{\alpha}=\alpha+\beta$, with $\beta$ closed $1$-form.
The map $f$ is exact with respect to $\tilde{\alpha}$ if and only if we have (4.1). Taking $1$-cohomology we get
(4.2) $f^{*}{\tilde{\alpha}}-\eta{\tilde{\alpha}}]=[f^{*}{\alpha}-\eta{\alpha}]+f^{\#1}[\beta]-\eta[\beta]=0,$
which is equivalent to
(4.3) $(f^{\#1}-\eta\textrm{Id})[\beta]=-[f^{*}\alpha-\eta\alpha].$
For part (i), as $\eta$ is not an eigenvalue of $f^{\#1}$, we can find a unique solution $\beta]$ for (4.3). This determines $\beta$ up to the addition of the differential of a function $G$. The corresponding change in the primitive function $P$ follows from (2.9).
Part (ii.a) follows from linear algebra.
For part (ii.b), if $\beta]$ and $\beta^{\prime}]$ are two solutions of (4.3), then $\beta]-[\beta^{\prime}]$ is in the eigenspace of $\eta$.
For part (iii), we have that $\eta=1$ is an eigenvalue of $f^{\#1}=\textrm{Id}$ and $\textrm{Range}(f^{\#1}-\eta\textrm{Id})=\{0\}$.
Then, we observe that for any closed form $\beta$,
$f^{\#1}(\alpha+\beta)-(\alpha+\beta)=[f^{\#1}\alpha-\alpha].$
Therefore, the map $f$ is exact for all $\alpha+\beta$ – all possible action forms – if and only if $f$ is exact for $\alpha$. ∎
Remark 4.5.
The assumption in part (iii) that $f^{\#1}$ is the identity is natural for symplectic maps on compact manifolds. This occurs when one considers a cohomology theory which is invariant under homotopy and $f$ lies in the connected component of the identity in the group of symplectic diffeomorphisms. This is the case, for instance, of the time-$1$ map of an (autonomous) Hamiltonian flow.
The dichotomy in (iii) of $f$ being exact for all actions forms or for none, is not true non-symplectic maps (see Example 5.1), or for symplectic maps not in the connected component of identity.
Remark 4.6.
The result in Lemma 4.2 does not depend on the fact that $\omega$ is non-degenerate. Hence, the topological obstruction applies just as well to the conformal factors for non-exact pre-symplectic forms, when the conformal factor is a constant.
Remark 4.7.
In unbounded manifolds there could be several different cohomology theories giving different results. Since we think of Lemma 4.2 as providing obstructions for possible conformal factors $\eta$, the existence of several theories provides different obstructions for each cohomology theory.
Some of these theories may lead to trivial obstructions, and several cohomology theories may lead to the same obstruction. In Example 4.15 we will apply simplicial cohomology to obtain obstructions for $\mathcal{R}$. For this example, we show that these obstructions are optimal (see Proposition 4.11). In Remark 4.18 we will apply Borel-Moore homology obtaining only trivial obstructions. The survey [BanyagaHS24] presents some other cohomology theories useful for (locally) conformally symplectic geometry.
Remark 4.8.
Many (co)-homologies theories are homotopy invariant, i.e., homotopic maps induce maps on (co)-homology that are isomorphic [^4] However, this is not the case for the Borel-Moore cohomology; see Remark 4.18. In such a case, the conformal factor $\eta$ for a non-exact symplectic remains constant under $\mathcal{C}^{1}$-homotopy.
Under assumption (U1), if $g,f$ are $\mathcal{C}^{1}$-close (hence $\mathcal{C}^{1}$ homotopic) conformally symplectic mappings for the same non-exact form, then the conformal factors have to be the same.
Hence, for a non-exact form, the set of conformally symplectic factors is locally constant for $f$ in the $\mathcal{C}^{1}$ topology.
Remark 4.9.
A particular case of Remark 4.8 is that a symplectic diffeomorphism for a non-exact form cannot be perturbed into a conformally symplectic one. The Example 5.2 presents a different argument for the same phenomenon.
In contrast, for exact forms $\omega=d\alpha$, given a function $H$, we can define a vector field $X$ by $i_{X}\omega=dH+\sigma\alpha$.
Assuming that the flow of $X$ exists [^5] This is not obvious in unbounded manifolds since $\alpha$ may be unbounded (see Section 9.4) and the standard Cauchy-Picard theorem does not apply. Indeed, there are examples of action forms with solutions that reach infinity in arbitrarily short times., the time-$t$ map of the flow is conformally symplectic for $\omega$ with a conformal factor $\exp(\sigma t)$.
Hence, if the hypothesis of exactness in Lemma 4.2 does not hold, many of the conclusions fail.
Remark 4.10.
Assume that we consider a cohomology theory that is homotopy invariant. If $f$ is homotopic to the identity, (which is implied by $f$ being the time-$1$ map of a time-dependent flow or by $f$ being $\mathcal{C}^{1}$-close to identity) then the induced maps $f^{\#}$ on $1$ and $2$-cohomology are the identity. If the conformal factor $\eta\neq 1$, then $\eta$ is not an eigenvalue of $f^{\#1}$, $f^{\#2}$ and Lemmas 4.3 and  4.4 (i) apply. In this case, we recover [ArnaudF24, Proposition 9].
Time-11 maps of conformally symplectic vector fields are exact because there are explicit formulas for the primitive function [ArnaudF24].
4.2. Topological obstructions to conformal factors
The composition of two conformally symplectic maps with conformal factors $\eta_{1}$, $\eta_{2}$, is a conformally symplectic map with conformal factor $\eta_{1}\cdot\eta_{2}$. Therefore the set of conformal factors forms a multiplicative subgroup $\mathcal{R}\subseteq{\mathbb{R}}_{+}^{*}$. The topology of the underlying manifold $M$ presents obstructions to the possible $\mathcal{R}$’s (for instance, when $M$ is compact $\mathcal{R}=\{1\}$). Also, conditions on the conformal factor $\eta$ imply conditions on the symplectic structure of $M$.
In this section we will address the following question formulated in [ArnaudF24, p. 160]:
“Can $\mathcal{R}$ be strictly between $\{1\}$ and $\mathbb{R}$$_{+}^{*}$?”
We will not attempt a general setting as in [ArnaudF24], rather we are just presenting some examples in a concrete manifold: the Cartesian product of a torus and an Euclidean space.
For the obstructions we present, one also needs to specify a cohomology theory. The manifold considered is a deformation retract of the torus, and many homology and cohomology theories agree. But there are others, such as Borel-Moore cohomology (see Remark 4.18) that give different answers. We will use the standard De Rham (co)homology.
Note also that the definition of $\mathcal{R}$ requires specifying a set of allowed diffeomorphisms (e.g., specifying the growth properties at $\infty$). We will consider diffeomorphisms with uniformly bounded derivatives and bounded forms.
Proposition 4.11.
There exists a symplectic manifold $(M,\omega)$ such that the set
$\mathcal{R}=\{\eta\in{\mathbb{R}}_{+}\,|\,\exists f:M\rightarrow M\quad\textrm{ s.t. }f^{*}\omega=\eta\omega\}$
satisfies:
  • $\mathcal{R}$ contains a number different from $1$;
  • $\mathcal{R}$ consists only of algebraic numbers of degree d(d1)/2d(d-1)/2 that are products of two algebraic numbers of degree $d$.
Hence for $(M,\omega)$, $\mathcal{R}$ is neither $\{1\}$ nor $\mathbb{R}_{+}$.
In what follows, we describe several constructions that lead to a proof of Proposition 4.11. The symplectic manifold $(M,\omega)$ claimed in this proposition is explicitly constructed in Example 4.15.
We will show, using cohomology, that all elements of $\mathcal{R}$ have to satisfy some conditions. Then, we will provide examples showing that for any number $\eta$ that satisfies the above restrictions, there is a diffeomorphism $f$ that is conformally symplectic for $\omega$ with conformal factor $\eta$.
4.2.1. Computation of $f^{\#2}$ on some manifolds
The matrix elements of $f^{\#2}$ in the standard De Rham cohomology can be computed explicitly when $M=\mathbb{R}^{d}\times\mathbb{T}^{d}$, where $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ and $d\geq 2$. We denote the coordinates on $M$ by $(I,\theta)$. (Later, in Remark 4.18, we will present results for Borel-Moore theory).
We note that $\tilde{f}$, the lift of $f$ to the universal cover has to satisfy for some $A\in GL(d,\mathbb{Z})$,
(4.4) $\tilde{f}(I,\theta+e)=\tilde{f}(I,\theta)+(0,Ae),\quad\forall e\in\mathbb{Z}^{d}.$
Since $A$ is an operator over integers, it is a topological invariant, as small perturbations of $f$ cannot change $A$. We will compute $f^{\#2}$ in terms of $A$.
Let:
$\{\sigma_{ij}\}_{1\leq i<j\leq d}=d\theta_{i}\wedge d\theta_{j}$
be a basis for the $2$-dimensional de Rham cohomology of $M$. The dimension of the $2$-cohomology is d(d1)/2d(d-1)/2.
For $1\leq k\leq l\leq d$, we define the $2$-cell $\gamma_{k,l}(t,s)\subset M$, $(t,s)\in[0,1]^{2}$, by setting $\theta_{k}=t$, $\theta_{l}=s$ (mod $1$), and all the other coordinates in $M$ to zero; that is, $\gamma_{k,l}$ is a $2$-torus embedded in $M$.
We have:
$\int_{\gamma_{kl}}\sigma_{ij}=\delta^{i}_{k}\delta^{j}_{l}$
where $\delta$ is the Kronecker symbol. Therefore, we can compute the matrix elements of $f^{\#2}$ by computing $\int_{\gamma_{kl}}f^{*}\sigma_{ij}$.
The following result will be proved next by a direct calculation as well as a more conceptual argument.
Lemma 4.12.
With the notations in (4.4), we have:
(4.5) $\int_{\gamma_{kl}}f^{*}\sigma_{ij}=-A_{il}A_{jk}+A_{ik}A_{jl}.$
Therefore, in this basis of cohomology, all the coefficients of the matrix of $f^{\#2}$ are integers and the eigenvalues of $f^{\#2}$ are algebraic numbers of degree equal to the dimension of $H^{2}(M)$. In our case, the degree is d(d1)/2d(d-1)/2.
Remark 4.13.
Since $f^{\#2}:H^{2}(M)\rightarrow H^{2}(M)$ is the wedge product of $f^{\#1}:H^{1}(M)\rightarrow H^{1}(M)$ with itself, it follows that the eigenvalues of $f^{\#2}$ are the product of two eigenvalues of $f^{\#1}$, which are algebraic numbers of degree $d$.
In our case, any conformal factor will be the product of two algebraic numbers of degree $d$. This is related to a question raised in [ArnaudF24, p. 165], where a very similar phenomenon is observed in some examples.
As a corollary, we have:
Corollary 4.14.
Let $M=\mathbb{R}^{d}\times{\mathbb{T}}^{d}$.
Assume there is a symplectic form $\omega$ and a conformally symplectic diffeomorphism $f$ with a conformal factor $\eta$ that is not the product of two algebraic numbers of degree $d$
Then, $\omega$ is exact.
Proof of Lemma 4.12.
To prove (4.5) let
$\mathcal{I}\equiv\int_{\gamma_{kl}}f^{*}\sigma_{ij}=\int_{\partial f(\gamma_{kl})}\theta_{i}d\theta_{j}.$
where the second integral is interpreted in the universal cover so that we can use the variable $\theta$.
Introduce the notation:
${\varphi}_{i}(t,s)=f_{\theta_{i}}(\gamma_{kl}(t,s)),$
where the subindex on $f$ indicates the $\theta_{i}$-component of $f(\gamma_{kl}(t,s))$. The periodicity conditions (4.4) give:
(4.6) $\begin{split}&{\varphi}_{i}(1,s)-{\varphi}_{i}(0,s)=A_{ik},\\ &{\varphi}_{i}(t,1)-{\varphi}_{i}(t,0)=A_{il},\\ &\partial_{2}{\varphi}_{i}(1,s)-\partial_{2}{\varphi}_{i}(0,s)=0,\\ &\partial_{1}{\varphi}_{i}(t,1)-\partial_{1}{\varphi}_{i}(t,0)=0,\end{split}$
as well as analogous formulas for $j$ taking the place of $i$.
We have $\partial f(\gamma_{kl})=f(\partial\gamma_{kl})$ consists of four segments and, using (4.6),
$\begin{split}\mathcal{I}=&\int_{0}^{1}{\varphi}_{i}(t,0)\partial_{1}{\varphi}_{j}(t,0)\,dt+\int_{0}^{1}{\varphi}_{i}(1,s)\partial_{2}{\varphi}_{j}(1,s)\,ds\\ &-\int_{0}^{1}{\varphi}_{i}(t,1)\partial_{1}{\varphi}_{j}(t,1)\,dt-\int_{0}^{1}{\varphi}_{i}(0,s)\partial_{2}{\varphi}_{j}(0,s)\,ds\\ =&\int_{0}^{1}{\varphi}_{i}(t,0)\partial_{1}{\varphi}_{j}(t,0)\,dt+\int_{0}^{1}{\varphi}_{i}(1,s)\partial_{2}{\varphi}_{j}(0,s)\,ds\\ &-\int_{0}^{1}{\varphi}_{i}(t,1)\partial_{1}{\varphi}_{j}(t,0)\,dt-\int_{0}^{1}{\varphi}_{i}(0,s)\partial_{2}{\varphi}_{j}(0,s)\,ds\\ =&\int_{0}^{1}[{\varphi}_{i}(t,0)-{\varphi}_{i}(t,1)]\partial_{1}{\varphi}_{j}(t,0)\,dt+\int_{0}^{1}[{\varphi}_{i}(1,s)-{\varphi}_{i}(0,s)]\partial_{2}{\varphi}_{j}(0,s)\,ds\\ =&\int_{0}^{1}-A_{il}\partial_{1}{\varphi}_{j}(t,0)\,dt+\int_{0}^{1}A_{ik}\partial_{2}{\varphi}_{j}(0,s)\,ds\\ =&-A_{il}[{\varphi}_{j}(1,0)-{\varphi}_{j}(0,0)]+A_{ik}[{\varphi}_{j}(0,1)-{\varphi}_{j}(0,0)]\\ =&-A_{il}A_{jk}+A_{ik}A_{jl}.\end{split}$
Now, we present a more conceptual (but more sophisticated) second proof of Lemma 4.12.
Alternative proof of Lemma 4.12.
Observe that we can define a homotopy in the space of differentiable maps connecting $F^{1}(I,\theta)=f(I,\theta)$ and $F^{0}(I,\theta)=(f_{I}(I,\theta),A\theta)$.
Since the action on de Rham cohomology remains constant under a homotopy, we obtain that the action of $F^{1}$ on 2-cohomology is the same as that of $F^{0}$. The latter is just $A^{\wedge 2}$ (the wedge product  [^6] We recall that $A^{\wedge 2}$ is defined by $A^{\wedge 2}(\alpha\wedge\beta)=(A\alpha)\wedge(A\beta)$ for all vectors $\alpha,\beta$ of $A$ with itself), which agrees with the direct calculation.
We work in the lift and we have
$\begin{split}&\tilde{F}^{1}(I,\theta)=(\tilde{f}_{I}(I,\theta),\tilde{f}_{\theta}(I,\theta)),\\ &\tilde{F}^{0}(I,\theta)=(\tilde{f}_{I}(I,\theta),A\theta).\end{split}$
where the subindex under $f$ indicates taking the component.
We have for all $e\in\mathbb{Z}^{d}$,
$\begin{split}&\tilde{f}_{\theta}(I,\theta+e)=\tilde{f}_{\theta}(I,\theta)+Ae,\\ &\tilde{f}_{I}(I,\theta+e)=\tilde{f}_{I}(I,\theta).\end{split}$
We set for $t\in[0,1]$,
$\tilde{F}^{t}(I,\theta)=(\tilde{f}_{I}(I,\theta),\tilde{f}^{t}_{\theta}(I,\theta)),$
with
$\tilde{f}^{t}_{\theta}(I,\theta)=t\tilde{f}_{\theta}(I,\theta)+(1-t)Ae.$
Note that $\tilde{F}^{t}$ remains uniformly differentiable if $\tilde{f}$ is.
The function $\tilde{f}^{t}_{\theta}(I,\theta)$ satisfies
$\tilde{f}^{t}_{\theta}(I,\theta+e)=\tilde{f}^{t}_{\theta}(I,\theta)+Ae.$
Therefore, $\tilde{F}^{t}$ is the lift of a function on the manifold. ∎
4.2.2. A concrete example
In the following example, we construct an explicit conformally symplectic map with a non-exact symplectic form. Note how it is important that the conformal factor has to be chosen to be an eigenvalue of the action in cohomology and, therefore be an algebraic number.
Example 4.15.
Consider $M=\mathbb{R}^{d}\times\mathbb{T}^{d}$.
Let $A\in SL(d,\mathbb{R})$ be such that it has a complete set of eigenvalues/eigenvectors $\lambda_{i},v_{i}$, some of them real and whose product is not $1$ (this is easy to arrange when $d\geq 3$).
For a pair of different real eigenvalues of $A$, $\lambda_{i}$, $\lambda_{j}$ denote:
$\begin{split}&\omega_{0}=\sum_{k}dI_{k}\wedge d\theta_{k},\\ &\omega_{1}=(v_{i}\cdot d\theta)\wedge(v_{j}\cdot d\theta).\end{split}$
Denote $\eta_{A}=\lambda_{i}\lambda_{j}$, which we assume is not $1$. The number $\eta_{A}$ is an eigenvalue of $A^{\wedge 2}$ which is the same as the action of $A$ in $H^{2}(\mathbb{T}^{d})=H^{2}(M)$. The dimension of $H^{2}(\mathbb{T}^{d})$ is $\ell=d(d-1)/2$.
We have the following elementary facts:
  • $A^{*}\omega_{1}=\eta_{A}\omega_{1}$.
  • For $|\varepsilon|\ll 1$, $\omega=\omega_{0}+\varepsilon\omega_{1}$, is not degenerate.
  • For $\varepsilon\neq 0$, $\omega$ is not exact.
We define
(4.7) $f_{A}(I,\theta)=(\eta_{A}A^{-t}I,A\theta).$
where $A^{-t}=(A^{t})^{-1}$ is the inverse of the transpose.
We have that $f_{A}^{*}\omega_{0}=\eta_{A}\omega_{0}$. Similarly, $f^{*}_{A}\omega_{1}=\eta_{A}\omega_{1}$.
Hence
$f^{*}_{A}\omega=\eta_{A}\omega.$
Example 4.15 depends on the choice of an automorphism $A$ of the torus. To emphasize this we denote the $f$ in (4.7) by $f_{\mid A}$. Note that in the definition of $f$, the factor $\eta$ is chosen depending on $A$, so it will be denoted also $\eta_{A}$.
Proof of Proposition 4.11.
Consider $(M,\omega)$ as in Example 4.15, where $\omega=\omega_{0}+\varepsilon\omega_{1}$ with $\varepsilon\neq 0$. Since $\omega$ is not exact, for any conformally symplectic map $f$ on $M$, the symplectic factor $\eta$ is an eigenvalue of $f^{\#2}$.
By Lemma 4.12, $\eta$ must be an algebraic number of degree d(d1)/2d(d-1)/2. Thus $\mathcal{R}\neq\mathbb{R}+^{*}$. For the conformally symplectic map $f_{A}$ from Example 4.15, $\eta_{A}\neq 1$. Thus, $\mathcal{R}$ is strictly between $\{1\}$ and $\mathbb{R}_{+}^{*}$. ∎
Remark 4.16.
If $\lambda_{i}$ is an eigenvalue of $A$ of multiplicity $2$, the construction of Example 4.15 leads to maps that are conformally symplectic with respect to several non-cohomologous symplectic forms.
Remark 4.17.
We will now explore the possibility of constructing more examples by optimizing the choice of $A$.
The main observation is that if $B$ is an automorphism of the torus with simple eigenvalues such that $BA=AB$, then $A$ and $B$ have a common set of eigenvectors [^7] If $Av=\sigma v$, multiplying by $B$ on the left and commuting, we have $A(Bv)=\sigma(Bv)$, Since we are assuming the space of eigenvectors of $A$ with eigenvalue $\sigma$ is $1$-dimensional, we have that there exists $\nu\in\mathbb{R}$ such that $Bv=\nu v$., even though the eigenvalues may be quite different and, indeed, independent over the rationals. Now, if $A,B$ have the same eigenvectors, then the maps $f_{A},f_{B}$ as in (4.7) are conformally symplectic with factors $\eta_{A},\eta_{B}$ respectively. Therefore, the set $\mathcal{R}$ contains the multiplicative group generated by $\eta_{A},\eta_{B}$. This group could well be dense.
The construction of integer matrices that commute and have simple eigenvalues is not obvious. We just point out that several such examples, with many extra properties, appear in [KatoKN11, p. 61 ff.] motivated by the theory of Abelian actions on the torus. For some of these examples, $\mathcal{R}$ is dense in $\mathbb{R}$.
Remark 4.18.
Example 4.15 can be also analyzed using the Borel-Moore (co)homology theory, which is different from the simplicial (co)homology we used so far.
We recall the Borel-Moore homology groups of the Euclidean space are $H^{BM}_{d}(\mathbb{R}^{d})=\mathbb{Z}$ and $H^{BM}_{k}(\mathbb{R}^{d})=\{0\}$ for $k\neq d$. The B-M homology groups of the torus $H^{BM}_{k}(\mathbb{T}^{d})=\mathbb{Z}^{d\choose{k}}$ (which is the same as the simplicial homology). The homology groups of $\mathbb{R}^{d}\times\mathbb{T}^{d}$ can be computed using Künneth formula.
Thus, in Example 4.15 the Borel-Moore is different from the usual simplicial (co)homology, and the obstructions provided by Borel-Moore to $\mathcal{R}$ are all trivial.
This concrete example that we have developed is special, but the main ingredients (finite dimensional cohomology, with some duality – via periods – to homology with integer coefficients) could hold in greater generality, even if they fail in certain manifolds (e.g., in cylinders with infinitely many handles attached, which has an infinite dimensional homology).
Conjecture 4.19.
We expect that for “many” manifolds with finite dimensional (co)homology theories we have:
$\mathcal{R}$ consists of algebraic numbers.
We hope that the precise hypotheses needed could be well known to experts.
5. Examples
In this section, we present several examples of systems that illuminate several of the issues addressed by the theorems. We also note that some of the further results depend on constructing examples.
5.1. An example of a conformally symplectic map that is exact for one action form but not for another
The next example stresses that the exactness property of a conformally symplectic map may depend on the choice of action form.
Example 5.1.
A paradigmatic example of conformally symplectic system is the dissipative standard map with two parameters $\varepsilon,\mu$ acting on $\mathbb{R}\times\mathbb{T}$
(5.1) $f(I,\theta)=(\eta I+\mu+\varepsilon V^{\prime}(\theta),\theta+\eta I+\mu+\varepsilon V^{\prime}(\theta))$
The symplectic form considered in (5.1) is $\omega=dI\wedge d\theta$. This is an exact form and we can take the action form $\tilde{\alpha}_{\sigma}=(I+\sigma)d\theta$ where $\sigma$ is any constant.
We have
$\begin{split}f^{*}&\tilde{\alpha}_{\sigma}=(\eta I+\mu+\varepsilon V^{\prime}(\theta)+\sigma)(d\theta+\eta dI+\varepsilon V^{\prime\prime}(\theta)d\theta)\\ &=\eta Id\theta+\eta^{2}IdI+\eta\varepsilon IV^{\prime\prime}(\theta)d\theta+(\mu+\sigma)d\theta+(\mu+\sigma)\eta dI\\ &+(\mu+\sigma)\varepsilon V^{\prime\prime}(\theta)d\theta+\varepsilon V^{\prime}(\theta)d\theta+\varepsilon\eta V^{\prime}(\theta)dI+\varepsilon^{2}V^{\prime}(\theta)V^{\prime\prime}(\theta)d\theta\\ &=\eta Id\theta+(\mu+\sigma)d\theta\\ &+d\left(\frac{\eta^{2}}{2}I^{2}+\varepsilon\eta IV^{\prime}(\theta)+\eta(\mu+\sigma)I+\varepsilon(\mu+\sigma)V^{\prime}(\theta)+\varepsilon V(\theta)+\varepsilon^{2}\frac{(V^{\prime}(\theta))^{2}}{2}\right).\end{split}$
Note that the term $d\theta$ on the right hand side prevents exactness  [^8] The integral of $d\theta$ over a non-contractible loop in the cylinder is not zero; $\theta$ cannot be made into a continuous variable over the manifold., so the map (5.1) is exact if and only if $\eta\sigma=\mu+\sigma$. Hence, if we choose $\sigma_{*}=\mu/(\eta-1)$, we have that the mapping is exact for the action form $\tilde{\alpha}_{\sigma_{*}}$, and the primitive function is
$P_{\sigma_{*}}=\frac{\eta^{2}}{2}I^{2}+\varepsilon\eta IV^{\prime}(\theta)+\frac{\eta^{2}\mu}{\eta-1}I+\frac{\varepsilon\eta\mu}{\eta-1}V^{\prime}(\theta)+\varepsilon V(\theta)+\varepsilon^{2}\frac{(V^{\prime}(\theta))^{2}}{2}+C.$
In particular, if $\sigma=0$ and the action form is the Liouville form $\tilde{\alpha}_{0}=Id\theta$, the mapping is exact if and only if $\mu=0$.
When $\eta=1$ and $\mu=0$ (5.1) becomes the conservative standard map. The drift parameter $\mu$ is fundamental in applications of the KAM theorem to conformally symplectic systems; one needs to adjust a drift parameter $\mu$ to find an invariant torus of preassigned frequency (see [CallejaCL13]).
When $\eta=1$, $\varepsilon=0$, (5.1) becomes the integrable area preserving twist map whose phase space is foliated by quasi-periodic orbits. For $\eta<1$, $\varepsilon=0$, there is only one quasi-periodic orbit, which illustrates that quasi-periodic orbits may disappear under arbitrarily small dissipation.
Note that the limit $\eta\to 1$ is singular. If we fix $\mu\neq 0$, the $\sigma_{*}(\eta)$ that makes (5.1) exact, goes to $\infty$ as $\eta$ converges to $1$.
An interesting variant of (5.1) (Fröeschlé map) is to take the variables $I,\theta$ to be $n$-dimensional and the drift parameter $\mu$ to be also an $n$-dimensional vector. In such a case, $V^{\prime}$ means the gradient. The calculations presented here show that when $\eta\neq 1$, for any $\mu$ we can find a unique $n$-vector $\sigma_{*}(\mu)$ so that the map is exact for $I\cdot d\theta+\sigma(\mu)\cdot d\theta$.
5.2. Symplectic systems that can not be perturbed into conformally symplectic ones
The second example shows some symplectic maps that cannot be perturbed into conformally symplectic ones. This is due to global properties of the manifold. If this was a model of a mechanical system, it would show that if one adds friction, the friction cannot be just proportional to the velocity due to the global shape of the manifold.
Example 5.2.
Consider the phase space $M={\mathbb{T}}^{2n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{m}$ (with coordinates $(\theta,x,y)$), endowed with the symplectic form $\omega=\sum_{i=1}^{n}d\theta_{i}\wedge d\theta_{i+n}+\sum_{i=1}^{m}dx_{i}\wedge dy_{i}$.
Define the dynamics on $M$ by
$f(\theta,x,y)=(A\theta,\lambda x,\lambda^{-1}y)$
where $A\in Sp(2n,\mathbb{Z})$ is a symplectic matrix whose spectrum is contained between $1/\mu$ and $\mu$, for some $\mu\geq 1$, and $0<\lambda<1$ is a sufficiently small number so that $\lambda\mu<1$ and therefore the set
$\Lambda:=\{(\theta,0,0)\,|\,\theta\in{\mathbb{T}}^{2n}\}\subset M$
is a normally hyperbolic invariant manifold (see Definition 2.16).
The map $f$ is symplectic for $\omega$.
There is no conformally symplectic $\mathcal{C}^{1}$-small perturbation of the map $f$ with a conformal factor different from $1$.
Proof.
If such a perturbation existed, by the theory of NHIM, there should be an invariant manifold $\mathcal{C}^{1}$-close to $\Lambda\sim{\mathbb{T}}^{2n}$ with rates $\lambda_{\pm}$, $\mu_{\pm}$ close to $\lambda$ and $\mu$, and conformally symplectic factor close to $1$. Then, for small enough perturbations, conditions (R) and (S) would be satisfied and, applying Theorem 3.1, the persistent NHIM would be symplectic and the dynamics on it would be conformally symplectic Therefore, the NHIM would have infinite volume.
On the other hand the manifold would be $\mathcal{C}^{1}$-close to $\Lambda\sim{\mathbb{T}}^{2n}$ and hence have finite volume, which contradicts the fact that the dynamics on it is conformally symplectic. ∎
5.3. Minimal set of constraints on rates for the existence of a symplectic NHIM
From (R), (2.22), (P), we see that the minimal set of constraints for the existence of a symplectic NHIM for a conformally symplectic map, in terms of the optimal rates (2.20) and the conformal factor $\eta$, is
(5.2) $\begin{split}\lambda_{+}^{*}\mu_{-}^{*}<1&\textrm{ and }\lambda_{-}^{*}\mu_{+}^{*}<1\\ \mu_{+}^{*}\mu_{-}^{*}\geq&1\\ \frac{\lambda_{+}^{*}}{\lambda_{-}^{*}}=\eta&\textrm{ and }\frac{\mu_{+}^{*}}{\mu_{-}^{*}}=\eta.\end{split}$
The following example shows that there are no other constraints besides (5.2) for the existence of a symplectic NHIM for a conformally symplectic map.
Example 5.3 (Flexibility).
Suppose that $0<\eta\leq 1$ and we are given a set of positive real numbers $\lambda_{+}^{*}$, $\lambda_{-}^{*}$, $\mu_{+}^{*}$ $\mu_{-}^{*}$ satisfying (5.2). Then, there exists a conformally symplectic map $f$, with symplectic factor $\eta$, possessing a symplectic NHIM $\Lambda$, such that the corresponding rates (R) are $\lambda_{+}^{*}$, $\lambda_{-}^{*}$, $\mu_{+}^{*}$ $\mu_{-}^{*}$, respectively.
We denote by $\textrm{Diag}(a_{1},\ldots,a_{k})$ a diagonal matrix of entries $a_{1},\ldots,a_{k}$.
Choose $0<\lambda_{n}<\ldots<\lambda_{1}:=\lambda_{+}^{*}$ and take
$A=\textrm{Diag}(\lambda_{1},\ldots,\lambda_{n}).$
Using (5.2), the condition $\mu_{+}^{*}\mu_{-}^{*}\geq 1$ in (5.2) can be replaced by the equivalent condition
(5.3) $(\mu^{*}_{+})^{2}\geq\eta.$
Choose: $0<\mu_{d}\leq\ldots\leq\mu_{1}:=\mu_{+}^{*}$ subject to the following condition
(5.4) $\mu_{d}\geq\frac{\eta}{\mu_{1}}=\frac{\eta}{\mu^{*}_{+}}.$
Choosing $\mu_{d}$ as in (5.4) is possible since by (5.3) we have $\mu_{1}\geq\frac{\eta}{\mu_{1}}$. Then take
$M=\textrm{Diag}(\mu_{1},\ldots,\mu_{d}).$
Consider the symplectic form on ${\mathbb{R}}^{2n+2d}$:
$\omega=\sum_{i=1}^{n}dy_{i}\wedge dx_{i}+\sum_{i=1}^{d}dv_{i}\wedge du_{i}.$
Define the conformally symplectic map:
(5.5) $\begin{split}f(x,y,u,v)=\left(Ax,\eta A^{-1}y,Mu,\eta M^{-1}v\right).\end{split}$
We have
$\begin{split}\Lambda&=\{(x,y,u,v)\,|\,x=y=0\}\\ E^{s}_{z}&=\{(x,y,u,v)\,|\,y=0,\,u=0,\,v=0\},\ \forall z=(0,0,u_{0},v_{0})\in\Lambda\\ E^{u}_{z}&=\{(x,y,u,v)\,|\,x=0,\,u=0,\,v=0\},\ \forall z=(0,0,u_{0},v_{0})\in\Lambda\\ T_{z}\Lambda&=\{(x,y,u,v)\,|\,x=y=0\},\ \forall z=(0,0,u_{0},v_{0})\in\Lambda.\end{split}$
We now show that $\Lambda$ is a symplectic NHIM with corresponding optimal rates $\lambda_{+}^{*}=\lambda_{1}$, $\lambda_{-}^{*}=\frac{\lambda_{1}}{\eta}$, $\mu_{+}^{*}=\mu_{1}$, $\mu_{-}^{*}=\frac{\mu_{1}}{\eta}$. For $z\in\Lambda$ and $n>0$
$\begin{split}\|Df^{n}(z)w\|&\leq C(\lambda_{1})^{n}\|w\|=C(\lambda_{+}^{*})^{n}\|w\|,\ \text{for}\ w=(x,0,0,0)\in E^{s}_{z},\\ \|Df^{-n}(z)w\|&\leq C(\frac{\lambda_{1}}{\eta})^{n}\|w\|=C(\frac{\lambda_{+}^{*}}{\eta})^{n}\|w\|,\ \text{for}\ w=(0,y,0,0)\in E^{u}_{z}.\end{split}$
Moreover, if $w=(0,0,u,v)\in T_{z}\Lambda$, since $\frac{\eta}{\mu_{d}}\leq\mu_{1}$ by (5.3) we obtain:
$\begin{split}\|Df^{n}(z)w\|&\leq C\left(\max{\left(\mu_{1},\frac{\eta}{\mu_{d}}\right)}\right)^{n}\|w\|=C(\mu_{+}^{*})^{n}\|w\|,\\ \|Df^{-n}(z)w\|&\leq C\left(\max{\left(\frac{1}{\mu_{d}},\frac{\mu_{1}}{\eta}\right)}\right)^{n}\|w\|=C\left(\frac{\mu_{+}^{*}}{\eta}\right)^{n}\|w\|.\end{split}$
5.4. Degenerate forms in a manifold and no paring rules
Now, we present an example illustrating the degeneracy of the forms in invariant manifolds that do not satisfy the pairing rules.
Example 5.4.
Fix the conformal factor $\eta>0$. Consider numbers
$0<a<b<c<d<1<d^{-1}\eta<c^{-1}\eta<b^{-1}\eta<a^{-1}\eta,$
and the map on $\mathbb{R}^{8}$
$f(x_{1},\ldots,x_{4},y_{1},\ldots,y_{4})=(ax_{1},bx_{2},cx_{3},dx_{4},a^{-1}\eta y_{1},b^{-1}\eta y_{2},c^{-1}\eta y_{3},d^{-1}\eta y_{4}).$
The map $f$ is conformally symplectic of factor $\eta$ for the symplectic form
$\omega=dy_{1}\wedge dx_{1}+dy_{2}\wedge dx_{2}+dy_{3}\wedge dx_{3}+dy_{4}\wedge dx_{4}.$
We consider the NHIM given by
$\Lambda_{0}=\{(0,0,0,x_{4},0,0,0,y_{4})\}$
and $\omega_{0}=\omega|_{\Lambda_{0}}=dy_{4}\wedge dx_{4}$ is non-degenerate on $\Lambda_{0}$. We have that
$f_{\mid\Lambda_{0}}(0,0,0,x_{4},0,0,0,y_{4})=(0,0,0,dx_{4},0,0,0,d\eta^{-1}y_{4}).$
The NHIM $\Lambda_{0}$ has $5$-dimensional stable and unstable manifolds given by
$W^{s}_{\Lambda}=\{(x_{1},x_{2},x_{3},x_{4},0,0,0,y_{4})\}\textrm{ and }W^{u}_{\Lambda}=\{(0,0,0,0,x_{4},y_{1},y_{2},y_{3},y_{4})\}.$
The optimal rates are $\lambda_{+}^{*}=c$, $\lambda_{-}^{*}=\eta^{-1}c$, $\mu_{+}^{*}=\eta d^{-1}$, and $\mu_{-}^{*}=d^{-1}$. Note that $\Lambda_{0}$ satisfies the pairing rules.
We also consider the NHIM
$\Lambda=\{(0,x_{2},x_{3},x_{4},0,0,0,y_{4})\}.$
We have that
$f_{\mid\Lambda}(0,x_{2},x_{3},x_{4},0,0,0,y_{4})=(0,bx_{2},cx_{3},dx_{4},0,0,0,d\eta^{-1}y_{4}).$
The manifold $\Lambda$ has $5$-dimensional stable and $7$-dimensional unstable manifolds given by
$W^{s}_{\Lambda}=\{(x_{1},x_{2},x_{3},x_{4},0,0,0,y_{4})\}\textrm{ and }W^{u}_{\Lambda}=\{(0,x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4})\}.$
The corresponding optimal rates are $\lambda_{+}^{*}=a$, $\lambda_{-}^{*}=\eta^{-1}c$, $\mu_{+}^{*}=\eta d^{-1}$, and $\mu_{-}^{*}=b^{-1}$, and they do not satisfy the pairing rules. The form $\omega$ is degenerate on $\Lambda$, as we have $\omega_{\mid\Lambda}=\omega_{0}$. Note that $\Lambda$ is contained in the stable manifold of $\Lambda_{0}$.
Even if $\Lambda$ is not symplectic, we can identify presymplectic forms in $\Lambda$.
Clearly $\Lambda_{0}\subset\Lambda$ and $\Lambda_{0}$ is NHIM for the dynamical system $f_{\mid\Lambda}$.
For every $(x_{4},y_{4})$ we can write the $2$-dimensional leaf:
${\mathcal{L}}_{(x_{4},y_{4})}=\{(0,x_{2},x_{3},x_{4},0,0,0,y_{4})\,|\,x_{2},x_{3}\in\mathbb{R}\}.$
As $(x_{4},y_{4})$ range over $\Lambda_{0}$, the leaves ${\mathcal{L}}_{(x_{4},y_{4})}$ foliate $\Lambda$.
Also $\omega_{\mid{\mathcal{L}}_{(x_{4},y_{4})}}=0$. So that the foliation of $\Lambda$ given by the leaves $\mathcal{L}$} is the foliation integrating the kernel of $\omega_{\mid\Lambda}$ and $\Lambda_{0}$ is the symplectic quotient.
From the dynamical point of view, we can think of ${\mathcal{L}}_{(x_{4},y_{4})}$ as the stable manifold of $(0,0,0,x_{4},0,0,0,y_{4})$ in $f_{\mid\Lambda}$.
However, considered as a dynamical system in $\mathbb{R}^{8}$, ${\mathcal{L}}_{(x_{4},y_{4})}$ is only a weak stable manifold for $(0,0,0,x_{4},0,0,0,y_{4})$. From the point of view of weak stable manifolds, integrability of the foliation is surprising (see [JiangPL1995]).
5.5. Unbounded forms and no pairing rules
In the next example we show that the standing assumption (U5) on the boundedness of the symplectic form is essential for the pairing rules.
Example 5.5.
Let $M={\mathbb{R}}\times\mathbb{T}^{1}\times{\mathbb{R}}^{2}$ be a manifold, and let the (unbounded) symplectic form on $M$ be
$\omega=e^{I}dI\wedge d\theta+dy\wedge dx,\textrm{ for }(I,\theta,x,y)\in{\mathbb{R}}\times\mathbb{T}^{1}\times{\mathbb{R}}^{2}.$
For $t>0$ define the map
$f(I,\theta,y,x)=\left(I+t,\theta,10e^{t}y,x/10\right).$
We have
$f^{*}\omega=e^{t}\omega$
that is, $f$ is conformally symplectic with factor $\eta=e^{t}$.
The set
$\Lambda=\{(I,\theta,0,0)\,|\,(I,\theta)\in{\mathbb{R}}\times\mathbb{T}^{1}\}$
is a NHIM and is symplectic. The optimal rates are:
$\mu_{+}^{*}=\mu_{-}^{*}=1,\,\lambda_{+}^{*}=1/10,\lambda_{-}^{*}=e^{-t}/10.$
The pairing rules (P) do not hold in this example, since $\frac{\mu_{+}^{*}}{\mu_{-}^{*}}=1\neq\eta$.
6. Vanishing lemmas
This section is devoted to formulating and proving vanishing lemmas, which are an important ingredient in the proofs of the main results of Section 3.
We will show that, under assumptions on the rates on convergence of the differential of the map and the conformal factor, several blocks of the symplectic form in the decomposition corresponding to the invariant spaces have to vanish. The idea is to use the invariance equation (2.3) for sufficiently high iterates.
This idea is very general and applies in many other contexts and other types of rates, for example Sacker-Sell spectrum and Lyapunov exponents, and for other geometries such as locally conformal systems.
We call attention to the fact that the proofs of the lemmas in this section do not use that the form $\omega$ is closed, and only Lemma 6.3 uses that the form $\omega$ is non-degenerate in the tangent bundle of the considered submanifold. We also include in this section Proposition 6.9, which gives results about isotropic manifolds. The proof of Proposition 6.9 requires that the form is symplectic and is an easy consequence of Theorems 3.1 and 3.3 whose proof appears later in Section 8.4.
6.1. A basic inequality
Most of the vanishing lemmas in this section rely on the following elementary result:
Lemma 6.1.
Let $f$ be a conformally symplectic diffeomorphism $f:M\to M$ with conformal factor $0<\eta$ as in (2.3). Then for all $x\in M$, $n\in\mathbb{Z}$, and $u,v\in T_{x}M$ we have
(6.1) $\begin{split}|\omega(x)(u,v)|\leq\eta^{-n}\|\omega(f^{n}(x))\|\|Df^{n}(x)u\|\|Df^{n}(x)v\|.\end{split}$
Proof.
Since $f$ is conformally symplectic we have
$\begin{split}\omega(f^{n}(x))(Df^{n}(x)u,Df^{n}(x)v)=\eta^{n}\omega(x)(u,v),\end{split}$
so
(6.2) $\begin{split}\omega(x)(u,v)=\frac{1}{\eta^{n}}\omega(f^{n}(x))(Df^{n}(x)u,Df^{n}(x)v),\end{split}$
which yields (6.1). ∎
6.2. General vanishing lemmas
In this section we will give two general vanishing lemmas for a conformally symplectic diffeomorphism. These lemmas will be the main ingredients of the proof of the pairing rules for a symplectic normally hyperbolic invariant manifold given in Section 7.2.
Lemma 6.2.
Let $f$ be a conformally symplectic diffeomorphism $f:M\to M$ with conformal factor $0<\eta$ as in (2.3). Let $L\subseteq M$ be a submanifold invariant under $f$. Assume that the symplectic form $\omega$ is uniformly bounded in a neighborhood of $L$.
Take $x\in L$, and assume that there exist two constants $C_{1}=C_{1}(x),C_{2}=C_{2}(x)>0$ and two vectors $u,v\in T_{x}L$, such that:
  • (A1)
    There exists $0<\alpha<1$, such that for all $n\geq 0$
    (6.3) $\|Df^{n}(x)u\|\leq C_{1}\alpha^{n}\|u\|,$
  • (A2)
    There exists $0<\beta$ with $\alpha\beta<\eta$ such that there exists an increasing sequence of positive integers $\{n_{j}\}_{j=1,\ldots,\infty}$ such that
    (6.4) $\|Df^{n_{j}}(x)v\|\leq C_{2}\beta^{n_{j}}\|v\|,\textrm{ for }j\geq 0.$
Then $\omega(x)(u,v)=0$.
Proof.
Using that
(6.5) $\begin{split}|\omega(f^{n_{j}}(x))(Df^{n_{j}}(x)u,Df^{n_{j}}(x)v)|\leq&\|\omega_{\mid L}\|C_{1}C_{2}(\alpha\beta)^{n_{j}}\|u\|\|v\|,\end{split}$
(6.1) gives:
$|\omega(x)(u,v)|\leq\|\omega_{\mid L}\|C_{1}C_{2}(\alpha\beta\eta^{-1})^{n_{j}}\|u\|\|v\|$
since $\alpha\beta\eta^{-1}<1$, taking $n_{j}\to\infty$, we obtain $\omega(x)(u,v)=0$. ∎
The next result is a converse of Lemma 6.2.
If we assume that $\omega_{\mid L}$ is non-degenerate, for any $x\in L$, and $0\neq u\in T_{x}L$ we have $\iota_{u}(\omega(x))\neq 0$, i.e., there exists $v\in T_{x}L$ such that $\omega(x)(u,v)\neq 0$. Hence, the hypothesis of the previous lemma have to fail. If for a point $x$ in an invariant manifold with a symplectic form there is a vector decreasing exponentially fast, there has to be another one growing exponentially with a rate that matches. This is the key to pairing rules, but one has to fix some details of uniformity.
Lemma 6.3.
Let $f$ be a conformally symplectic diffeomorphism $f:M\to M$ with conformal factor $0<\eta$ as in (2.3), and $L\subseteq M$ a submanifold invariant under $f$.
Assume that the symplectic form $\omega$ is uniformly bounded in a neighborhood of $L$, and that $\omega_{\mid L}$ is non-degenerate. Take $x\in L$ and assume that for some $0<\alpha<1$, there exists $u\in T_{x}L$ that satisfies (A1) of Lemma 6.2.
Then, for any $\beta$ with $\alpha\beta<\eta$, there exists $v\in T_{x}L$ that fails (A2) of Lemma 6.2
The negation of (A2) is very strong. It means that, for the point $x\in L$ and for the vector $v$, we have that for every $C_{2}>0$ there are only finitely $n_{j}$’s such that $\|Df^{n_{j}}(x)v\|\leq C_{2}\beta^{n_{j}}\|v\|$. Therefore, there exists $n_{0}(C_{2})$ such that
$\|Df^{n}(x)v\|\geq C_{2}\beta^{n}\|v\|\ \ \forall n\geq n_{0}(C_{2}).$
Increasing the constant $C_{2}$, we obtain the previous inequality for all $n\geq 0$.
(6.6) $\|Df^{n}(x)v\|\geq\tilde{C}_{2}(x)\beta^{n}\|v\|\ \ \forall n\geq 0.$
A subtle point is that the $\tilde{C}_{2}$ appearing in (6.6) may depend on the point $x\in L$ even if the assumptions in (A1) hold with uniform constants. The reason is that the failure of (A2) may happen for different sequences depending on the point.
This will be enough for our purposes in Section 7.2 which only need the lower bounds for large enough $n$ and some $x$.
Remark 6.4.
For the experts in Fenichel theory, we point out that the uniformity lemma [Fenichel74] allows to go from (6.6) to bounds with uniform constants. Unfortunately, one of the assumptions of the uniformity lemma is compactness of the manifold, which is not true in our setting.
6.3. Vanishing lemmas on NHIMs
This subsection gives a useful vanishing lemma for a NHIM under assumptions (S) on the rates and the conformal factor, and assuming the form $\omega$ is bounded. It will be crucial to prove that the NHIM is symplectic in Section 7.1.
Lemma 6.5 (Infinitesimal Vanishing Lemma).
We take the standing assumptions from Section 3.1.
Let $x\in\Lambda$. Then, we have:
(6.7) $\begin{split}\mu_{+}\lambda_{+}\eta^{-1}<1&\implies\omega(x)(v_{t},v_{s})=0,\quad\forall v_{t}\in T_{x}\Lambda,\,\forall v_{s}\in E_{x}^{s},\\ \mu_{-}\lambda_{-}\eta<1&\implies\omega(x)(v_{t},v_{u})=0,\quad\forall v_{t}\in T_{x}\Lambda,\,\forall v_{u}\in E_{x}^{u},\\ \lambda_{+}^{2}\eta^{-1}<1&\implies\omega(x)(v^{1}_{s},v^{2}_{s})=0,\quad\,\forall v^{1}_{s},v^{2}_{s}\in E_{x}^{s},\\ \lambda_{-}^{2}\eta<1&\implies\omega(x)(v^{1}_{u},v^{2}_{u})=0,\quad\forall v^{1}_{u},v^{2}_{u}\in E_{x}^{u}.\end{split}$
Proof.
For $v_{t}\in T_{x}\Lambda$, $v_{s}\in E^{s}_{x}$, by(H) we have
(6.8) $\begin{split}\|Df^{n}(x)v_{t}\|\leq&D_{+}\mu_{+}^{n}\|v_{t}\|\textrm{ for }n\geq 0,\\ \|Df^{n}(x)v_{s}\|\leq&C_{+}\lambda_{+}^{n}\|v_{s}\|\textrm{ for }n\geq 0,\end{split}$
so, by (6.1) and (U5),
$|\omega(x)(v_{t},v_{s})|\leq M_{\omega}(C_{+}D_{+})(\lambda_{+}\mu_{+}\eta^{-1})^{n}\|v_{t}\|\|v_{s}\|\textrm{ for }n\geq 0.$
As $\lambda_{+}\mu_{+}\eta^{-1}<1$, and $n$ is arbitrary, we obtain $\omega(x)(v_{t},v_{s})=0$.
Analogously, from
(6.9) $\begin{split}\|Df^{n}(x)v_{t}\|\leq&D_{-}\mu_{-}^{|n|}\|v_{t}\|\textrm{ for }n\leq 0,\\ \|Df^{n}(x)v_{u}\|\leq&C_{-}\lambda_{-}^{|n|}\|v_{u}\|\textrm{ for }n\leq 0,\end{split}$
it follows
$|\omega(x)(v_{t},v_{u})|\leq M_{\omega}(C_{-}D_{-})(\lambda_{-}\mu_{-}\eta)^{|n|}\|v_{t}\|v_{u}\|\|\textrm{ for }n\leq 0$
and, since by assumption, $\lambda_{-}\mu_{-}\eta<1$, we obtain $\omega(x)(v_{t},v_{u})=0$.
Similarly
$|\omega(x)(v^{1}_{s},v^{2}_{s})|\leq\ M_{\omega}(C_{+})^{2}(\lambda_{+}^{2}\eta^{-1})^{n}\|v^{1}_{s}\|\|v^{2}_{s}\|\textrm{ for }n\geq 0,$
and $\lambda_{+}^{2}\eta^{-1}<1$ imply $\omega(x)(v^{1}_{s},v^{2}_{s})=0$.
An analogous argument shows that $\lambda_{-}^{2}\eta<1$ implies $\omega(x)(v^{1}_{u},v^{2}_{u})=0$. ∎
A corollary of Lemma 6.5 is that the manifold $W^{s,u,\textrm{loc}}$ are co-isotropic.
Remark 6.6.
In the neighborhood $\mathcal{O}_{\rho}$ (see (2.13)), it is natural to obtain a system of coordinates in $W^{s,\mathrm{loc}}_{\Lambda}$ to a neighborhood of the zero section of $E^{s}_{\Lambda}$.
We could use any system of coordinates whose coordinate is tangent to the stable bundle. For example the coordinate system in Section 2.8, which has useful geometric properties. In a sufficiently small neighborhood, we can trivialize the stable bundle.
We want to express the symplectic form $\omega$ in the $x,y$ system of coordinates. Note that $(x,0)$ corresponds to points in $\Lambda$.
Assuming (S), the Lemma 6.5 gives us that $\omega(x,0)$ – the form $\omega$ in the manifold $\Lambda$, has the representation
$\omega(x,0)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}$.
Therefore, if $\omega$ is differentiable and satisfies (U5) we have, by the mean value theorem:
(6.10) $\omega(x,s)=\begin{pmatrix}\omega_{xx}(x)&0\\ 0&0\end{pmatrix}+e(x,y),\textrm{ with }\|e(x,y)\|\leq CM^{\prime}_{\omega}\|y\|.$
We will use these estimates in Section 8.2.7.
Remark 6.7.
As a consequence of Lemma 6.5, we obtain that the stable and unstable bundles have equal dimension, i.e., $d_{s}=d_{u}$. The reason is that relative to $T_{\Lambda}M\oplus E^{s}\oplus E^{u}$ we have
$\omega=\begin{pmatrix}\omega_{xx}&0&0\\ 0&0&\omega_{y_{s}y_{u}}\\ 0&\omega_{y_{u}y_{s}}&0\\ \end{pmatrix}$
and $\omega_{y_{s}y_{u}}$, $\omega_{y_{u}y_{s}}$ are non-degenerate. Therefore, $E^{s}$ is isomorphic to $E^{u}$.
6.4. Vanishing lemmas on stable/unstable manifolds
The following lemma can be considered as an analogue of Lemma 6.5 for the stable and unstable manifolds of a NHIM. It will be used in the proof of part (B) of Theorem  3.3.
Lemma 6.8 (Vanishing Lemma).
We adopt the standing assumptions from Section 3.1.
Let $y\in W^{s,\mathrm{loc}}_{x}$ for some $x\in\Lambda$. Then, we have:
(6.11) $\begin{split}\mu_{+}\lambda_{+}\eta^{-1}<1&\implies\omega(y)(v_{t},v_{s})=0,\quad\forall v_{t}\in T_{y}W^{s,\mathrm{loc}}_{\Lambda},\,\forall v_{s}\in T_{y}W_{x}^{s,\mathrm{loc}},\\ \lambda_{+}^{2}\eta^{-1}<1&\implies\omega(y)(v^{1}_{s},v^{2}_{s})=0,\quad\,\forall v^{1}_{s},v^{2}_{s}\in T_{y}W^{s,\mathrm{loc}}_{x}.\end{split}$
Analogously, let $y\in W^{u,\mathrm{loc}}_{x}$ for some $x\in\Lambda$. Then, we have:
(6.12) $\begin{split}\mu_{-}\lambda_{-}\eta<1&\implies\omega(y)(v_{t},v_{u})=0,\quad\forall v_{t}\in T_{y}W^{u,\mathrm{loc}}_{\Lambda},\,\forall v_{u}\in T_{y}W^{u,\mathrm{loc}}_{x},\\ \lambda_{-}^{2}\eta<1&\implies\omega(x)(v^{1}_{u},v^{2}_{u})=0,\quad\forall v^{1}_{u},v^{2}_{u}\in T_{y}W^{u,\mathrm{loc}}_{x}.\end{split}$
Consequently:
(6.13) $\begin{split}\lambda_{+}^{2}\eta^{-1}<1&\implies W^{s,\mathrm{loc}}_{x}\textrm{ is isotropic},\ \forall x\in\Lambda,\\ \lambda_{-}^{2}\eta<1&\implies W^{u,\mathrm{loc}}_{x}\textrm{ is isotropic},\ \forall x\in\Lambda.\end{split}$
Note that in (6.11), we have $\lambda_{+}\mu_{+}\eta^{-1}<1$ implies $\lambda_{+}^{2}\eta^{-1}<1$, since $\lambda_{+}<\mu_{+}$.
The conclusions in the first lines of (6.11), (6.12) can be stated geometrically as saying that, for all $x\in\Lambda$:
(6.14) $\begin{split}&\forall y\in W^{s}_{x},v_{s}\in T_{y}W^{s}_{x}\implies i_{v_{s}}(\omega_{\mid W^{s}_{\Lambda}})=0,\\ &\forall y\in W^{u}_{x},v_{u}\in T_{y}W^{u}_{x}\implies i_{v_{u}}(\omega_{\mid W^{u}_{\Lambda}})=0.\end{split}$
In other words, $W^{s,u}_{\Lambda}$ is presymplectic and the foliation given by the kernel is the foliation of strong stable/unstable manifolds.
Proof of Lemma 6.8.
We consider the case of the stable manifold. For points in the unstable manifold we proceed analogously. Take $y\in W^{s,\textrm{loc}}_{\Lambda}$ and $x\in\Lambda$ such that $y\in W^{u,\mathrm{loc}}_{x}$.
For $v_{t}\in T_{y}W^{s,\mathrm{loc}}_{\Lambda}$, $v_{s}\in T_{y}W^{s,\mathrm{loc}}_{x}$ we have, using the bounds (B.2) in Lemma B.3, we have that there exist $D_{+},C_{+}$, such that:
(6.15) $\begin{split}\|Df^{n}(y)v_{t}\|\leq&D_{+}\mu_{+}^{n}\|v_{t}\|\textrm{ for }n\geq N,\\ \|Df^{n}(y)v_{s}\|\leq&C_{+}\lambda_{+}^{n}\|v_{s}\|\textrm{ for }n\geq N.\end{split}$
Using (6.1) we obtain:
$|\omega(y)(v_{t},v_{s})|\leq M_{\omega}(C_{+}D_{+})(\lambda_{+}\mu_{+}\eta^{-1})^{n}\|v_{t}\|\|v_{s}\|\textrm{ for }n\geq N.$
Since $\lambda_{+}\mu_{+}\eta^{-1}<1$ by assumption, and $n\geq N$ is arbitrary, we obtain $\omega(y)(v_{t},v_{s})=0$.
Analogously, for $v^{1}_{s},v^{2}_{s}\in T_{y}W^{s,\mathrm{loc}}_{x}$:
$|\omega(y)(v^{1}_{s},v^{2}_{s})|\leq C_{+}^{2}M_{\omega}(\lambda_{+}^{2}\eta^{-1})^{n}\|v^{1}_{s}\|\|v^{2}_{s}\|,\ \forall n\geq N,$
and therefore, if $\lambda_{+}^{2}\eta^{-1}<1$, we have $\omega(y)(v^{1}_{s},v^{2}_{s})=0$ for any $v^{1}_{s},v^{2}_{s}\in T_{y}{W^{s,\mathrm{loc}}_{x}}$. As this is true for any $y\in W^{s,\mathrm{loc}}_{x}$, we obtain that $\omega_{\mid W^{s,\mathrm{loc}}_{x}}=0$, and therefore $W^{s,\mathrm{loc}}_{x}$ is isotropic.
6.5. Results on isotropic and coisotropic manifolds
The next Proposition 6.9 gives results which ensure that the form $\omega$ vanishes on some manifolds. This result will not be used in the proofs of Theorems 3.1 and 3.3. The proof of Proposition 6.9 is given in Section 8.4 after the proofs of these theorems.
Proposition 6.9 ((Co)isotropic submanifolds).
We take the standing assumptions from Section 3.1.
  • (i)
    If $N\subset\Lambda$ is an isotropic submanifold (not necessarily invariant), that is, $\omega_{\mid N}=0$, then we have:
    (6.16) $\begin{split}\mu_{+}\lambda_{+}\eta^{-1}<1&\implies W^{s}_{N}\textrm{ is isotropic},\ \text{that is}\ \omega_{\mid W^{s}_{N}}=0,\\ \mu_{-}\lambda_{-}\eta<1&\implies W^{u}_{N}\textrm{ is isotropic},\ \text{that is}\ \omega_{\mid W^{u}_{N}}=0.\end{split}$
  • (ii)
    The stable and unstable manifolds of $\Lambda$ satisfy:
    (6.17) $\begin{split}\mu_{+}\lambda_{+}\eta^{-1}<1&\implies W^{s}_{\Lambda}\textrm{ is coisotropic},\\ \mu_{-}\lambda_{-}\eta<1&\implies W^{u}_{\Lambda}\textrm{ is coisotropic}.\end{split}$
We note that $\omega_{\mid W^{s,u}_{\Lambda}}$ is presymplectic and its kernel $K_{x}(\omega)$ has constant rank equal to $d_{u}=d_{s}$; see Remark 3.8. Recall that $d_{s}=d_{u}$ from Remark 6.7.
6.6. Some properties of rates in isotropic invariant manifolds
In this section we start to explore the interaction between rates and isotropic invariant manifolds.
In this section we assume that $\Lambda$ is an invariant manifold satisfying conditions (B) and (H), but not necessarily R. In particular, $\Lambda$ is not necessarily normally hyperbolic. It is easy to see that Theorem 3.1 (A) still holds under these hypotheses [^9] Its proof uses Lemma 6.5 which only requires (B), (H) and (S). A consequence of Corollary 3.2 is that if $\Lambda$ is not symplectic, then it does not satisfy (S). One interesting case is when $\Lambda$ is an isotropic manifold. By the vanishing lemmas, there are assumptions on rates that imply that $\Lambda$ is isotropic. Hence, we have some inequalities on rates (involving only $\mu_{\pm},\eta$) that imply other inequalities on rates (involving $\mu_{\pm},\lambda_{\pm},\eta$) by passing through isotropic manifolds.
This reveals some relation between rates, isotropy, normal hyperbolicity that we illustrate in an example. A fuller theory is being developed incorporating other ingredients.
Corollary 6.10.
Assume the setting in Section 3.1 without (iii) and (iv), and that $\Lambda$ is an invariant manifold satisfying (B) and (H).
Then:
(6.18) $\begin{split}&\ (\mu_{+}^{*})^{2}\eta^{-1}<1\quad\textrm{OR}\quad\ (\mu_{-}^{*})^{2}\eta<1\\ &\phantom{AAAAA}\implies\\ &\omega|_{\Lambda}=0\\ &\phantom{AAAAA}\implies\\ &\mu_{+}^{*}\lambda_{+}^{*}\eta^{-1}\geq 1\quad\textrm{OR}\quad\mu_{-}^{*}\lambda_{-}^{*}\eta\geq 1.\end{split}$
In particular, $\Lambda$ is not normally hyperbolic.
Proof.
We start by proving the first implication in (6.18)
If $(\mu_{+}^{*})^{2}\eta^{-1}<1$, then for some $\mu_{+}$ with $(\mu_{+})^{2}\eta^{-1}<1$ and some $C>0$, we have for all $v\in T_{x}\Lambda$,
$\|Df^{n}(x)v\|\leq C\mu_{+}^{n}\|v\|\quad\textrm{ for }n>0.$
Hence, using (6.1) and taking the limit as $n\to\infty$, we have that $\omega(x)(u,v)=0$, $\forall u,v\in T_{x}\Lambda$, $\forall x\in\Lambda$.
The identical argument for $(\mu_{-}^{*})^{2}\eta<1$, taking $n\to$-$\infty$ is left to the reader.
The second implication in (6.18) is just the failure of (S).
Finally, we note that if $\Lambda$ is a NHIM, then by (2.11) and (2.12), we have $\lambda_{+}^{*}<\mu_{+}^{*}$, therefore $(\mu_{+}^{*})^{2}\eta^{-1}<1$ implies $\mu_{+}^{*}\lambda_{+}^{*}\eta^{-1}<1$; similarly, $(\mu_{-}^{*})^{2}\eta<1$ implies $\mu_{-}^{*}\lambda_{-}^{*}\eta<1$. This contradicts the last conclusion of (6.18). Hence, $\Lambda$ cannot be normally hyperbolic. ∎
Isotropic, specially Lagrangian manifolds have extra properties among rates that are incompatible with normal hyperbolicity. In this paper, we will only mention an example, and postpone a fuller exploration involving other concepts to future work.
Example 6.11.
Consider $\mathbb{R}^{10}$ endowed with the form $\omega=\sum_{j}dx_{j}\wedge dy_{j}$ and the map $f$ given by:
$\begin{split}f&(x_{1},x_{2},x_{3},x_{4},x_{5},y_{1},y_{2},y_{3},y_{4},y_{5})\\ =&\left(\lambda_{+}x_{1},\mu_{-}^{-1}x_{2},\mu x_{3},\mu_{+}x_{4},\lambda_{-}^{-1}x_{5},\lambda_{+}^{-1}\eta y_{1},\mu_{-}\eta y_{2},\mu^{-1}\eta y_{3},\mu_{+}^{-1}\eta y_{4},\lambda_{-}\eta y_{5}\right).\end{split}$
We just assume
$\lambda_{+}<\mu_{-}^{-1}\leq\mu\leq\mu_{+}<\lambda_{-}^{-1}.$
The $3$-D manifold $\Lambda$ corresponding to the variables $x_{2},x_{3},x_{4}$ (and the other variables set to zero) is an invariant manifold which is isotropic.
The key point of the example is that we have introduced an intermediate rate $\mu$ in $\Lambda$. The presence of a rate $\mu$ along the manifold forces the presence of a rate $\mu^{-1}\eta$ in the normal bundle.
If $\mu_{-}^{-1}<\mu^{-1}\eta<\mu_{+}$ (which in our case could well happen) we obtain that the presence of vectors with an intermediate rate $\mu$ as above is incompatible with $\Lambda$ being normally hyperbolic.
Similar phenomena appear in the use of automatic reducibility in whiskered tori [CallejaCL20].
6.7. Vanishing lemmas for derivatives of a general $2$-form $\omega$
For a general $2$-form $\omega$ (which may be non-closed or be degenerate) we have that
$f^{*}\omega=\eta\omega\quad\implies\quad f^{*}(d\omega)=\eta(d\omega).$
Hence, procedures similar to those used to prove Lemma 6.8 can be applied to obtain a vanishing Lemma 6.12, where we assume that $d\omega$ is bounded and some adequate assumptions on rates, and we conclude that $d\omega$ vanishes on several blocks.
This will be enough to give a proof of a variant of part (B) of Theorem 3.3 in Section 8.2.6 under the assumptions of Lemma 6.12. In particular, the proof in Section 8.2.6 does not assume that $\omega$ is closed and can be extended to cases studied in [wojtkowski1998conformally].
Lemma 6.12.
We make the standing assumptions from Section 3.1 with (U5’) instead of (U5). We assume only that $\omega$ is a $2$-form not necessarily closed or non-degenerate.
We assume that $f$ satisfies (2.3) and hence satisfies also:
(6.19) $f^{*}(d\omega)=\eta(d\omega).$
Then we have:
  • (A)
    If
    (6.20) $\mu_{+}^{2}\lambda_{+}\eta^{-1}<1,$
    then for every $y\in W^{s}_{\Lambda}$ and $x\in\Lambda$ such that $y\in W^{s}_{x}$, for all $v_{t},w_{t}\in T_{y}W^{s}_{\Lambda}$, and for all $u_{s}\in T_{y}W^{s}_{x}$, we have:
    (6.21) $d\omega(y)(v_{t},w_{t},u_{s})=0.$
    Analogously, if:
    (6.22) $\mu_{-}^{2}\lambda_{-}\eta<1.$
    then, for every $y\in W^{u}_{\Lambda}$ and $x\in\Lambda$ such that $y\in W^{u}_{x}$, for all $v_{t},w_{t}\in T_{y}W^{u}_{\Lambda}$, and for all $u_{u}\in T_{y}W^{u}_{x}$, we have:
    (6.23) $d\omega(y)(v_{t},w_{t},u_{u})=0.$
  • (B)
    (6.24) $\begin{split}\lambda_{+}^{3}\eta^{-1}<1&\implies\forall x\in\Lambda,\ d(\omega_{\mid W^{s}_{x}})=0,\\ \lambda_{-}^{3}\eta<1&\implies\forall x\in\Lambda,\ d(\omega_{\mid W^{u}_{x}})=0.\end{split}$
Proof.
As in the previous lemmas, we observe that, because of (6.19), we have for all $n\in\mathbb{Z}$ (and for all $y$ and all $(v,w,u)$),
$d\omega(y)(v,w,u)=\eta^{-n}d\omega(f^{n}(y))(Df^{n}(y)v,Df^{n}(y)w,Df^{n}(y)u)$
With the respective assumptions on rates and uniform boundedness of the derivative $d\omega$, we obtain the desired result taking the limit $n\to\pm\infty$ in the different cases, as we did in the proofs of Lemma 6.5 and Lemma 6.8. ∎
6.8. Vanishing lemmas for some examples of unbounded symplectic forms
In this section, we develop a result Lemma 6.13,which is very similar to Lemma 6.5, but which applies to some unbounded symplectic forms. The system is assumed to have a compact invariant set $\mathcal{A}$ – that serves as the origin to measure distances, e.g. $\mathcal{A}$ could be a fixed point. We also assume that the symplectic form at a point $x$ is bounded by a power $\alpha$ of the distance from $x$ to $\mathcal{A}$ and that the hyperbolicity rates together with $\alpha$ and $\eta$ satisfy some relations.
We hope that Lemma 6.13 indicates the ingredients needed in a systematic theory dealing with unbounded forms. However, this result will not be used in this paper.
Lemma 6.13 (Infinitesimal Vanishing Lemma for Some Unbounded Forms).
We take as granted the standing assumptions from Section 3.1 without (U5).
We assume that there exists a compact invariant set $\mathcal{A}\subset\Lambda$ such that for some $A,B,\alpha>0$ we have for all $x\in\Lambda$
(6.25) $\|\omega(x)\|\leq B+A\cdot d(x,\mathcal{A})^{\alpha},$
where $d$ is the Riemannian distance measured along $\Lambda$.
Then, for all $x\in\Lambda$, $v^{1}_{s},v^{2}_{s}\in E^{s}_{x}$, $v^{1}_{u},v^{2}_{u}\in E^{u}_{x}$,, $v^{1}_{t},v^{2}_{t}\in T_{x}\Lambda$, we have:
(6.26) $\begin{split}\mu_{+}^{1+\alpha}\lambda_{+}\eta^{-1}<1&\implies\omega(x)(v^{1}_{t},v^{1}_{s})=0,\\ \mu_{-}^{1+\alpha}\lambda_{-}\eta<1&\implies\omega(x)(v^{1}_{t},v^{1}_{u})=0,\\ \mu_{+}^{\alpha}\lambda_{+}^{2}\eta^{-1}<1&\implies\omega(x)(v^{1}_{s},v^{2}_{s})=0,\\ \mu_{-}^{\alpha}\lambda_{-}^{2}\eta<1&\implies\omega(x)(v^{1}_{u},v^{2}_{u})=0,\\ \mu_{+}^{2+\alpha}\eta^{-1}<1&\implies\omega(x)(v^{1}_{t},v^{2}_{t})=0,\\ \mu_{-}^{2+\alpha}\eta<1&\implies\omega(x)(v^{1}_{t},v^{2}_{t})=0.\end{split}$
Remark 6.14.
One may wonder whether the assumptions of Lemma 6.13 are contradictory. An example in $M={\mathbb{R}}\times{\mathbb{T}}$ is obtained by choosing any function $h:{\mathbb{R}}\to{\mathbb{R}}$, satisfying $|h(I)|\leq|I|^{\alpha}$, and setting $\omega=h(I)dI\wedge d\theta$. We consider a map of the form $f(I,\theta)=(g(I),\theta)$. If $h$, $g$ satisfy the separable differential equation $h(g(I))g^{\prime}(I)=\eta h(I)$, with $g(0)=0$ then the map $f$ is conformally symplectic for $\omega$ and the set $\mathcal{A}=\{(0,\theta)\}$ satisfies the hypotheses of the lemma. We need to choose $h$ so that the solution $g$ gives a diffeomorphism.
Proof.
Since $\mathcal{A}$ is a compact set, for every $x\in\Lambda$ we have
$d(x,\mathcal{A})=\inf_{y\in\mathcal{A}}d(x,y)<+\infty.$
Condition (H) implies that, for some constants $\tilde{D}_{+},\tilde{D}_{-}>0$ independent of $x$, we have
[…]
From (6.27) and (6.25), for $n\geq 0$, we have
$\begin{split}\sup_{x\in\Lambda}\|\omega(f^{n}(x))\|\leq&B+A\cdot d(f^{n}(x),\mathcal{A})^{\alpha}\leq B+A(\tilde{D}_{+})^{\alpha}\mu_{+}^{\alpha n}d(x,\mathcal{A})^{\alpha}\\ =&B+A^{+}_{x}\mu_{+}^{\alpha n},\\ \sup_{x\in\Lambda}\|\omega(f^{-n}(x))\|\leq&B+A\cdot d(f^{-n}(x),\mathcal{A})^{\alpha}\leq B+A(\tilde{D}_{-})^{\alpha}\mu_{-}^{\alpha n}d(x,\mathcal{A})^{\alpha}\\ =&B+A^{-}_{x}\mu_{-}^{\alpha n},\end{split}$
where $A^{+}_{x}=A(\tilde{D}_{+})^{\alpha}\cdot d(x,\mathcal{A})^{\alpha}$ and $A^{-}_{x}=A(\tilde{D}_{-})^{\alpha}\cdot d(x,\mathcal{A})^{\alpha}$.
Using (6.1), for $v^{1}_{s},v^{2}_{s}\in E^{s}_{x}$, $v^{1}_{t},v^{2}_{2}\in T_{x}\Lambda$, we have for $n\geq 0$:
(6.28) $\begin{split}&|\omega(x)(v^{1}_{s},v^{2}_{s})|\leq\left[\tilde{B}(\eta^{-1}\lambda_{+}^{2})^{n}+\tilde{A}_{x}(\eta^{-1}\lambda_{+}^{2}\mu_{+}^{\alpha})^{n}\right]\|v^{1}_{s}\|\|v^{2}_{s}\|,\\ &|\omega(x)(v^{1}_{s},v^{1}_{t})|\leq\left[\tilde{B}(\eta^{-1}\lambda_{+}\mu_{+})^{n}+\tilde{A}_{x}(\eta^{-1}\lambda_{+}\mu_{+}^{1+\alpha})^{n}\right]\|v^{1}_{s}\|\|v^{1}_{t}\|,\\ &|\omega(x)(v^{1}_{t},v^{2}_{t})|\leq\left[\tilde{B}(\eta^{-1}\mu_{+}^{2})^{n}+\tilde{A}_{x}(\eta^{-1}\mu_{+}^{2+\alpha})^{n}\right]\|v^{1}_{t}\|\|v^{2}_{t}\|,\\ &|\omega(x)(v^{1}_{t},v^{2}_{t})|\leq\left[\tilde{B}(\eta\mu_{-}^{2})^{n}+\tilde{A}_{x}(\eta\mu_{-}^{2+\alpha})^{n}\right]\|v^{1}_{t}\|\|v^{2}_{t}\|,\end{split}$
for some constant $\tilde{B}>0$ and some $\tilde{A}_{x}>0$ that is independent of $n$ but depends on $d(x,\mathcal{A})^{\alpha}$.
Under the hypothesis on the rates and using (N), the limit as $n\rightarrow+\infty$ of the right-hand-size of (6.28) is zero.
A similar argument gives the vanishing of the symplectic form in the case when one of the tangent vectors is unstable (or both are unstable). ∎
7. Proof of Theorem 3.1
7.1. Proof of Theorem 3.1 (A) on the symplecticity of the NHIM
We want to show that $\omega_{\Lambda}:=\omega_{\mid\Lambda}$ is a symplectic form and hence $\Lambda$ is symplectic.
Since the exterior derivative commutes with restriction to submanifolds, we have:
$d(\omega_{\mid\Lambda})=(d\omega)_{\mid\Lambda}=0,$
so to prove that $\omega_{\Lambda}$ is symplectic we only have to prove that $\omega_{\Lambda}$ is non-degenerate.
For $x\in\Lambda$, if $v_{*}\in T_{x}\Lambda$ and $\omega(x)(v_{*},v_{t})=0$ for all $v_{t}\in T_{x}\Lambda$, then, as (S) is satisfied, we can apply Lemma 6.5, and we have $\omega(x)(v_{*},v)=0$ for all $v\in T_{x}M$. By the non-degeneracy of $\omega$ in $TM$ we conclude $v_{*}=0$.
The dynamics on $\Lambda$ is conformally symplectic because $f^{*}\omega=\eta\omega$ and the restriction commutes with the pullback. Hence, $(f_{\mid\Lambda})^{*}\omega_{\Lambda}=\eta\omega_{\Lambda}$.
7.2. Proof of Theorem 3.1 (B) on pairing rules
In this section we show that the geometry imposes certain symmetries on the possible rates. In the case of symplectic maps, these symmetries (and their proofs) have been folklore but we have not been able to locate a specific reference. Here we derive the symmetries for conformally symplectic maps, and note that the proof also applies to the symplectic case. For conformally symplectic systems, there are arguments for periodic orbits and for Lyapunov exponents [dettmann1996proof, wojtkowski1998conformally], but the argument here is different and is based on the vanishing lemmas.
We are under the assumption that $\Lambda$ is symplectic and therefore $\omega(x)$ is non-degenerate for any $x\in\Lambda$.
For the optimal rate $\mu_{+}^{*}$ we have that $\forall\varepsilon>0,\,\exists D_{+}=D_{+}(\varepsilon)$ such that:
$\forall x\in\Lambda\,\forall u\in T_{x}\Lambda\,\|Df^{n}(x)u\|\leq D_{+}(\mu_{+}^{*}+\varepsilon)^{n}\|u\|,\ \forall n\geq 0.$
Taking $x\in\Lambda$ and applying Lemma 6.3 for $L=\Lambda$, $\alpha=\mu_{+}^{*}+\varepsilon$ and $\beta=\frac{\eta-\varepsilon}{\mu_{+}^{*}+\varepsilon}$, as $\alpha\beta<\eta$, we obtain there exists $v\in T_{x}\Lambda$ where $\omega(x)(u,v)\neq 0$ and there exists $D_{2}>0$ such that
$\|Df^{n}(x)v\|\geq D_{2}\left(\frac{\eta-\varepsilon}{\mu_{+}^{*}+\varepsilon}\right)^{n}\|v\|,\ \forall n\geq 0.$
Since $\mu_{-}^{*}$ is defined as an optimal rate, by Lemma B.1 we have
$\frac{1}{\mu_{-}^{*}}\geq\frac{\eta-\varepsilon}{\mu_{+}^{*}+\varepsilon},$
and, since this holds for all $\varepsilon>0$, we obtain
$\frac{1}{\mu_{-}^{*}}\geq\frac{\eta}{\mu_{+}^{*}}.$
Applying the same argument for the inverse map $f^{-1}$ we also have
$\frac{1}{\mu_{+}^{*}}\geq\frac{\eta^{-1}}{\mu_{-}^{*}}.$
We conclude
(7.1) $\frac{\mu_{+}^{*}}{\mu_{-}^{*}}=\eta.$
A similar argument, which we now detail, yields
$\frac{\lambda_{+}^{*}}{\lambda_{-}^{*}}=\eta.$
For the optimal rate $\lambda_{+}^{*}$ we have that $\forall\varepsilon>0,\,\exists C_{+}=C_{+}(\varepsilon)$ such that:
$\forall x\in\Lambda\,\forall u\in E^{s}_{x}\,\|Df^{n}(x)v\|\leq C_{+}(\lambda_{+}^{*}+\varepsilon)^{n}\|v\|,\ \forall n\geq 0.$
Taking $x\in\Lambda$ and applying Lemma 6.3 for $\alpha=\lambda_{+}^{*}+\varepsilon$ and $\beta=\frac{\eta-\varepsilon}{\lambda_{+}^{*}+\varepsilon}$, as $\alpha\beta<1$, we conclude that there is a vector $w\in T_{x}M$ such that $\omega(x)(v,w)\neq 0$ and a constant $C_{2}>0$ such that
(7.2) $\|Df^{n}(x)w\|\geq C_{2}\left(\frac{\eta-\varepsilon}{\lambda_{+}^{*}+\varepsilon}\right)^{n}\|w\|,\ \forall n\geq 0.$
Let $w=w^{u}+w^{ts}$, where $w^{u}\in E^{u}_{x}$ and $w^{ts}\in T_{x}W^{s}_{\Lambda}$.
Using (7.2), B.2 and (7.1) we obtain
$\begin{split}\|Df^{n}(x)w^{u}\|\geq&\|Df^{n}(x)w\|-\|Df^{n}(x)w^{ts}\|\\ \geq&C_{2}\left(\frac{\eta-\varepsilon}{\lambda_{+}^{*}+\varepsilon}\right)^{n}\|w\|-D_{+}(\mu^{*}_{+}+\varepsilon)^{n}\|w^{ts}\|\\ =&C_{2}\left(\frac{\eta-\varepsilon}{\lambda_{+}^{*}+\varepsilon}\right)^{n}\|w\|-D_{+}(\eta\mu^{*}_{-}+\varepsilon)^{n}\|w^{ts}\|\\ \geq&C_{3}\left(\frac{\eta-\varepsilon}{\lambda_{+}^{*}+2\varepsilon}\right)^{n}\|w^{u}\|,\end{split}$
for $n\geq 0$ sufficiently large and $\varepsilon$ sufficiently small, where the last inequality is due to the fact that $\lambda^{*}_{+}\mu^{*}_{-}<1$. Note that $w^{u}\neq 0$ because we have upper bounds for the growth of $w^{ts}$ which are incompatible with the lower bounds for the growth of $w$.
Since any uniform bound $\lambda_{-}$ with
$\|Df^{n}(x)w^{u}\|\geq\tilde{C}(\lambda_{-})^{-n}\|w^{u}\|$
has to satisfy
$\lambda_{-}^{*}\leq\lambda_{-}$
we conclude, by the same argument as before (letting $\varepsilon\to 0$) that
$\lambda_{-}^{*}\leq\eta^{-1}\lambda_{+}^{*}.$
Applying this result to $f^{-1}$ in place of $f$ we obtain the desired result. ∎
Remark 7.1.
It is interesting to compare the proofs of pairing rules for rates above with the proofs of pairing rules for periodic orbits or for Lyapunov exponents in [Dessler88, dettmann1996proof, wojtkowski1998conformally]. The proofs in the above references are based on defining the operator $J_{x}:T_{x}M\rightarrow T_{x}M$ by $\omega(x)(u,v)=g_{x}(u,J_{x}v)$ where $g$ is the Riemannian metric.
Then, the conformal symplectic property of the map is translated into $Df^{n}(x)^{T}J_{f^{n}(x)}Df^{n}(x)=\eta^{n}J_{x}$, where the transpose is with respect to the metric. Hence,
(7.3) $Df^{n}(x)=\eta^{n}J^{-1}_{f^{n}(x)}(Df^{n}(x))^{-T}J_{x}.$
We think of (7.3) as a relation among linear operators in tangent spaces. In the literature, sometimes, (7.3) is described as a relation among matrices using a global frame (introduced already in the setup). We emphasize that (7.3) has an intrinsic meaning without a global frame.
The equation (7.3) relates the rates of growth of $Df^{n}(x)$ and those of $(Df^{n}(x))^{-T}$ leading to pairing rules. Using (7.3) to relate asymptotic rates, seems to require that $J^{-1}_{f^{n}(x)}$ is uniformly bounded.
The method we use here to obtain pairing rules does not require that $\|J^{-1}_{x}\|$ is uniformly bounded nor the existence off a global frame.
8. Proof of Theorem 3.3
8.1. Proof of Theorem 3.3 (A) on symplecticity of the homoclinic channel
We first prove that, if $\Gamma$ is a homoclinic channel (see Definition 2.22), then $\omega_{\mid\Gamma}$ is non-degenerate, hence $(\Gamma,\omega_{\mid\Gamma})$ is a symplectic manifold.
Conditions (S) allow to apply part (A) of Theorem 3.1 obtaining that $\Lambda$ is symplectic.
If $\Gamma$ is sufficiently $\mathcal{C}^{1}$-close to $\Lambda$, from $\omega_{\mid\Lambda}$ being non-degenerate we deduce $\omega_{\mid\Gamma}$ is non-degenerate.
If $\Gamma$ is not $\mathcal{C}^{1}$-close to $\Lambda$, by the Fiber Contraction Theorem (see Lemma C.1) we have
(8.1) $d_{\mathcal{C}^{1}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{+}\mu_{-})^{n},\textrm{ for }n\geq 0.$
Then, there exists $N>0$ such that $f^{N}(\Gamma)$ is sufficiently $\mathcal{C}^{1}$-close to $\Lambda$ so that $\omega_{\mid f^{N}(\Gamma)}$ is non-degenerate as in the previous case. Since $f$ is conformally symplectic we have
$\omega_{\mid f^{N}(\Gamma)}=(f^{*})^{N}\omega_{\mid\Gamma}=\eta^{N}\omega_{\mid\Gamma}.$
Since $\omega_{\mid f^{N}(\Gamma)}$ is non-degenerate it follows that $\omega_{\mid\Gamma}$ is non-degenerate.
8.2. Proof of Theorem 3.3 (B) on symplecticity of the scattering map
In this section we give seven different proofs of Theorem 3.3 (B) or of some versions of it (some versions do not assume that $\omega$ is closed or use different assumptions on rates or boundedness of the derivatives of $\omega$). We note that some of these proofs do not use that $\omega$ is non-degenerate, so they work without change in the presymplectic case (see Section 10) just taking into account that the conformal factor can be a function.
The first proof, given in Section 8.2.1 is based on vanishing lemmas. A proof adapting the one from [DLS08] from the symplectic case to the conformally symplectic case is given in Section 8.2.2. In Section 8.2.3 we give a proof based on the system of coordinates defined in Section 2.8. In Section 8.2.4 we give a proof based on Cartan’s magic formula. These four proofs use the standing assumptions from Section 3.1 and conditions (S). They use strongly that $\omega$ is closed, but they do not use that $\omega$ is non-degenerate, so that these proofs apply to presymplectic forms.
A fifth proof, given in Section 8.2.5, uses the study of graphs, but requires (U5) and different rate conditions. We give a sixth proof based on vanishing lemmas in Section 8.2.6, which does not use that $\omega$ is a closed form, but also requires (U5) and different rate conditions. In Section 8.2.7 we give a seventh proof, based on iterations, which also uses (U5).
We also give two proofs of part (C) in Section 8.3. We note that they are based on vanishing lemmas. The first one, given in Section 8.3.1 and based on Stokes’ Theorem, uses that $\omega$ is exact (hence closed) but it does not use that $\omega$ is non-degenerate. The second one, given in Section 8.3.2 also uses that $\Omega$ is exact and uses Cartan’s magic formula.
Remark 8.1.
To prove that
$(\Omega_{+})^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\textrm{loc}}_{\Lambda}}$
it is enough to work on $W^{s,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$. The reason is that, taking $n>0$ big enough but fixed, $f^{n}(W^{s,\textrm{loc}}_{\Lambda})\subset\mathcal{O}_{\rho}$ and, by (2.24) we obtain that:
$(\Omega_{+})_{\mid W^{s,\textrm{loc}}_{\Lambda}}=f^{-n}_{\mid\Lambda}\circ(\Omega_{+})_{\mid f^{n}(W^{s,\textrm{loc}}_{\Lambda})}\circ f^{n}_{\mid W^{s,\textrm{loc}}_{\Lambda}}$
and therefore we obtain the equality in all $W^{s,\textrm{loc}}_{\Lambda}$ and indeed on $W^{s}_{\Lambda}$.
8.2.1. A proof of Theorem 3.3 (B) based on Stokes’ theorem
We prove that the $2$-form $\omega$ is invariant under the pullback of $\Omega_{+}$, as the proof for $\Omega_{-}$ is analogous.
Figure 4.
It is enough to take any $y\in W^{s,\textrm{loc}}_{\Lambda}$, and any two tangent vectors $v^{1},v^{2}\in T_{y}W^{s}_{\Lambda}$. We will prove that
(8.2) $(\Omega_{+}^{*}\omega)(y)(v^{1},v^{2})=\omega(y)(v^{1},v^{2}).$
We define a $2D$-cell [^10] A concrete, but slightly more costly in the regularity is to write explicitly the cell as $\Sigma(t_{1},t_{2})=\exp_{y}(\varepsilon(t_{1}v^{1}+t_{2}v^{2})),\quad 0\leq t_{1},t_{2}\leq 1$, for $0<\varepsilon$ sufficiently small, where now $\exp$ denotes the exponential mapping for the metric restricted to $W^{s}_{\Lambda}$. $\Sigma\subseteq W^{s,\textrm{loc}}_{\Lambda}$ in such a way that it is tangent to $v_{1}$ and $v_{2}$ using the uniform system of coordinates assumed to exist in (U1) and restricting it to $W^{s,\mathrm{loc}}_{\Lambda}$.
By the transversality assumption on the vectors $v^{1},v^{2}$, it follows that different points $a\in{\Sigma}$ project onto different points $\Omega_{+}(a)\in\Lambda$.
Define a $3D$-cell $\tilde{\Sigma}$ in $W^{s,\textrm{loc}}_{\Lambda}$ by
(8.3) $\tilde{\Sigma}(t,t_{1},t_{2})=\gamma(t;\Sigma(t_{1},t_{2}),\Omega_{+}(\Sigma(t_{1},t_{2}))),\quad 0\leq t,t_{1},t_{2}\leq 1$
where
[…]
The family of paths is chosen so that they depend smoothly on $a\in\Sigma$ and that $\tilde{\Sigma}$ forms a $3D$-cell inside $W^{s}_{\Lambda}$. The projection $\Omega_{+}(\Sigma)$ is a $2D$-cell inside $\Lambda$. See Fig. 4. Let $c(s)$ be a piecewise smooth parametrization of $\partial\Sigma$, for $s\in\partial([0,1]^{2})$.
Since $\omega$ is a closed form, using Stokes Theorem we compute
(8.5) $0=\int_{\tilde{\Sigma}}d\omega=\int_{\partial\tilde{\Sigma}}\omega=\int_{\Sigma}\omega-\int_{\Omega_{+}(\Sigma)}\omega+\int_{\Upsilon}\omega,$
where $\Upsilon$ is the $2D$-cell completing the boundary of $\tilde{\Sigma}$. We consider it parameterized by
$\Upsilon(t,s)=\gamma(t,c(s)),\textrm{ for }(t,s)\in[0,1]\times\partial([0,1]^{2}).$
Now we compute the integral along $\Upsilon$:
$\begin{split}\int_{\Upsilon}\omega=&\int_{s\in\partial([0,1]^{2})}\int_{t\in[0,1]}\omega(\Upsilon(t,s))(\partial_{t}\Upsilon(t,s),\partial s\Upsilon(t,s))dtds=0,\end{split}$
where we have used that
$\left(\partial_{t}\Upsilon(t,s),\partial s\Upsilon(t,s)\right)\in T_{\Upsilon(t,s)}W^{s}_{\Omega_{+}(\Upsilon(t,s))}\times T_{\Upsilon(t,s)}W^{s,\mathrm{loc}}_{\Lambda},$
and therefore, conditions (S) allow to apply Lemma 6.8 obtaining that $\omega(\Upsilon(t,s))(\partial_{t}\Upsilon(t,s),\partial_{s}\Upsilon(t,s))=0$.
In conclusion
(8.6) $\int_{\Sigma}\omega=\int_{\Omega_{+}(\Sigma)}\omega.$
Since (8.6) holds for any $2$-cell $\Sigma$ that is transverse to the fiber $W^{s}_{\Omega_{+}(y)}$, and for any $y\in W^{s,\textrm{loc}}_{\Lambda}$, it follows that
$\Omega_{+}:W^{s,\textrm{loc}}_{\Lambda}\to\Lambda$
satisfies $\Omega_{+}^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\textrm{loc}}_{\Lambda}}$.
Finally, as $\Omega_{+}^{\Gamma}$ is a restriction of this map to the symplectic manifold $\Gamma$, it is symplectic.
An analogous reasoning gives that $\Omega_{-}^{\Gamma}$ is symplectic, and then, $(\Omega_{-}^{\Gamma})^{-1}$ is also symplectic and therefore $S=\Omega_{+}^{\Gamma}\circ(\Omega_{-}^{\Gamma})^{-1}$ is symplectic.
8.2.2. A proof of Theorem 3.3 (B) by adapting the proof of [DLS08] from the symplectic case to the conformally symplectic case
The proof of [DLS08] uses a similar geometric construction as the proof in Section 8.2.1. The paper [DLS08] starts from the same cell depicted in Fig. 4 and obtains the desired result by showing that the integral of $\omega$ over $\Upsilon$ is zero.
The vanishing of this 2D integral is obtained using that for every $n>0$
$\int_{\Upsilon}\omega=\eta^{-n}\int_{\Upsilon}({f^{n}})^{*}\omega=\eta^{-n}\int_{f^{n}(\Upsilon)}\omega.$
We now observe that the Riemannian area of $f^{n}(\Upsilon)$ is bounded from above by $C(\lambda_{+}\mu_{+})^{n}$. Since $\omega$ is bounded, under the rate conditions (S), we obtain that $\eta^{-n}\left|\int_{f^{n}(\Upsilon)}\omega\right|$ can be made as small as desired by taking $n$ large.
The proof in Section 8.2.1, can be considered as a “disintegration” of the argument in [DLS08]. We can think of the vanishing lemma as dividing $\Upsilon$ into infinitesimal cells and showing that each infinitesimal integral is exactly zero. Proving first the infinitesimal result gives more flexibility and the vanishing lemmas are used also to prove the pairing rules.
8.2.3. A proof of Theorem 3.3 (B) based on a system of coordinates
A more explicit version of Lemma 6.8 can be obtained by using the system of coordinates defined in Section 2.8, which exists thanks to hypothesis (U1). Using these coordinates we can make the symplectic form $\omega$ explicit.
Recall that the coordinate system $\varphi$ is so that
$\{\varphi(x,y)\,|\,y\in B_{\rho}(0)\}=W^{s,\mathrm{loc}}_{x}.$
Note that the representation of $\Omega_{+}$ in this system of coordinates is given by:
(8.7) $\Omega_{+}(x,y)=(x,0)$
The coordinate $x$ is defined precisely as taking the projection $\Omega_{+}$, therefore $\Omega_{+}$ is represented by setting the $y$ coordinate to 0.
We can identify:
$T_{(x,y)}W^{s}_{\Lambda}=\{(t,w),\,t\in\mathbb{R}^{d_{c}},\,w\in\mathbb{R}^{d_{s}}\},$
so we can choose a basis $(t_{1},0),\dots(t_{d_{c}},0),(0,w_{1}),\dots(0,w_{d_{s}})$ of $T_{(x,y)}W^{s}_{\Lambda}$ independent of the point $(x,y)$.
It is important to remark that, in these coordinates, we can use the results of Lemma B.3 so that
(8.8) $\begin{split}\|Df^{n}(x,y)(t,0)\|&\leq D_{+}\mu_{+}^{n}\|t\|,\\ \|Df^{n}(x,y)(0,w)\|&\leq C_{+}\lambda_{+}^{n}\|w\|,\end{split}\quad n\geq 1.$
By hypotheses (S), we can use (6.11) of the Vanishing Lemma 6.8, to obtain that the symplectic form can be represented as:
(8.9) $\omega(x,y)=\begin{pmatrix}\omega_{xx}(x,y)&0\\ 0&0\end{pmatrix}$.
When $\omega_{\mid\Lambda}$ is non-degenerate, the kernel of $\omega|_{W^{s,\textrm{loc}}_{\Lambda}}$ in this neighborhood is the tangent to the $W^{s}_{x}$ leaves of the strong stable foliation  [^11] This is consistent with Lemma 3.6 that shows that the kernel of a presymplectic form integrates to a foliation..
Now we proceed to the proof of Theorem 3.3 (B).
We have the representation of $\omega$ by (8.9).
The following is the key observation: if the symplectic form $\omega$ is closed, expressing the differential in coordinates we then have:
(8.10) $\partial_{y}\omega_{xx}(x,y)=0.$
To show this, take a sufficiently small patch $U\subset\Lambda$, where we can trivialize $E^{s}$. Since $d\omega=0$, we have $d(\omega_{\mid W^{s}_{\Lambda}})=0$ which expressed in coordinates gives
$\begin{split}0=\sum_{i}\sum_{j<k}\partial_{x_{i}}\omega_{x_{j}x_{k}}dx_{i}\wedge dx_{j}\wedge dx_{k}+\sum_{l}\sum_{j<k}\partial_{y_{l}}\omega_{x_{j}x_{k}}dy_{l}\wedge dx_{j}\wedge dx_{k}.\end{split}$
As the terms in the above sum are linearly independent, it follows that $\partial_{y_{l}}\omega_{x_{j}x_{k}}(x,y)=0$ for all $l$ and $j<k$. This shows (8.10).
Therefore, $\omega_{xx}$ depends only on $x$. Since in this system of coordinates we have $\Omega_{+}(x,y)=(x,0)$ (see (8.7)), we obtain directly that, if $n$ sufficiently large and $\mathscr{C}\subset f^{n}(\Gamma)$ is a $2$-cell, then
$\omega(\Omega_{+}(\mathscr{C}))=\omega(\mathscr{C}).$
8.2.4. Proof of part (B) of Theorem 3.3 based on Cartan’s magic formula
The following proof is similar in spirit to the one in Section 8.2.3 but avoids the construction of a system of coordinates.
To prove that $\Omega_{+}:W^{s,\mathrm{loc}}_{\Lambda}\to\Lambda$ satisfies $(\Omega_{+})^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\mathrm{loc}}_{\mid\Lambda}}$, we proceed as follows. Take any section $\Psi\subset W^{s,\mathrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$ transversal to the foliation (2.16) (see condition (2.29)) and consider the restricted wave map
$\Omega_{+}^{\Psi}\equiv(\Omega_{+})_{\mid\Psi}:\Psi\to\Lambda.$
We will see that $\Omega_{+}^{\Psi}$ satisfies: $(\Omega_{+}^{\Psi})^{*}(\omega_{\mid\Lambda})=\omega_{\mid\Psi}$.
As usual, we can assume that $\Psi$ is $\mathcal{C}^{1}$ close to $\Lambda$ and use a finite number of iterates to get to others.
For $x\in\Omega_{+}(\Psi)\subset\Lambda$, using the implicit function theorem, we can associate unique $\gamma(x)\in\Psi$ and $v(x)\in T_{x}W^{s}_{x}=E^{s}_{x}$ in such a way that $\gamma(x)=\exp_{x}(v(x))$, where the exponential is along $W^{s}_{x}$ and $v(x)$ is required to be in a sufficiently small ball. Both $\gamma$ and $v$ depend on $x\in\Omega_{+}(\Psi)\subset\Lambda$ in a continuously differentiable way. As we mention, we can always restrict $\Psi$ so that $\Omega_{+}(\Psi)$ is bounded.
Consider the $\mathcal{C}^{1}$ family of mappings $\phi_{t}:\Omega_{+}(\Psi)\to W^{s}_{\Lambda}$, indexed by $t\in[0,1]$:
$\phi_{t}(x)=\exp_{x}(tv(x)).$
Clearly, $\phi_{0}(x)=x,\phi_{1}(x)=\gamma(x)$, and, more succinctly, $\phi_{0}={\rm Id}$ and $\phi_{1}=(\Omega_{+}^{\Psi})^{-1}$.
We let $\frac{d}{dt}\phi_{t}=V\circ\phi_{t}$, where $V(\phi_{t}(x))$ is tangent to $W^{s}_{x}$ at $\phi_{t}(x)$. This defines $V$ as a $\mathcal{C}^{1}$ vector field on some domain in $W^{s}_{\Lambda}$.
We now compute, using Cartan’s magic formula
$\frac{d}{dt}\left(\phi^{*}_{t}\omega\right)=\phi_{t}^{*}\left[i(V)d\omega+di(V)\omega\right].$
The first term above is zero because $\omega$ is closed. The second term is also zero by the Vanishing Lemma 6.8. Therefore
$\omega_{\mid\Lambda}=\phi_{0}^{*}(\omega_{\mid\Lambda})=\phi_{1}^{*}(\omega_{\mid\Psi})=((\Omega_{+}^{\Gamma})^{-1})^{*}(\omega_{\mid\Psi}).$
8.2.5. Proof of part (B) of Theorem 3.3 based on graphs in products of manifolds
In this section we present another proof of part (B) of Theorem 3.3, with the standing assumptions from Section 3.1 but with (U5’).
We also assume that the rates (R) do not satisfy (S) but they satisfy a different condition:
(8.11) $\mu_{-}\mu_{+}^{2}\lambda_{+}\eta^{-1}<1,\quad\mu_{+}\mu_{-}^{2}\lambda_{-}\eta<1.$
This proof is based on the study of graphs.
First, we recall some standard results. Given a pair of manifolds $(M_{1},\omega_{1})$, $(M_{2},\omega_{2})$ where $\omega_{i}$ are two $2$-forms, and a pair of maps $g_{1}:M_{1}\to M_{1}$, $g_{2}:M_{2}\to M_{2}$. We define
(8.12) $\begin{split}&\tilde{M}=M_{1}\times M_{2},\quad\tilde{\omega}=(-\omega_{1})\oplus\omega_{2},\\ &\textrm{ that \ is, \ for }\ x_{1}\in M_{1},\ x_{2}\in M_{2},\ v_{1},w_{1}\in T_{x_{1}}M_{1},\ v_{2},w_{2}\in T_{x_{2}}M_{2}:\\ &\tilde{\omega}(x_{1},x_{2})((v_{1},v_{2}),(w_{1},w_{2}))=-\omega_{1}(x_{1})(v_{1},w_{1})+\omega_{2}(x_{2})(v_{2},w_{2}),\\ &\tilde{g}:\tilde{M}\rightarrow\tilde{M},\\ &\tilde{g}(x_{1},x_{2})=(g_{1}(x_{1}),g_{2}(x_{2})),\end{split}$
Given a map $f:M_{1}\rightarrow M_{2}$, we define its graph $\mathcal{G}(f)\subset\tilde{M}$ by:
$\mathcal{G}(f)=\{(x,f(x))\,|\,x\in M_{1}\}.$
The following result is well known:
Lemma 8.2.
With the notations above, the diffeomorphism $f$ satisfies $f^{*}(\omega_{2})=\omega_{1}$ if and only if $\tilde{\omega}$ vanishes on $\mathcal{G}(f)\subset\tilde{M}$.
Proof.
The standard and easy proof of Lemma 8.2 is just to observe that $T_{(x,f(x))}\mathcal{G}(f)=\{(u,Df(x)u)\,|u\in T_{x}M_{1}\}$. Hence, $\tilde{\omega}$ vanishes on $\mathcal{G}(f)\subset\tilde{M}$ is the same as having for all $x\in M_{1}$, $u,v\in T_{x}M_{1}$,
$\begin{split}0&=\tilde{\omega}(x,f(x))((u,Df(x)u),(v,Df(x)v))\\ &=-\omega_{1}(x)(u,v)+\omega_{2}(f(x))(Df(x)u,Df(x)v)\end{split}$
If $f^{*}(\omega_{2})\neq\omega_{1}$, the form ${\tilde{\omega}}_{\mid\mathcal{G}(f)}$ does not vanish. The size of $\|\tilde{\omega}_{\mid\mathcal{G}(f)}\|$ is a measure of the failure of $f^{*}\omega_{1}=\omega_{2}$.
We now proceed with the proof of part (B) of Theorem 3.3. In this case, $M_{1}=M_{2}=M$, $\tilde{\omega}=(-\omega)\oplus\omega$, and $g_{1}=g_{2}=f:M\to M$ which satisfies: $f^{*}(\omega)=\eta\omega$, and therefore
$\tilde{f}^{*}\tilde{\omega}=\eta\tilde{\omega}$.
To prove that $\Omega_{+}:W^{s,\mathrm{loc}}_{\Lambda}\to\Lambda$ satisfies $(\Omega_{+})^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\mathrm{loc}}_{\mid\Lambda}}$, we proceed as we did in section 8.2.4, taking any section $\Psi\subset W^{s,\mathrm{loc}}_{\Lambda}$ transversal to the foliation (2.16) (see condition (2.29)) and considering the restricted wave map
$\Omega_{+}^{\Psi}\equiv(\Omega_{+})_{\mid\Psi}:\Psi\to\Lambda.$
We will see that $\Omega_{+}^{\Psi}$ satisfies: $(\Omega_{+}^{\Psi})^{*}(\omega_{\mid\Lambda})=\omega_{\mid\Psi}$.
To prove it, we take the graph of $\Omega_{+}^{\Psi}$
$\mathcal{G}(\Omega_{+}^{\Psi})\subseteq\Psi\times\Lambda\subseteq M\times M=\tilde{M}$
and prove that $\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{\Psi})}=0$.
To this end, we also consider $\textrm{Id}_{\Lambda}:\Lambda\to\Lambda$ whose graph $\mathcal{G}(\textrm{Id}_{\Lambda})=\tilde{\Lambda}=\Lambda\times\Lambda\subseteq\tilde{M}$, and note that $\tilde{\omega}_{\mid\tilde{\Lambda}}=0$.
Observe that the equivariance relation for the wave maps (2.24) when restricted to a transversal manifold $\Psi$ to the foliation gives a relation similar (2.31), that is:
(8.13) $\begin{split}&\Omega_{+}^{\Psi}=f_{\mid\Lambda}^{-n}\circ\Omega_{+}^{f^{n}(\Psi)}\circ f^{n},\quad n\geq 0\end{split}$
If we reformulate the equivariance relation (8.13) in terms of graphs, we obtain:
${\tilde{f}}^{n}(\mathcal{G}(\Omega_{+}^{\Psi}))=\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)}),\quad n\geq 0$
where $\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)})\subseteq f^{n}(\Psi)\times\Lambda\subseteq\tilde{M}$ is the graph of $\Omega^{f^{n}(\Psi)}_{+}$.
Therefore, we have for all $n\geq 0$
(8.14) $\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{\Psi})}=\eta^{-n}({\tilde{f}}^{n})^{*}(\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)})}).$
Using Lemma C.1, we have that $d_{\mathcal{C}^{1}}(f^{n}(\Psi),\Lambda)\leq C(\lambda_{+}\mu_{-})^{n}$, hence
$d_{\mathcal{C}^{1}}\left(\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)}),\mathcal{G}(\textrm{Id}_{\Lambda})\right)\leq C(\lambda_{+}\mu_{-})^{n}.$
Using (U5) and that $\tilde{\omega}_{\mathcal{G}(\textrm{Id}_{\Lambda})}=0$, we have:
$\begin{split}\|\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)})}\|_{\mathcal{C}_{0}}&=\|\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)})}-\tilde{\omega}_{\mid\mathcal{G}(\textrm{Id}_{\Lambda})}\|_{\mathcal{C}_{0}}\\ &\leq\|\tilde{\omega}\|_{\mathcal{C}^{1}}d_{\mathcal{C}^{1}}(\mathcal{G}(\Omega_{+}^{f^{n}(\Psi)}),\mathcal{G}(\textrm{Id}_{\Lambda}))\\ &\leq C(\lambda_{+}\mu_{-})^{n}\end{split}$
Hence, we estimate (8.14), using the obvious estimates for $f^{*}$ and the previous estimates:
$\|\tilde{\omega}_{\mid\mathcal{G}(\Omega_{+}^{\Psi})}\|_{\mathcal{C}^{0}}\leq C\eta^{-n}\mu^{2n}_{+}(\lambda_{+}\mu_{-})^{n}=C(\mu^{2}_{+}\lambda_{+}\mu_{-}\eta^{-1})^{n},\quad n\geq 0.$
We conclude that, under the assumptions (8.11), the left hand side of the above vanishes and we obtain that the form $\tilde{\omega}$ vanishes on $\mathcal{G}(\Omega_{+}^{\Psi})$, or, equivalently by Lemma 8.2, that $(\Omega_{+}^{\Psi})^{*}(\omega)=\omega$. As this is true for any section $\Psi$ the map $\Omega_{+}$ satisfies:
$\Omega^{*}_{+}(\omega_{\mid\Lambda})=\omega_{\mid W^{s,\mathrm{loc}}_{\Lambda}}$
An analogous proof works for the map $\Omega_{-}$
Remark 8.3.
The proof in this section, based on the study of graphs, as well as the proof based in iteration given in Section 8.2.7, use only the convergence of $f^{n}(\Psi)$ to $\Lambda$. We use only the most elementary bounds.
One advantage of the use of elementary bounds is that the proofs work for largely arbitrary forms. This allows us to obtain results for more models. See Section D, in particular Section D.2.
When $\omega$ is indeed a symplectic form, $\mathcal{G}(\Omega_{+}^{\Psi})$ is a Lagrangian manifold and this gives extra properties to study the convergence of approximations and perturbation theory.
8.2.6. Proof of Theorem 3.3 (B) for non-closed forms based on vanishing lemmas
In this section we present a version of part (B) of Theorem 3.3 assuming the standing assumptions of Section 3.1 without the hypothesis that the form $\omega$ is closed but assuming that it satisfies (U5). We also assume that the rates (R) satisfy (6.20) and (6.22).
The main tool will be the second item in Lemma 6.12, which claims that $d\omega$ vanishes on the leaves of the stable and unstable manifolds of $\Lambda$.
Theorem 8.4.
In the setup of Theorem 3.3, do not assume $d\omega=0$. Assume that $\omega$ satisfies (U5’). Assume that the hyperbolicity rates of $\Lambda$ satisfy (6.20) and (6.22). Then, we have
$\Omega_{+}^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{s}_{\Lambda}},\ \Omega_{-}^{*}(\omega_{\mid\Lambda})=\omega_{\mid W^{u}_{\Lambda}}.$
Proof.
We do the proof for $\Omega_{+}$. We start as in the proof of part (B) of Theorem 3.3 in Section 8.2.1
Using the same notation, an application of Stokes’ Theorem gives identity (8.5), that we write here:
$\int_{\tilde{\Sigma}}d\omega=\int_{\partial\tilde{\Sigma}}\omega=\int_{\Sigma}\omega-\int_{\Omega_{+}(\Sigma)}\omega+\int_{\Upsilon}\omega,$
where $\tilde{\Sigma}$ is the $3D$-cell defined in (8.3). As we are not assuming $\omega$ is closed, we need an argument to show that the left hand side of this equality vanishes.
We have that
(8.15) $\int_{\tilde{\Sigma}}d\omega=\int_{[0,1]^{3}}d\omega\left(\tilde{\Sigma}(\tilde{t})\right)\left(\partial_{t}\tilde{\Sigma}(\tilde{t}),\partial_{t_{1}}\tilde{\Sigma}(\tilde{t}),\partial_{t_{2}}\tilde{\Sigma}(\tilde{t})\right)dt\,dt_{1}\,dt_{2},$
[…]
As we assume (6.20), we can apply Lemma 6.12 observing that
$\begin{split}\partial_{t}\tilde{\Sigma}(\tilde{t})&\in T_{y}W^{s}_{x},\\ \partial_{t_{1}}\tilde{\Sigma}(\tilde{t}),\partial_{t_{2}}\tilde{\Sigma}(\tilde{t})&\in T_{y}W^{s}_{\Lambda},\end{split}$
Therefore, by (6.21) of Lemma 6.12, the integrand in (8.15) vanishes and we obtain that $\int_{\tilde{\Sigma}}d\omega=0$.
From there, the proof does not need any change from the proof in Section 8.2.
The proof for $\Omega_{-}$ is analogous. ∎
8.2.7. A proof of Theorem 3.3 (B) for non-closed forms based on iteration
In this section we present a proof of a version of part (B) of Theorem 3.3 assuming the standing assumptions of section 3.1, without the hypothesis that the form $\omega$ is closed but assuming that the form $\omega$ satisfies (U5). We also assume that the rates (R) satisfy (6.20) and (6.22).
We will also assume that the manifold $W^{s,\textrm{loc}}_{\Lambda}$ is a $\mathcal{C}^{2}$-manifold, which is stronger than the standing assumption (H2).
We hope that this new proof provides some insights that can be used to develop perturbation theories or to extend the theory to other contexts involving non-closed forms as in [wojtkowski1998conformally].
To prove that the wave maps $\Omega_{\pm}$ preserve the form $\omega$, as we did in section 8.2.5, we take any section $\Psi$ transversal to the foliation (2.16), and prove that $\Omega^{\Psi}_{\pm}$ preserve the form $\omega$.
We use again that $\Omega^{\Psi}_{\pm}$ satisfy (8.13), to relate the projection on $\Psi$ to the projection on $f^{n}(\Psi)$. We will focus in $\Omega_{+}^{\Psi}$. The heuristic idea is that, by Lemma C.1, $f^{n}(\Psi)$ approaches $\Lambda$ for $n\geq N_{+}$ sufficiently large, so that we can approximate $\Omega_{+}^{f^{n}(\Psi)}$ by the identity map.
The errors of symplecticity in the approximation amazingly wash away when put through the (2.24). Let us emphasize that (2.24) is an exact formula for every $n$ and that we do not need to take limits in the formula, only on the estimates obtained by applying it.
For a 2-cell $\mathscr{D}$, we denote $\omega(\mathscr{D})=\int_{\mathscr{D}}\omega$ and $|\mathscr{D}|=\textrm{Area}(\mathscr{D})$ the Riemannian area.
Remark 8.5.
In general, we have $\omega(\mathscr{D})\leq C|\mathscr{D}|$. The converse inequality
(8.16) $|\mathscr{D}|\leq C|\omega(\mathscr{D})|$
is true in bounded neighborhoods when the dimension of $\Lambda$ is $2$, but (8.16) is false when $\Lambda$ has dimension $\geq 4$.
The fact that (8.16) is true when the dimension of $\Lambda$ is $2$, will be developed in Section 8.2.7.1.
We use the coordinate system $(x,y)$ on $W^{s}_{\Lambda}$ described in Section 2.8, and the approximation of $\omega$ near $\Lambda$ given by (6.10). More concretely, by (U5) and (S) – which is implied by (6.20) – we can apply (6.10), obtaining
$|(\Omega^{\Psi}_{+})^{*}\omega_{\mid\Lambda}(x,y)-\omega_{\mid\Psi}(x,y)|\leq M_{\omega}\|y\|.$
Then, for any $2$-cell $\mathscr{D}$ in a neighborhood given by $\|y\|\leq\rho$ we have:
(8.17) $|\omega(\Omega_{+}^{\Psi}(\mathscr{D}))-\omega(\mathscr{D})|=\left|\int_{\mathscr{D}}(\Omega^{\Psi}_{+})^{*}\omega-\omega\right|\leq C\rho|\mathscr{D}|.$
Given a $2$-cell $\mathscr{D}$ contained in $\Psi$, we have, by the conformally symplectic property,
$\omega(f^{n}(\mathscr{D}))=\eta^{n}\omega(\mathscr{D})$
and, by (8.17) applied to $f^{n}(\mathscr{D})$ using (8.8) and (6.10):
(8.18) $\omega(\Omega_{+}^{f^{n}(\Psi)}(f^{n}(\mathscr{D})))=\eta^{n}\omega(\mathscr{D})+e_{n}|f^{n}(\mathscr{D})|$
with
$|e_{n}|\leq CC_{+}\rho\lambda_{+}^{n},\quad|f^{n}(\mathscr{D})|\leq D_{+}^{2}\mu_{+}^{2n}|\mathscr{D}|.$
Finally, using (2.24):
$\begin{split}\omega(\Omega_{+}^{\Psi}(\mathscr{D}))&=\omega(f_{\mid\Lambda}^{-n}\circ\Omega_{+}^{f^{n}(\Psi)}\circ f^{n}(\mathscr{D}))=\eta^{-n}\left(\eta^{n}\omega(\mathscr{D})+e_{n}|f^{n}(\mathscr{D})|\right)\\ &=\omega(\mathscr{D})+\eta^{-n}e_{n}|f^{n}(\mathscr{D})|.\end{split}$
Using the previous bounds we obtain:
$\eta^{-n}e_{n}|f^{n}(\mathscr{D})|\leq D_{+}^{2}C_{+}\rho(\lambda_{+}\mu_{+}^{2}\eta^{-1})^{n}|\mathscr{D}|.$
Under the assumption (6.20), by taking the limit as $n\to\infty$ we conclude that $\omega(\Omega^{\Psi}_{+}(\mathscr{D}))=\omega(\mathscr{D})$ for all 2-cells $\mathscr{D}$ in $\Psi$.
Remark 8.6.
Note that the only ingredients entering in the proof are the equivariance relation (2.24), the infinitesimal vanishing lemma (Lemma 6.5), and the fact that the stable manifold is tangent to the stable bundle. None of those use the fact that the form $\omega$ is closed nor that it is non-degenerate.
8.2.7.1 The case when $\Psi$ is $2$-dimensional: Proof of part (B) of Theorem 3.3 without assumption (S)
When $\Psi$ is $2$-dimensional, if we assume (8.16) with a $C$ uniform on the whole manifold, we can obtain a stronger result.
Suppose that $\omega(\mathscr{D})\geq 0$ (otherwise change $\mathscr{D}$ into the cell with opposite orientation). Then our assumption (8.16) can be written
(8.19) $\forall\mathscr{D}\textrm{ $2$-cell},\,|\mathscr{D}|\leq C\omega(\mathscr{D}).$
By the conformal symplectic property, we have:
$\omega(f^{n}(\mathscr{D}))=\eta^{n}\omega(\mathscr{D}),$
and, by (8.18) and (8.19),
$\begin{split}\omega(\Omega_{+}^{f^{n}(\Psi)}\circ f^{n}(\mathscr{D}))&\leq\eta^{n}\omega(\mathscr{D})+Ce_{n}\omega(f^{n}(\mathscr{D}))\\ &=\eta^{n}(1+Ce_{n})\omega(\mathscr{D})\end{split}$
with $|e_{n}|\leq CC_{+}\rho\lambda_{+}^{n}$.
Finally:
$\begin{split}\omega(\Omega_{+}^{\Psi}(\mathscr{D}))&=\omega(f^{-n}\circ\Omega_{+}^{f^{n}(\Psi)}\circ f^{n}(\mathscr{D}))\\ &\leq\eta^{-n}\left(\eta^{n}(1+e_{n})\omega(\mathscr{D})\right)=(1+e_{n})\omega(\mathscr{D}).\end{split}$
Taking the limit as $n\to\infty$ we get the result.
So, we proved the desired result with the same rate assumptions (S) (but of course we need (8.19) with uniform bounds, which are guaranteed by (U5), as well that the manifold $\Upsilon$ is 22-dimensional).
8.2.7.2 Systematic construction of approximations to scattering map
The method developed above shows that we can reconstruct the scattering map through approximations that get washed away by (2.24). Using the procedure above we can pass some properties of the approximations of $\Omega_{+}^{f^{n}(\Gamma)}$ from the approximations to $\Omega_{+}^{\Gamma}$.
Even if Lemma 8.7 is not used in this paper, we could use it to control perturbations (some version of this was used in [GideaLM21]) or to generate numerical approximations.
Lemma 8.7.
Assume that $\Upsilon^{n}:f^{n}(\Gamma)\rightarrow\Lambda,n>0$ satisfies
[…]
Then,
$\lim_{n\to\infty}\|\Omega_{+}^{\Gamma}-f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}\|_{\mathcal{C}^{0}}=0$
Proof.
The proof of Lemma 8.7 is immediate using formula (2.24) and the mean value theorem. ∎
Lemma 8.8.
Assume that $\Upsilon^{n}:f^{n}(\Gamma)\rightarrow\Lambda,n>0$ is $\mathcal{C}^{1}$ and satisfies
[…]
Then,
$\lim_{n\to\infty}\|\Omega_{+}^{\Gamma}-f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}\|_{\mathcal{C}^{1}}=0$
Proof.
Using the estimates on the rates of growth of higher derivatives from [DLS08, Proposition 15] – recall that in this paper we are assuming condition (N) – and condition (U3) we obtain:
$\|D^{2}f_{\mid\Lambda}^{-n}\|\leq C_{1}\mu_{-}^{2n}n^{2}$
and, therefore,
$\begin{split}\|f_{\mid\Lambda}^{-n}\circ\Omega_{+}^{f^{n}\Gamma}\circ f^{n}-f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}\|_{\mathcal{C}^{1}}&\leq C_{1}\mu_{-}^{2n}n^{2}\|\Omega_{+}^{f^{n}\Gamma}\circ f^{n}-\Upsilon^{n}\circ f^{n}\|_{\mathcal{C}^{1}}\\ \leq&C\mu_{-}^{2n}n^{2}\|\Omega_{+}^{f^{n}\Gamma}-\Upsilon^{n}\|_{\mathcal{C}^{1}}C_{2}\mu_{+}^{n}\end{split}$
Now, we prove that if the approximated map $\Upsilon^{n}$ is approximately symplectic in the weak sense this implies that the map $\Omega_{+}^{\Gamma}$ is symplectic. The following norm is natural
$(\Upsilon^{n})^{*}\omega-\omega]]\equiv\sup_{\mathcal{A}}\frac{|\int_{\Upsilon^{n}(\mathcal{A})}\omega-\int_{\mathcal{A}}\omega|}{|\mathcal{A}|}$
where the supremum is taken over all $\mathcal{A}$, $\mathcal{C}^{1}$ 2-cells in $\Gamma$, and $|\cdot|$ is the Riemannian area.
Lemma 8.9.
Assume that we are in the conditions of Lemma 8.8
Assume furthermore:
(8.20) $\begin{split}&\lim_{n\to\infty}(\eta^{-1}\mu_{+}^{2})^{n}[[(\Upsilon^{n})^{*}\omega-\omega]]=0\end{split}$
Then $\Omega_{+}^{\Gamma}$ is symplectic.
Proof.
Let $\mathcal{A}$ be a 2-cell in $\Gamma$. We have $\omega(f(\mathcal{A}))=\eta\omega(\mathcal{A})$.
Denote $\omega(\mathcal{A})=a$. We have:
  • $\omega(f^{n}(\mathcal{A}))=\eta^{n}a$,
  • By hypothesis 8.20, $\omega(\Upsilon^{n}\circ f^{n}(\mathcal{A}))=\eta^{n}a+\varepsilon_{n}|f^{n}(\mathcal{A})|$ with $\varepsilon_{n}\rightarrow 0$.
  • Then, $\omega(f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}(\mathcal{A}))=\eta^{-n}(\eta^{n}a+\varepsilon_{n}|f^{n}(\mathcal{A})|)=a+\bar{\varepsilon}_{n}$ with $|\bar{\varepsilon}_{n}|\leq C\eta^{-n}\mu_{+}^{2n}\varepsilon_{n}|\mathcal{A}|$. By the assumption $(\eta^{-1}\mu_{+}^{2})^{n}\varepsilon_{n}\to 0$, we obtain that $\omega(f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}(\mathcal{A}))\to a$ as $n\to\infty$.
  • Since by Lemma 8.8 we have $\|\Omega_{+}^{\Gamma}-f_{\mid\Lambda}^{-n}\circ\Upsilon^{n}\circ f^{n}\|_{\mathcal{C}^{1}}\to 0$ this implies that $\omega(\Omega_{+}^{\Gamma}(A))=a=\omega(A)$, which is the integral version of $\Omega_{+}^{\Gamma}$ being symplectic.
8.3. Proof of part (C) of Theorem 3.3 on the exact symplecticity of the scattering map
We give two proofs that the scattering map is exact symplectic (even if the map $f$ is not). The first proof is based on Stokes theorem, and the second one on Cartan’s magic formula.
8.3.1. Proof of part (C) of Theorem 3.3 based on Stokes’ theorem
To prove that the scattering map is exact symplectic, we prove this property for $\Omega_{+}$ (a similar argument applies to $\Omega_{-}$).
We perform a construction similar to that in the proof of part (B) of Theorem 3.3 given in section 8.2.1.
Let $\sigma\subset W^{s,\mathrm{loc}}_{\Lambda}$ be a 1D-cell parameterized by
$\begin{split}\sigma:[0,1]&\to W^{s,\mathrm{loc}}_{\Lambda}\\ u&\mapsto\sigma(u).\end{split}$
We complete $\sigma$ to a 2D-cell $\tilde{\sigma}$ contained in $W^{s,\mathrm{loc}}_{\Omega_{+}(\sigma)}$ by
$\tilde{\sigma}(t,u)=\gamma(t;\sigma(u),\Omega_{+}(\sigma(u))),\quad(t,u)\in[0,1]\times[0,1],$
where
$\tilde{\sigma}(0,u)=\sigma(u),\ \tilde{\sigma}(1,u)=\Omega_{+}(\sigma(u))$
and the path $\gamma$ is defined as in (8.4). See Fig. 5.
Figure 5. A 1-cell $\sigma$ and its completion to a 2D-cell $\tilde{\sigma}$.
We note that, $\tilde{\sigma}$ is contained in $W^{s,\mathrm{loc}}_{\Omega_{+}(\sigma)}$ and, by Proposition 6.9, we know that $\omega$ vanishes on $W^{s,\mathrm{loc}}_{\Omega_{+}^{\Gamma}(\sigma)}$, so
$\int_{\tilde{\sigma}}\omega=0.$
Then, using Stokes’ Theorem, we obtain
(8.21) $0=\int_{\tilde{\sigma}}\omega=\int_{\tilde{\sigma}}d\alpha=\int_{\partial\tilde{\sigma}}\alpha=\int_{\sigma}\alpha-\int_{\Omega_{+}(\sigma)}\alpha+\int_{\gamma_{1}}\alpha-\int_{\gamma_{0}}\alpha$
where
$\begin{split}\gamma_{1}(t)=\tilde{\sigma}(t,1)=\gamma(t;\sigma(1),\Omega_{+}(\sigma(1)),\\ \gamma_{0}(t)=\tilde{\sigma}(t,0)=\gamma(t;\sigma(0),\Omega_{+}(\sigma(0)),\ t\in[0,1].\end{split}$
Now, we define the following function on $W^{s,\mathrm{loc}}_{\Lambda}$:
(8.22) $P^{+}(x)=\int_{\gamma(\cdot;x,\Omega_{+}(x))}\alpha,\textrm{ for }x\in W^{s,\mathrm{loc}}_{\Lambda}.$
Note that $P^{+}=0$ on $\Lambda$. The function $P^{+}$, clearly depends on $\alpha$, but when $\alpha$ is fixed, we will not include it in the notation. See Remark 8.12 for the effects of changing $\alpha$.
Lemma 8.10.
The integral defining $P^{+}(x)$ in (8.22) does not depend on the path $\gamma(\cdot;x,\Omega_{+}(x))$ in $W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$ chosen to connect $x$ to $\Omega_{+}(x)$.
Lemma 8.10 shows that $P^{+}$ is indeed well defined as a function on $W^{s,\mathrm{loc}}_{\Lambda}$.
Proof.
Take another path $\tilde{\gamma}(\cdot;x,\Omega_{+}(x))$ contained in $W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$, and call
$\tilde{P}(x)=\int_{\tilde{\gamma}(\cdot;x,\Omega_{+}(x))}\alpha.$
We know that $d\alpha_{\mid W^{s}_{\Omega_{+}(x)}}=0$. Since $W^{s}_{\Omega_{+}(x)}$ is simply connected (by Theorem A.1(IV)-(ii)) we have that $\gamma\cup\tilde{\gamma}$ bounds a 2D-cell $\mathcal{B}$ in $W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$. Applying Stokes theorem we get:
$\begin{split}0=\int_{\mathcal{B}}d\alpha=\int_{\gamma\cup\tilde{\gamma}}\alpha=P^{+}(x)-\tilde{P}(x).\end{split}$
Using the function $P^{+}$ in (8.22), by (8.21) we obtain:
$\int_{\sigma}\Omega_{+}^{*}(\alpha)=\int_{\sigma}\alpha+P^{+}(\sigma(1))-P^{+}(\sigma(0))=\int_{\sigma}\alpha+\int_{\sigma}dP^{+}.$
As this is true for any $1$-cell $\sigma$, we therefore have proved that
(8.23) $\Omega_{+}^{*}\left(\alpha_{\mid\Lambda}\right)=\alpha_{\mid W^{s,\textrm{loc}}_{\Lambda}}+dP^{+},$
where $P^{+}$ given by the formula (8.22) is the primitive function of $\Omega_{+}$ with respect to $\alpha$ which satisfies $P^{+}_{\mid\Lambda}=0$.
More important for our purposes, we can restrict (8.23) to the homoclinic channel $\Gamma$ and obtain:
$\left(\Omega_{+}^{\Gamma}\right)^{*}\left(\alpha_{\mid\Lambda}\right)=\alpha_{\mid\Gamma}+dP^{+}_{\mid\Gamma}$,
A similar argument yields $\Omega_{-}$ is exact, with the primitive function $P^{-}$, given by an integral formula similar to (8.22).
Using the elementary calculation from Lemma 2.13, it follows that the scattering map $S=\Omega^{\Gamma}_{+}\circ(\Omega^{\Gamma}_{-})^{-1}$ is exact. The derivation below (which uses (2.5) and (2.7)) gives a formula for the primitive function for $S$:
$\begin{split}P^{S}&=P^{\Omega_{+}^{\Gamma}\circ(\Omega_{-}^{\Gamma})^{-1}}=P^{({\Omega_{-}^{\Gamma}})^{-1}}+P^{{\Omega_{+}^{\Gamma}}}\circ({\Omega_{-}^{\Gamma}})^{-1}\\ &=-P^{\Omega_{-}^{\Gamma}}\circ(\Omega_{-}^{\Gamma})^{-1}+P^{\Omega_{+}^{\Gamma}}\circ({\Omega_{-}^{\Gamma}})^{-1},\end{split}$
so, using the notation from above,
$P^{S}=(P^{+}-P^{-})\circ(\Omega_{-}^{\Gamma})^{-1}.$
Remark 8.11.
Assume that the symplectic form is exact $\omega=d\alpha$. Since $\omega$ vanishes on $W^{u,s}_{x}$, and $W^{u,s}_{x}$ is simply connected (see Theorem A.1), by applying the Poincaré Lemma we obtain that the restrictions of $\alpha$ to the stable/unstable fibers $\alpha_{\mid W^{u,s}_{x}}$ are exact. Then (8.22) shows that $P^{\pm}_{\mid W^{s,u}_{x}}$ is a primitive of $\alpha_{\mid W^{s,u}_{x}}$.
Remark 8.12.
The integral formulas (8.22) are for a fixed action function $\alpha$.
For the primitive $\tilde{\alpha}$ related to the $\alpha$ by a gauge transformation, that is $\tilde{\alpha}=\alpha+dG$ with $G$ a real valued function on $M$, on $W^{s}_{\Lambda}$ we have:
(8.24) $P^{+}_{\tilde{\alpha}}=P^{+}_{\alpha}+G\circ\Omega_{+}-G.$
A similar formula holds for $P^{-}$ on $W^{u}_{\Lambda}$.
The generating function $P^{S}_{\tilde{\alpha}}$ satisfies the following on the domain $H^{-}\subset\Lambda$ of $S$:
$P^{S}_{\tilde{\alpha}}=P^{S}_{\alpha}+G\circ S-G.$
8.3.2. Proof of part (C) of Theorem 3.3 based on Cartan’s magic formula
With the same notation as in Section 8.2.4 we compute:
$\frac{d}{dt}\phi_{t}^{*}\alpha=\phi^{*}_{t}\left[i_{V}d\alpha+d(i_{V}\alpha)\right]=\phi_{t}^{*}(di_{V}\alpha)=d(\phi^{*}_{t}(i_{V}\alpha)),$
where we have used $i_{V}(d\alpha)=i_{V}(\omega)=0$. The last equality is because of the Vanishing Lemma 6.8.
Therefore we have
$((\Omega_{+}^{\Gamma})^{-1})^{*}\alpha-\alpha=\phi_{1}^{*}\alpha-\phi^{*}_{0}\alpha=d\int_{0}^{1}\phi^{*}_{t}(i_{V}\alpha)\,dt,$
showing that $(\Omega_{+}^{\Gamma})^{-1}$ is exact. The fact that $\Omega_{+}^{\Gamma}$ is exact follows from applying (2.7) to the map $(\Omega_{+}^{\Gamma})^{-1}$ for $\eta=1$. We also obtain the direct formula:
$(\Omega_{+}^{\Gamma})^{*}\alpha-\alpha=-d\left[\int_{0}^{1}\left(\phi_{t}\circ(\Omega_{+}^{\Gamma}))^{*}(i_{V}\alpha)\,dt\right)\right].$
Remark 8.13.
Note that this proof also gives that, if we have a presymplectic manifold and consider the foliation by the kernel, the holonomy maps between two transversals (which are symplectic manifolds since the kernel is excluded) are symplectic maps.
8.4. Proof of Proposition 6.9
The first item of Proposition 6.9 is a direct consequence of part (B) of Theorem 3.3. For $\Omega_{+\mid W^{s}_{N}}:W^{s}_{N}\to N$ with $N\subset\Lambda$, as $(\Omega_{+})^{*}(\omega_{|\Lambda})=\omega_{|W^{s,\mathrm{loc}}_{\Lambda}}$, using that $\omega_{\mid N}=0$, we get:
$0=(\Omega_{+\mid W^{s}_{N}})^{*}\omega_{\mid N}=\omega_{\mid W^{s}_{N}}.$
To prove the second item, let any $y\in W^{s,\mathrm{loc}}_{\Lambda}$. We have that $y\in W^{s,\mathrm{loc}}_{x}$ for $x=\Omega_{+}(y)\in\Lambda$. Using that $\Lambda$ is symplectic (Part (A) of Theorem 3.1) we can construct as isotropic manifold $N\subset\Lambda$ of dimension $\frac{d_{c}}{2}$ (using, for example Darboux theorem) with $x\in N$. Using Proposition 6.9 part (i), we obtain that $W^{s}_{N}$ is isotropic. We note that $y\in W^{s}_{N}\subset W^{s}_{\Lambda}$ and that
$\text{dim}(W^{s}_{N})=\frac{d_{c}}{2}+d_{s}=\frac{1}{2}(d_{c}+d_{s}+d_{u}).$
Therefore $W^{s}_{N}$ is a Lagrangian submanifold of $M$.
Therefore, $T_{y}W^{s}_{N}$ is a Lagrangian subspace of $T_{y}M$. Since $T_{y}W^{s}_{N}\subset T_{y}W^{s}_{\Lambda}$, we conclude that $T_{y}W^{s}_{\Lambda}$ is coisotropic, and since $y$ was an arbitrary point, the manifold $T_{y}W^{s}_{\Lambda}$ is coisotropic but not Lagrangian.
9. Formulas for the primitive functions of wave maps and scattering map when $f$ is exact
When $f$ is exact conformally symplectic, in this section we obtain formulas for the primitive functions of the wave maps and the scattering map in terms of the primitive function of $f$. The variational formulation for conformally symplectic systems is given in (9.21). The formulas (9.6) (9.7) and (9.9) provide a link with the calculus of variations for conformally symplectic systems. In the symplectic case, [Angenent93, Lomeli97, AmbrosettiB98] develop variational descriptions of heteroclinic connections. Such formulas have been used in numerical calculations of orbits homoclinic (or heteroclinic) to periodic orbits [Tabacman95] of twist maps.
Fixing an action form $\alpha$, we rearrange the definition of an exact conformally symplectic map (2.4) as
(9.1) $\alpha=\eta^{-1}f^{*}\alpha-d\eta^{-1}P^{f}_{\alpha}.$
Applying formula (9.1) repeatedly, we obtain for any $N\in\mathbb{N}$
(9.2) $\alpha=\eta^{-N}(f^{*})^{N}\alpha-d\left(\sum_{j=0}^{N-1}\eta^{-j-1}P^{f}_{\alpha}\circ f^{j}\right).$
Similarly, rearranging (2.4) as
$\alpha=\eta(f^{*})^{-1}\alpha+dP^{f}_{\alpha}\circ f^{-1}$
and iterating we have:
(9.3) $\alpha=\eta^{N}(f^{*})^{-N}\alpha+d\left(\sum_{j=1}^{N}\eta^{j-1}P^{f}_{\alpha}\circ f^{-j}\right).$
Integrating (9.2) and (9.3) over a path $\sigma$ and remembering that the integral of a differential over a path is just the difference of the values at the ends, we obtain for any path $\sigma$:
(9.4) $\int_{\sigma}\alpha=\eta^{-N}\int_{f^{N}(\sigma)}\alpha-\sum_{j=0}^{N-1}\eta^{-j-1}\left(P^{f}_{\alpha}\circ f^{j}(\sigma(1))-P^{f}_{\alpha}\circ f^{j}(\sigma(0))\right),$
(9.5) $\int_{\sigma}\alpha=\eta^{N}\int_{f^{-N}(\sigma)}\alpha+\sum_{j=1}^{N}\eta^{j-1}\left(P^{f}_{\alpha}\circ f^{-j}(\sigma(1))-P^{f}_{\alpha}\circ f^{-j}(\sigma(0))\right).$
Given $x\in W^{s,\mathrm{loc}}_{\Lambda}$, when $\sigma$ is chosen to be the path $\gamma_{+}(\cdot;x,\Omega_{+}(x))\subset W^{s,\mathrm{loc}}_{\Omega_{+}(x)}$ given in (8.4), and denoting by
$\gamma^{N}_{+}:=f^{N}(\gamma_{+}(\cdot;x,\Omega_{+}(x))=\gamma_{+}(\cdot;f^{N}(x),\Omega_{+}(f^{N}(x))),$
using the formula (8.22) for the primitive function $P^{+}_{\alpha}$ (and, similarly, in the analogous formula for $P^{-}_{\alpha}$), we obtain:
Lemma 9.1.
The primitive functions $P^{\pm}_{\alpha}$ of $\Omega_{\pm}$ for the action form $\alpha$ are given by:
(9.6) $\begin{split}P^{+}_{\alpha}(x)=&\eta^{-N}\int_{\gamma^{N}_{+}}\alpha+\sum_{j=0}^{N-1}\eta^{-j-1}[P^{f}_{\alpha}\circ f^{j}(\Omega_{+}^{\Gamma}(x))-P^{f}_{\alpha}\circ f^{j}(x)],\end{split}$
(9.7) $\begin{split}P^{-}_{\alpha}(x)=&\eta^{N}\int_{\gamma^{N}_{-}}\alpha+\sum_{j=1}^{N}\eta^{j-1}[P^{f}_{\alpha}\circ f^{-j}(\Omega_{-}^{\Gamma}(x))-P^{f}_{\alpha}\circ f^{-j}(x)].\end{split}$
The primitive function $P^{S}_{\alpha}$ of the scattering map is given by (see (2.5) and (2.7)):
(9.8) $\begin{split}P^{S}_{\alpha}(x)&=\left[\eta^{-N}\int_{\gamma^{N}_{+}}\alpha+\sum_{j=0}^{N-1}\eta^{-j-1}[P_{\alpha}^{f}\circ f^{j}(\Omega^{\Gamma}_{+})-P_{\alpha}^{f}\circ f^{j}]\right]\circ(\Omega^{\Gamma}_{-})^{-1}(x)\\ &-\left[\eta^{N}\int_{\gamma^{N}_{-}}\alpha+\sum_{j=1}^{N}\eta^{j-1}[P_{\alpha}^{f}\circ f^{-j}(\Omega^{\Gamma}_{-})-P_{\alpha}^{f}\circ f^{-j}]\right]\circ(\Omega^{\Gamma}_{-})^{-1}(x)\\ =&\eta^{-N}\int_{\gamma^{N}_{+}}\alpha\circ(\Omega^{\Gamma}_{-})^{-1}(x)+\sum_{j=0}^{N-1}\eta^{-j-1}[P_{\alpha}^{f}\circ f^{j}\circ S(x)-P_{\alpha}^{f}\circ f^{j}\circ(\Omega^{\Gamma}_{-})^{-1}(x)]\\ &-\eta^{N}\int_{\gamma^{N}_{-}}\alpha\circ(\Omega^{\Gamma}_{-})^{-1}(x)-\sum_{j=1}^{N}\eta^{j-1}[P_{\alpha}^{f}\circ f^{-j}(x)-P_{\alpha}^{f}\circ f^{-j}\circ(\Omega^{\Gamma}_{-})^{-1}(x)].\end{split}$
Remark 9.2.
The function $P^{+}_{\alpha}$ in (9.6) is well defined for all $x\in W^{s}_{\Lambda}$. Therefore, the series on the right-hand side of (9.6) is convergent if and only if the sequence $\eta^{-N}\int_{\gamma^{N}_{+}}\alpha$ converges to zero. Analogously for $P^{-}_{\alpha}$ in (9.7).
The convergence of the series in (9.6) and (9.7) is very easy if the orbits of $f$ in $\Lambda$ are bounded. In such a case we have uniform bounds on $\alpha$ and the paths $\gamma^{N}_{\pm}$ have lengths bounded by $\lambda_{\pm}^{|N|}$ (when $N\rightarrow\pm\infty$).
However, even if $\omega$ is uniformly bounded, $\alpha$ may be unbounded (see Section 9.4 for lower bounds for all forms). If $f^{N}(x)$ escapes to infinity, it could happen that $\|\alpha_{f^{N}(x)}\|$ grows so fast that it overtakes the decrease of the length of $\gamma^{N}_{\pm}$ (and the powers of $\eta$).
In Section 9.1, we show that there is a gauge in which the formulas (9.6) and (9.7) converge very fast. Indeed, with the construction of Section 9.1, the formulas become finite sums.
In Section 9.2 we will show that, if there are orbits that escape to infinity, there is always a gauge that makes (9.6) and (9.7) divergent.
9.1. Construction of gauges yielding convergence of the series for the primitives of the wave maps and the scattering map
Lemma 9.3.
Given an action form $\alpha$ for $\omega$, there exists an action form $\tilde{\alpha}=\alpha+dG$, for some smooth function $G:M\to\mathbb{R}$, such that
(9.9) $\begin{split}P^{S}_{\tilde{\alpha}}=&\sum_{j=0}^{\infty}\eta^{-j-1}[P^{f}_{\tilde{\alpha}}\circ f^{j}\circ S-P^{f}_{\tilde{\alpha}}\circ f^{j}\circ(\Omega_{-})^{-1}]\\ &-\sum_{j=1}^{\infty}\eta^{j-1}[P^{f}_{\tilde{\alpha}}\circ f^{-j}-P^{f}_{\tilde{\alpha}}\circ f^{-j}\circ(\Omega_{-})^{-1}].\end{split}$
Proof.
Let $G^{+}=P^{+}_{\alpha}$ given in (8.22) ,defined on $W^{s,\textrm{loc}}_{\Lambda}$, and note that $G^{+}=0$ on $\Lambda$, so, by (8.24) we have $P^{+}_{\alpha+dG^{+}}=P^{+}_{\alpha}-G^{+}=0$ on $W^{s,\textrm{loc}}_{\Lambda}$.
Note also that (9.6)can be written as:
$P^{+}_{\alpha}(x)=P^{+}_{\alpha}(F^{N}(x))+\sum_{j=0}^{N-1}\eta^{-j-1}[P^{f}_{\alpha}\circ f^{j}(\Omega_{+}^{\Gamma}(x))-P^{f}_{\alpha}\circ f^{j}(x)].$
Applying this formula for $\alpha+dG^{+}$ we obtain that the sum is zero.
Similarly, for $G^{-}=P^{-}_{\alpha}$ on $W^{u,\textrm{loc}}_{\Lambda}$, we have $P^{-}_{\alpha+dG^{-}}=P^{-}_{\alpha}-G^{-}=0$ on $W^{u,\textrm{loc}}_{\Lambda}$.
So, we have accomplished our goal of making the series convergent (in fact zero) using gauge functions $G^{\pm}$ defined on $W^{s,\textrm{loc}}_{\Lambda}\cup W^{s,\textrm{loc}}_{\Lambda}$. What remains is to show that this partially defined function can be extended to the whole $M$ so that it is a well defined gauge function and makes the series convergent (but not zero). We now give the details, which are fairly standard.
Let $\mathcal{O}_{\rho}$ be the uniform neighborhood of $\Lambda$ defined in (2.13). Choose $\rho^{\prime}>\rho$ such that $\mathcal{O}_{\rho^{\prime}}$ is disjoint from the homoclinic channel $\Gamma$.
Now we construct a function $G^{ext}:M\to\mathbb{R}$ which agrees with $G^{+}$ on $W^{s,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$ and with $G^{-}$ on $W^{u,\textrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$. We now give the details of this extension by using partitions of unity, and this finishes the proof of Lemma 9.3.
We can cover $\mathcal{O}_{\rho}$ by a countable collection of uniform balls $\mathcal{B}_{i}$, $i=1,\ldots,\infty$, such that:
  • (C1)
    $\mathcal{O}_{\rho}\subseteq\bigcup_{i}\mathcal{B}_{i}\subseteq\mathcal{O}_{\rho^{\prime}}$;
  • (C2)
    Each point $x\in\mathcal{O}_{\rho}$ is contained in only finitely many balls $\mathcal{B}_{i_{1}},\ldots,\mathcal{B}_{i_{L+1}}$ (here $L$ is the covering dimension of the manifold $M$, which equals the dimension of the manifold);
  • (C3)
    On each $\mathcal{B}_{i}$ we have a local trivialization of $E^{s}\oplus E^{u}$, that is,
    (9.10) $(E^{s}\oplus E^{u})_{\Lambda\cap\mathcal{B}_{i}}\simeq(\Lambda\cap\mathcal{B}_{i})\oplus E^{s}\oplus E^{u}.$
By (C3), on each open set $\mathcal{B}_{i}$ we can choose a system of coordinates $(c_{i},s_{i},u_{i})$ such that
(9.11) $\begin{split}\Lambda\cap\mathcal{B}_{i}=&\{(c_{i},s_{i},u_{i})\,|\,s_{i}=u_{i}=0\},\\ W^{s}_{\Lambda}\cap\mathcal{B}_{i}=&\{(c_{i},s_{i},u_{i})\,|\,u_{i}=0\},\\ W^{u}_{\Lambda}\cap\mathcal{B}_{i}=&\{(c_{i},s_{i},u_{i})\,|\,s_{i}=0\}.\end{split}$
We note that the systems of coordinates associated to two open sets $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ that have non-empty intersection do not have to agree with one another.
There exists a smooth partition of unity $\{\Psi_{i}\}$ subordinate to $\{\mathcal{B}_{i}\}$, with $\Psi_{i}:M\to\mathbb{R}$, such that:
  • For each $x\in\mathcal{O}_{\rho}$ and each $i$ we have $0\leq\Psi_{i}(x)\leq 1$;
  • For each $x\in\mathcal{O}_{\rho}$ there is an open neighborhood of $x$ such that all but finitely many $\Psi_{i}$’s are 0 on that open neighborhood of $x$;
  • For each $x\in\mathcal{O}_{\rho}$ we have $\Sigma_{i}\Psi_{i}(x)=1$ (by the previous condition this sum is finite on an open neighborhood of $x$);
  • For each $i$, $\textrm{supp}(\Psi_{i})\subseteq\mathcal{B}_{i}$.
This is a direct consequence of [Spivak, Theorem 3-11].
Now define $G^{0}_{i}:\mathcal{B}_{i}\to\mathbb{R}$ by
$G^{0}_{i}(c_{i},s_{i},u_{i})=G^{+}(c_{i},s_{i})+G^{-}(c_{i},u_{i}),$
where the underlying coordinate system corresponds to $\mathcal{B}_{i}$. Since $G^{+}_{\mid\Lambda}=G^{-}_{\mid\Lambda}=0$, we then have $G^{0}_{i}=0$ on $\Lambda\cap\mathcal{B}_{i}$. The function $G^{0}_{i}$ is a local extension of both $G^{+}$ and $G^{-}$, and depends on the underlying coordinate system.
Since $\textrm{supp}(\Psi_{i})\subseteq\mathcal{B}_{i}$, we define a global extension $G^{ext}_{i}:M\to\mathbb{R}$ of $G^{0}_{i}$ given by:
$G^{ext}_{i}(x)=\left\{\begin{array}[]{ll}G^{0}_{i}(x)\Psi_{i}(x)&\hbox{ on $\textrm{supp}(\Psi_{i})\subseteq\mathcal{B}_{i}$},\\ 0&\hbox{ on $M\setminus\textrm{supp}(\Psi_{i})$}.\end{array}\right$.
Then, we combine the functions $G^{ext}_{i}$ into a single global extension $G^{ext}:M\to\mathbb{R}$ by
(9.12) $\begin{split}G^{ext}(x)=&\sum_{i}G^{ext}_{i}(x).\end{split}$
Although a point $x$ may be covered by finitely many open sets $\mathcal{B}_{i_{1}},\ldots,B_{i_{L+1}}$, in which case the point $x$ has different coordinate representations $x=x(c_{i_{j}},s_{i_{j}},u_{i_{j}})$, $j=1,\ldots,L+1$, we have that
$G^{ext}(x)=\sum_{j=1}^{L+1}(G^{+}(c_{i_{j}},s_{i_{j}})+G^{-}(c_{i_{j}},u_{i_{j}}))\Psi_{i_{j}}(c_{i_{j}},s_{i_{j}},u_{i_{j}})$
is independent of the local system of coordinates.
By the uniformity assumption (U1), for some $C>0$ we also have
$\|G^{ext}\|_{\mathcal{C}^{r}(M)}\leq C\left[\|G^{+}\|_{\mathcal{C}^{r}(W^{s,\mathrm{loc}}_{\mid\Lambda})}+\|G^{-}\|_{\mathcal{C}^{r}(W^{u,\mathrm{loc}}_{\mid\Lambda})}\right].$
Note that
(9.13) $G^{ext}=\left\{\begin{array}[]{ll}G^{+}&\hbox{ on $W^{s,\mathrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$},\\ G^{-}&\hbox{ on $W^{u,\mathrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$}.\end{array}\right$.
and
(9.14) $\begin{split}G^{ext}=&0\textrm{ on }M\setminus\bigcup_{i}\mathcal{B}_{i}\implies G^{ext}=0\textrm{ on }M\setminus\mathcal{O}_{\rho^{\prime}}.\end{split}$
We now show that for the modified action form $\tilde{\alpha}=\alpha+dG^{ext}$ the series in (9.6) is convergent (in fact, it becomes a finite sum for every point).
Let $x$ be a point in $W^{s}_{\Lambda}$. Since $d(f^{n}(x),\Lambda)\to 0$ as $n\to\infty$, there exists $N$ depending on $x$ such that $f^{N}(x)\in\mathcal{O}_{\rho}$, and so $\gamma^{N}_{+}\subseteq\mathcal{O}_{\rho}$. Then
(9.15) $\begin{split}\int_{\gamma^{N}_{+}}\alpha+dG^{ext}=&\int_{\gamma^{N}_{+}}\alpha+dG^{+}\\ =&P^{+}_{\alpha}(f^{N}(x))-G^{+}(f^{N}(x))=0,\end{split}$
and therefore, using (9.6) yields
(9.16) $P^{+}_{\tilde{\alpha}}(x)=\sum_{j=0}^{N-1}\eta^{-j-1}[P^{f}_{\tilde{\alpha}}\circ f^{j}\circ\Omega_{+}(x)-P^{f}_{\tilde{\alpha}}\circ f^{j}(x)].$
Thus
(9.17) $P^{+}_{\tilde{\alpha}}=\sum_{j=0}^{\infty}\eta^{-j-1}[P^{f}_{\tilde{\alpha}}\circ f^{j}\circ\Omega_{+}-P^{f}_{\tilde{\alpha}}\circ f^{j}],$
where for each $x$ the above sum is finite.
Similarly
(9.18) $P^{-}_{\tilde{\alpha}}=\sum_{j=1}^{\infty}\eta^{j-1}[P^{f}_{\tilde{\alpha}}\circ f^{-j}\circ\Omega_{-}-P^{f}_{\tilde{\alpha}}\circ f^{-j}].$
By substituting (9.17) and (9.18) in (9.9), we obtain the desired conclusion. ∎
9.2. Construction of gauges yielding divergence of the series for the primitives of the wave maps when there are escaping orbits
The purpose of this section is to show that, if $f_{\mid\Lambda}$ has orbits that escape to infinity121212the orbit $f^{n}(y)$ escapes to infinity if every compact $K\subset\Lambda$ contains only finitely many points of the orbit, there is always a function $G$ (in fact, plenty of them) such that the series (9.6), which give the primitive $P_{\alpha}^{+}$ of the wave map $\Omega^{+}$, corresponding to the modified $1$-form $\alpha+dG$,
(9.19) $\begin{split}&\sum_{j\geq 0}\eta^{-j-1}[P_{\alpha+dG}\circ f^{j}(\Omega_{+}(x))-P_{\alpha+dG}\circ f^{j}(x)]\end{split}$
is divergent for some $x$.
More concretely, we have the following:
Lemma 9.4.
Assume that for a given $\alpha$ the series (9.19) with $G=0$ is convergent. Assume also that there exists a point $y\in\Lambda$ such that its orbit $y_{n}=f^{n}(y)$ escapes to infinity as $n\to\infty$.
Then, there exists $G:M\to\mathbb{R}$, such that the series (9.19) for $\alpha+dG$ is divergent at the point $y$.
Proof.
If the orbit $y_{n}=f^{n}(y)$ escapes to infinity, then it has a subsequence $f^{k_{n}}(y)$ such that for some $\delta>0$ we have
$d(y_{k_{n}},y_{k_{m}})\geq\delta$
for all $n,m$ with $m\neq n$.
Take $x\in W^{s}_{y}$, and denote $x_{n}=f^{n}(x)$ and $y_{n}=f^{n}(\Omega_{+}(x))=f^{n}(y)$. The points $x_{n},y_{n}$ are the endpoints of a curve $\gamma^{n}_{+}(\cdot)=f^{n}(\gamma_{+}(\cdot;x,\Omega_{+}(x)))$ in $W^{s}_{y_{n}}$. Assume that the sequence $\eta^{-n}\int_{\gamma^{n}_{+}}\alpha$ is convergent, otherwise there is nothing to prove (see Remark 9.2).
For any point $z_{n}=\gamma^{n}_{+}(t)$ on this curve, we know by Theorem A.1 part (II)-(iv) that there exists $\tilde{C}>0$ such that:
$d(z_{k_{n}},y_{k_{n}})\leq\tilde{C}\lambda_{+}^{k_{n}}\to 0,\ n\to\infty.$
Since $d(y_{k_{n}},y_{k_{n^{\prime}}})\geq\delta$ for all $n\geq 0$, it follows that, changing $\delta$ if necessary:
$d_{H}(\gamma^{k_{n}}_{+},\gamma^{k_{n^{\prime}}}_{+})\geq\delta,\textrm{ for }n\in\mathbb{N},\,n\neq n^{\prime}$
where $d_{H}$ is the Hausdorff distance.
Therefore, we can choose small, open neighborhoods $V_{n}$ of the curves $\gamma^{k_{n}}_{+}$ in $M$, such that $V_{k_{n}}\cap V_{k_{n^{\prime}}}=\emptyset$, and construct smooth functions $G_{n}$ with support in $V_{n}$ such that:
$G_{n}(x_{k_{n}})-G_{n}(y_{k_{n}})=(\eta\beta)^{n},\ n\geq 0,$
for some $\beta>1$. We write $G=\sum_{n}G_{n}$ (recall that the supports of $G_{n}$ are disjoint). Using that $x_{k_{n}},y_{k_{n}}$ are in the support of $G_{k_{n}}$ and not in the support of any other, we have
$G(x_{k_{n}})-G(y_{k_{n}})=(\eta\beta)^{k_{n}},\ n\geq 0.$
Then the series
(9.20) $\begin{split}&\eta^{-k_{n}}\int_{\gamma^{k_{n}}_{+}}\left(\alpha+dG\right)\\ &=\eta^{-k_{n}}\int_{\gamma^{k_{n}}_{+}}\alpha+\eta^{-k_{n}}(G(y_{k_{n}})-G(x_{k_{n}}))\\ &=\eta^{-k_{n}}\int_{\gamma^{k_{n}}_{+}}\alpha+\beta^{k_{n}}.\end{split}$
Since $\beta>1$, the sequence is divergent.
Going back to formula (9.6) and using that (9.19) is convergent, we obtain that the series for $P_{\alpha+dG}$ is divergent. ∎
9.3. Variational interpretation of the iterative formulas for primitive functions of scattering map
In this section, we discuss the variational interpretation of (9.8). The material in this section will not be used in this paper. The sole purpose of this section is to point out a possible bridge between variational and geometric approaches to heteroclinic jumps.
Let $T^{*}Q$ be a cotangent bundle of a manifold – the standard example of a symplectic manifold – with the canonical $1$-form defined in coordinates $(p,q)$ as $\alpha_{0}=pdq$  [^13] A geometrically natural definition of $\alpha_{0}$ is standard, see, for example: [AbrahamM78, Proposition 3.2,11 p. 180].. Let $f$ be a mapping on $T^{*}Q$ that is homotopic to the identity, exact conformally symplectic ($f^{*}\alpha-\eta\alpha=dP^{f}_{\alpha}$ for some function $P^{f}_{\alpha}$ on $T^{*}Q$), and satisfies a twist condition,  [^14] The twist condition is clearly non-generic – it is not verified by the identity– but it is verified by the geodesic flow of a compact manifold at time $t>0$ for small enough $t$ [Gole94]. meaning that if $f(p,q)=(p^{\prime},q^{\prime})$ then $(q,q^{\prime})$ gives a system of coordinates on $Q\times Q$. The primitive function $P^{f}_{\alpha}$ has the interpretation of an action and that the orbits of $f$ are critical points of the formal action
[…]
where $x=(x_{n})_{n}=(f^{n}(x))_{n}$. See [Haro00] for the symplectic case $\eta=1$.
Remark 9.5.
The variational principle in (9.21), often called discounted variational principle, appears naturally in finance. Each $x_{n}$ is a transaction at time $n$ and $P^{f}_{\alpha}(x_{n})$ is the cost of the transaction in currency. If there is constant inflation, to evaluate the cost of a strategy, it is natural to add the costs at different times by reducing them to a common time [Bensoussan88]. The conversion of the cost $P^{f}_{\alpha}(x_{n})$ to currency at $n=0$ is $\eta^{-n}P^{f}_{\alpha}(x_{n})$.
The variational principles (9.21) also appear in control theory under the name finite horizon approximation [KaliseKR17].
We can write the function $P^{f}_{\alpha}$ on the manifold in a coordinate patch as $S(q,q^{\prime})$, so that the conformally symplectic property can be written as
$p^{\prime}dq^{\prime}=\eta pdq+dS(q,q^{\prime}).$
This is equivalent to
(9.22) $\begin{split}&p^{\prime}=\partial_{2}S(q,q^{\prime}),\\ &p=-\eta^{-1}\partial_{1}S(q,q^{\prime}).\end{split}$
[…]
Given any real valued functions $\phi_{n}$ we can consider instead of the formal variational principle (9.21), the variational principle
[…]
Hence, taking the variational principle consisting of the primitive of the scattering map and the primitive of the map gives a variational principle for orbits of the map that include homoclinic excursions.
The variational approach to homoclinic orbits has also been used as a numerical tool for symplectic systems [Tabacman95, MacKayMS89]. Using that (9.4), (9.5) can be computed by other methods and considered as boundary terms, it seems that the method could be adapted to conformally symplectic systems.
9.4. Unbounded action forms
It is well know that a $(2n)$-dimensional, connected, Riemannian manifold $M$ that is closed (i.e., boundaryless and compact), cannot have a symplectic form $\omega$ that is exact. Indeed, if $\omega=d\alpha$ then $\omega^{n}=d(\alpha\wedge\omega^{n-1})$, and Stokes’ Theorem implies
(9.24) $\int_{M}\omega^{n}=\int_{M}d(\alpha\wedge\omega^{n-1})=\int_{\partial M}\alpha\wedge\omega^{n-1}=0,$
which contradicts the fact that $\omega^{n}$ is a volume form on $M$. Therefore, if such a manifold has an exact symplectic form, the manifold cannot be closed.
Below we show that if the manifold has an exact symplectic form $\omega=d\alpha$ and satisfies some additional conditions, then $\|\alpha\|$ must be unbounded. Moreover, we can provide some quantitative estimates on the growth of $\|\alpha\|$ along geodesic balls $B_{R}$.
Remark 9.6.
A symplectic manifold with an exact symplectic form can be with boundary or non-compact, but does not need to be unbounded. Of course, such phenomenon can only happen for manifolds which do not satisfy (U1) or (U1’)
For instance, $M$ can be a bounded cylinder (with or without boundary)
$M=\{(I,\theta)\,|\,I\in B^{n}_{R},\,\theta\in\mathbb{T}^{n}\},$
where $B^{n}_{R}$ is a ball in $\mathbb{R}^{n}$ (open or closed). The standard symplectic form $\omega=dI\wedge d\theta$ is exact with action form $\alpha=Id\theta$. Moreover, $\alpha=Id\theta$ is bounded on $M$.
We recall that if the Riemannian metric is complete, then every geodesic can be extended to a all times. Fixing a point $o\in M$, the geodesic ball $B_{R}$ is the set of points $x\in M$ for which $d(o,x)\leq R$. The distance $d(0,x)$ is a smooth function in $x$ except for the cut locus [^16] The cut locus consists of points that are conjugate to $o$ and points that have multiple minimal geodesics connecting them to $o$. of $o$.
Denote by $\textrm{Vol}_{d}$ the Riemannian volume on a $d$ dimensional manifold.
The form $\omega$ is said to be uniformly nondegenerate, if there exists a constant $C>0$ such that
(9.25) $|\omega^{n}|\geq C\cdot d{\textrm{Vol}_{2n}}$
where $d{\textrm{Vol}_{2n}}$ is the Riemannian volume form and we recall that volumes can be compared.
Lemma 9.7.
Assume that $M$ is a $2n$-dimensional, connected, complete Riemannian manifold, and $\omega=d\alpha$ is an exact symplectic form on $\omega$.
Assume that $\omega$ is bounded on $M$ and uniformly nondegenerate. Let $B_{R}$ be a geodesic ball of radius $R>0$ in $M$ such that $B_{R}$ and $\partial B_{R}$ are piecewise differentiable manifolds.
Then, there exists a constant $\bar{C}>0$, depending on $\omega$ but not on $\alpha$, such that
(9.26) $\begin{split}&\int_{\partial B_{R}}|\alpha|\geq\bar{C}\cdot\textrm{Vol}_{2n}(B_{R}),\\ &\sup_{x\in\partial B_{R}}\|\alpha(x)\|\geq\bar{C}\cdot\textrm{Vol}_{2n}(B_{R})/\textrm{Vol}_{2n-1}(\partial B_{R}).\end{split}$
Proof.
Using the uniform non-degeneracy of $\omega$, (9.25), the assumption that $\omega$ is bounded, and Stokes’ Theorem, we have
$\begin{split}C\cdot\textrm{Vol}_{2n}(B_{R})\leq&\left|\int_{B_{R}}\omega^{n}\right|=\left|\int_{\partial B_{R}}\alpha\wedge\omega^{n-1}\right|\leq C^{\prime}\int_{\partial B_{R}}\|\alpha\|\\ \leq&C^{\prime}\cdot\sup_{x\in\partial B_{R}}\|\alpha(x)\|\cdot\textrm{Vol}_{2n-1}(\partial B_{R})\end{split}$
for some $C^{\prime}>0$, depending on the norm of $\omega$. ∎
Note that right hand sides of (9.26) has a factor depending on $R$ that depends only on the Riemannian metric. The symplectic properties enter only as a constant. Hence, we obtain that any $\alpha$ has to be unbounded using only properties of the Riemannian metric.
Example 9.8.
An application of the Lemma 9.7 is when $M=\mathbb{R}^{n}\times\mathbb{T}^{n}$ with the standard symplectic form $dI\wedge d\theta$, where $(I,\theta)\in\mathbb{R}^{n}\times\mathbb{T}^{n}$. In such a case, $\textrm{Vol}_{2n}(B_{R})\approx C_{1}R^{2n}$, $\textrm{Vol}_{2n-1}(\partial B_{R})\approx C_{2}R^{2n-1}$. Hence we obtain
(9.27) $\sup_{x\in\partial B_{R}}\|\alpha(x)\|>C\cdot R\textrm{ for all $R$ large.}$
We conclude that any action form in the manifold has to grow linearly. The standard symplectic form and action form saturate the bound and show that the result cannot be improved.
Remark 9.9.
When $M=T^{*}Q$ – the symplectic manifold is the cotangent bundle of a compact Riemannian manifold $Q$ – we see that there is $C^{\prime}>0$ such that $\sup_{x\in\partial B_{R}}\|\alpha(x)\|\approx C^{\prime}\cdot R\textrm{ for all $R$ large}$. This shows that the inequality (9.26) is sharp in this case.
Similarly, we can consider other action forms on cotangent bundles
(9.28) $\alpha=\alpha_{0}+\pi^{*}A$
where $A$ is a closed $1$-form on $Q$, and $\pi$ is the projection in the bundle $T^{*}Q$, and $\alpha_{0}$ is the standard action form.
The bound (9.26) also applies to this case (which appears in the study of magnetic fields).
10. Proof of Theorem 3.9
Part (A) is automatic in the pre-symplectic case.
The proof of Theorem 3.9 is basically walking through the proofs of Theorem 3.1 and Theorem 3.3. The proofs of the vanishing lemmas Lemma 6.5 and Lemma 6.8 do not require any change since the proofs are just iterating the definition, and neither closedness nor nondegeneracy play a role.
In the case the conformal factor is not constant (6.1) has to be adapted to:
[…]
Therefore the same proof of the vanishing lemma 6.8 works in the presymplectic case, under the conditions $|\lambda_{+}\mu_{+}\eta_{-}^{-1}|<1$ for the stable case and $|\lambda_{-}\mu_{-}\eta_{+}|<1$ in the unstable one.
Observe that, when the presymplectic factor $\eta$ is constant, the rate conditions (S’) entering in Theorem 3.9 are implied by the hypothesis (S) in Theorem 3.1.
On the other hand, the pairing rules (P) may fail to hold for presymplectic NHIMs, as in Example 5.4.
To obtain the proof of part (B) of Theorem 3.9 we can apply the proof of part (B) of Theorem 3.3 given in section 8.2.1. These proofs use the vanishing lemma 6.8 and the fact that the form is closed. Hence, go through without change. The proofs in Section 8.2.5, 8.2.7, which do not use any geometry, (but use different hyperbolicity rates) do not require any adaptation.
Analogously, the proofs of part (C) of Theorem 3.9 are basically the same as the ones to prove part (C) of Theorem 3.3 given in section 8.3.
Of course, part (D) on the leaf dynamics in Theorem 3.9 does not have an analogue in Theorem 3.1 and in Theorem 3.3 but it is an easy consequence of the conformal dynamics and the scattering map preserving the kernel of the presymplectic form.
Appendix A Summary of the theory on properties of NHIMs
In this appendix, we collect, without detailed proofs but with references, several results on the theory of NHIMs paying special care to the case of unbounded manifolds and the needed explicit uniformity assumptions.
The theory of NHIMs is very rich and there are many results we do not use in this paper (e.g existence of locally invariant foliations, linearization, persistence, etc.) and, hence, we do not mention them in this appendix. We will concentrate in the properties of (un)stable and strong (un)stable manifolds of NHIMs including dynamical characterizations and regularity properties.
Note that, in the unbounded case that we are considering in this paper, it is important to make assumptions that make it explicit that the properties of the manifold are uniform.
For us, the most important result of the theory of NHIM is the characterization of invariant objects and their regularity.
We consider the setting from Section 3.1 without assuming a symplectic structure on the manifold and that the map is conformally symplectic.
Theorem A.1.
Consider a manifold $M$ satisfying the assumption (U1), and $f:M\to M$ a $C^{r}$ diffeomorphism on $M$.
Let $\Lambda\subset M$ invariant under $f$, satisfy Definition 2.16 for rates $\lambda_{\pm},\mu_{\pm}$, which, moreover, satisfy condition (N).
Assume, furthermore, that the manifold $\Lambda$ and the stable and unstable bundles satisfy the uniformity assumption (U2).
Then, there exist (rather explicit) $\ell,\tilde{\ell},m,(\tilde{m},m)$ – called regularities of invariant objects – depending only on $r$, the regularity of the map, and the hyperbolicity rates $\lambda_{\pm},\mu_{\pm}$ such that:
  • (I)
    $\Lambda$ is a $\mathcal{C}^{\ell}$-manifold;
  • (II)
    There exist $0<\tilde{\rho}<\rho$, $\tilde{C},\tilde{D}>0$, and a $\mathcal{C}^{\tilde{\ell}}-$manifold $W^{s,\textrm{loc}}_{\Lambda}$ in $\mathcal{O}_{\tilde{\rho}}(\Lambda)$ described by the following equivalent conditions:
    • (i)
      $y\in W^{s,\textrm{loc}}_{\Lambda}$;
    • (ii)
      $f^{n}(y)\in\mathcal{O}_{\tilde{\rho}}(\Lambda)$, i.e., $d(f^{n}(y),\Lambda)\leq\tilde{\rho}$ for all $n\geq 0$;
    • (iii)
      $f^{n}(y)\in\mathcal{O}_{\tilde{\rho}}(\Lambda)$, i.e., $d(f^{n}(y),\Lambda)\leq\tilde{\rho}$ for all $n\geq 0$; and
      $\lim_{n\to+\infty}d(f^{n}(y),\Lambda)=0;$
    • (iv)
      $d(f^{n}(y),\Lambda)\leq\tilde{C}(\lambda_{+})^{n}$ for all $n\geq 0$;
    • (v)
      $d(f^{n}(y),\Lambda)\leq\tilde{D}\left(\mu_{-}\right)^{-n}$ for all $n\geq 0$;
    • (vi)
      There exists a unique $x\in\Lambda$ such that
      $d(f^{n}(x),f^{n}(y))\leq\tilde{C}(\lambda_{+})^{n}\quad\textrm{ for all }n\geq 0;$
      We denote such $x=\Omega_{+}(y)$.
    • (vii)
      For $x=\Omega_{+}(y)\in\Lambda$ we have.
      $d(f^{n}(x),f^{n}(y))\leq\tilde{D}\left(\mu_{-}\right)^{-n}\quad\textrm{ for all }n\geq 0;$
      (that is, if the orbit of $y$ converges to the orbit of $x\in\Lambda$ at a certain rate, it converges to another faster rate).
  • (III)
    Given $x\in\Lambda$, we denote for $0<\rho$ sufficiently small:
    $\begin{split}W^{s,\rm{\textrm{loc}}}_{x}=&\{y\in\mathcal{O}_{\rho}\,|\,d(f^{n}(y),f^{n}(x))\leq\tilde{C}(\lambda_{+})^{n}\textrm{ for all $n\geq 0$}\}\\ =&\{y\in\mathcal{O}_{\rho}\,|\,d(f^{n}(y),f^{n}(x))\leq\tilde{D}\left(\mu_{-}\right)^{-n}\textrm{ for all $n\geq 0$}\}.\end{split}$
    We refer to $W^{s,\textrm{loc}}_{x}$ as the local strong stable manifolds.
  • (IV)
    • (i)
      The manifold $W^{s,\textrm{loc}}_{\Lambda}$ is diffeomorphic to a neighborhood of the zero section in $E^{s}_{\Lambda}$.
    • (ii)
      Moreover, $W^{s,\rm{\textrm{loc}}}_{x}$ is a $\mathcal{C}^{m}$ manifold, diffeomorphic to a ball in $E^{s}_{x}$ centered at 0 and tangent to $E^{s}_{x}$ at 0.
    • (iii)
      As a consequence of II.(iv) and II.(v), $\{W^{s,\textrm{loc}}_{x}\}_{x\in\Lambda}$ is a $C^{\tilde{m},m}$ foliation of a neighborhood of $\Lambda$ in $W^{s,\textrm{loc}}_{\Lambda}$, in the sense of Definition 2.3.
There are similar characterizations of $W^{u,\textrm{loc}}_{\Lambda}$, $W^{u,\rm{\textrm{loc}}}_{x}$, involving negative times, which we leave to the reader. They can be obtained by noting that these unstable objects are the stable objects for the inverse map $f^{-1}$.
The regularities $\ell,\tilde{\ell},m,(\tilde{m},m)$ can be made as large as desired by making assumptions on $\lambda_{\pm},\mu_{\pm},r$. Hence, the assumptions $\bf(H1),(H2),(H3),(H4)$} are assumptions on the rates.
Remark A.2.
In some treatments, the statements of Theorem A.1 are given with the rates in (II)(iv), (II)(v), (II)(vi), (II)(vii), and (III) being $\lambda_{+}+\varepsilon$, $\mu_{-}+\varepsilon$ rather than $\lambda_{+}$, $\mu_{-}$, respectively, where $\varepsilon>0$ can be chosen arbitrarily, and $\tilde{\rho}$ depends on $\varepsilon$. If we used such a statement, Lemma B.3 would imply the statement of Theorem A.1, that is, we can get rid of the $\varepsilon$ terms by choosing the constants $\tilde{C}$, $\tilde{D}$ a little larger.
Remark A.3.
For the sake of readability, we have decided to consider only $\mathcal{C}^{r}$ regularity for integer $r$. In many settings, $\mathcal{C}^{r}$ is also defined for non-integer $r$ with the fractional part interpreted as Hölder regularity. The definition of Hölder regularity on manifolds is delicate since it involves comparing geometric objects at separate points. Even if the notion of Hölder function is non-controversial, the notion of Hölder distance or Hölder norm (needed to work out proofs) is cumbersome. Using fractional regularities is needed to obtain sharp regularity results.
Note that the mapping $(f,g)\rightarrow f\circ g$ is not continuous from $C^{0}\times C^{0}\rightarrow C^{0}$ unless $f$ is uniformly continuous. If one does not use fractional regularities for $f$, the only way to obtain uniform continuity for $f$ is to assume $f\in C^{1}$. This may lead to extra losses of regularity in the conclusions. One possibility used in several references is to include the uniform continuity of the highest derivative in the definition of $C^{r}$ but note that, when the domain is unbounded, the uniform continuity is not preserved under uniform limits, so that the space thus defined is not a Banach space.
The way that Theorem A.1 is usually proved is by representing the objects of interest using functions and solving functional equations that express invariance. Many of these functional equations involve the composition operator whose properties involve subtleties related to uniformity.
Theorem A.1 gives several equivalent characterizations of the local stable manifolds of $\Lambda$. The weaker ones are boundedness, others are just convergence and convergence with fast rates. The fact that these characterizations are equivalent is rather remarkable. Describing homoclinic excursions by intersections of manifolds, allows to take advantage of the regularity of the manifolds and their geometric properties to compute homoclinic excursions.
For a point $y$ to be in $W^{s,\textrm{loc}}_{x}$ it is important that the convergence of its orbit to the orbit of $x$ happens with a rate $\lambda_{+}$ faster than the rate of convergence $\mu_{+}$ of points in $\Lambda$ to the orbits of $x$. It is not a characterization by topological properties such as convergence of orbits.
The fact that if we fix $y\in W^{s,\textrm{loc}}_{\Lambda}$, $x\in\Lambda$, is determined uniquely, as claimed in (II.vi) of Theorem A.1, is crucial for this paper, since the projection from $y$ to $x$ – defined by $\Omega_{+}$ – is an important ingredient. This property requires assumption (U2). [Eldering12, Example 2.9] (See Figure 2) provides an example of NHIM for which $W^{s}_{x}=W^{s}_{x^{\prime}}$ for some $x\neq x^{\prime}$. Therefore, the decomposition of $W^{s}_{\Lambda}$ into strong stable leaves $W^{s,loc}_{x}$ in (2.15) is not a disjoint union  [^17] Since the relation given by $y\approx\tilde{y}\iff d(f^{n}(y),f^{n}(\tilde{y}))\leq C_{y,\tilde{y}}\lambda_{+}^{n},n\geq 0$ is an equivalence relation, the equivalence classes under this relation are leaves of a local foliation. Nevertheless, in [Eldering12, Example 2.9] a leaf can contain two points in $\Lambda$, so that the leaves cannot be labelled by their point of intersection with $\Lambda$. For our purposes, this prevents us from defining $\Omega_{+}$. .
To understand the phenomenon of non-unique projection $\Omega_{+}$ mentioned above, the following remark may be useful.
Remark A.4.
In the study of NHIM, it is natural to consider two distances on $\Lambda$: $d_{\Lambda}(x,x^{\prime})$ is the shortest length of paths in $\Lambda$ joining $x,x^{\prime}$, and $d_{M}(x,x^{\prime})$ is the shortest length of paths in $M$. Clearly, $d_{M}(x,x^{\prime})\leq d_{\Lambda}(x,x^{\prime})$. The distance that enters in the definition of $W^{s}_{x}$ is $d_{M}$. It is easy to see that for $\varepsilon>0$, $x,x^{\prime}\in\Lambda$, $d_{\Lambda}(f^{n}(x),f^{n}(x^{\prime}))\leq C(\lambda_{+}-\varepsilon)^{n},n>0$ implies $x=x^{\prime}$. On the other, hand, if $\Lambda$ folds into itself (as in [Eldering12, Example 2.9]) it is possible that for two different points $x,x^{\prime}\in\Lambda$, $d_{M}(f^{n}(x),f^{n}(x^{\prime}))\leq C\lambda_{+}^{n}$.
Remark A.5.
Note that the characterization $(IV)$ of local stable manifolds and locally strong stable manifolds involves the choice of a sufficiently small $\rho$. Clearly, using the characterization by rates, if we choose $\rho_{1}<\rho_{2}$, the set of manifolds corresponding to $\rho_{1}$ will be contained in those corresponding to $\rho_{2}$.
Remark A.6.
Note that the regularity of the manifolds $\Lambda$, $W^{u,\mathrm{loc}}_{\Lambda}$, $W^{s,\mathrm{loc}}_{\Lambda}$, is limited not just by the regularity of the map $f$. Next Example A.7 shows that the regularity of these objects also depends on relations between the rates. It shows that, even for analytic (indeed polynomial) maps, the NHIM could be only finitely differentiable. The differentiability is an expression in terms of the hyperbolic rates.
Example A.7.
Consider the map $f:{\mathbb{T}}^{d}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{T}}^{d}\times{\mathbb{R}}^{2}$ given by
(A.1) $f(\theta,s,u)=(A\theta,\lambda_{+}s+a_{s}(\theta),(1/\lambda_{-})u+a_{u}(\theta))$
where $a_{s},a_{u}:{\mathbb{T}}^{d}\rightarrow{\mathbb{R}}$ are continuous functions (the concrete example is a trigonometric polynomial), $A\in SL(d,\mathbb{Z})$ has spectrum contained in $1/\mu_{-},\mu_{+}$.
To make the example easier to analyze, we will assume that the leading modulus eigenvalues are simple and irrational. Hence, $\lambda_{\pm},\mu_{\pm}$ are real numbers as in Figure 1
$\lambda_{+}<1/\mu_{-}<1<\mu_{+}<1/\lambda_{-}.$
It is standard that $A$ can be interpreted either as a diffeomorphism of ${\mathbb{T}}^{d}$ (taking the action mod $1$) or as a linear map on ${\mathbb{Z}}^{d}$ (acting on the Fourier coefficients) [KatokH95].
We search for an invariant set of (A.1) of the form or a graph
$\Lambda=\{(\theta,b_{s}(\theta),b_{u}(\theta))\,|\theta\in{\mathbb{T}}^{d}\}.$
The image of a point on $\Lambda$ under $f$ is
$(A(\theta),\lambda_{+}b_{s}(\theta)+a_{s}(\theta),(1/\lambda_{-})b_{u}(\theta)+a_{u}(\theta)).$
This image is in $\Lambda$ if and only if
[…]
The above equations for $b_{s},b_{u}$ can be rearranged as
[…]
The equations (A.2) can be thought of as fixed point equations for an operator given by the right hand side. They admit an unique bounded solution obtained by iteration of the operator given by the right hand side.
Analyzing these solutions reveals that, in many cases, they possess only a finite number of continuous derivatives. We will see that the analysis is very similar to the analysis of the classical Weierstrass example 181818[Hardy16] considers the harder problem of no derivative of Weierstrass function at any point. .
A version of the argument close to the one here appears in [Llave92, Section 6.2]. Similar arguments appears often in hyperbolic systems.
We give the explicit formulas only for the stable case. The unstable one is very similar.
We observe that if (A.2) is to hold, substituting the right-hand side of (A.2) repeatedly, we obtain that for any finite $N$ we have:
$\begin{split}b_{s}(\theta)&=a_{s}(A^{-1}\theta)+\lambda_{+}a_{s}(A^{-2}\theta)+\lambda_{+}^{2}a_{s}(A^{-3}\theta)+\cdots+\lambda_{+}^{N}a_{s}(A^{-(N+1)}\theta)\\ &+\lambda_{+}^{N+1}b_{s}(A^{-(N+1)}\theta)\end{split}$
If $b_{s}$ were bounded (or in any $L^{p}({\mathbb{T}}^{d},\mathbb{R})$ ) the last term in the above formula would tend to zero. Hence the only possible bounded $b_{s}$ solving (A.2) is
(A.3) $b_{s}(\theta)=\sum_{j=0}^{\infty}(\lambda_{+})^{j}a_{s}(A^{-j-1}\theta)$
The series in (A.3) is uniformly convergent because the general term is bounded by a geometric series.
$\|(\lambda_{+})^{j}a_{s}\circ A^{-j-1}\|_{\mathcal{C}^{0}}\leq(\lambda_{+})^{j}\|a_{s}\|_{\mathcal{C}^{0}}$
Hence, (A.3) defines a continuous function on the torus and we can compute its Fourier coefficients term by term.
Now we analyze (A.3) to show that it cannot be very differentiable.
The chain rule gives $D^{m}(a_{s}\circ A^{-j-1})=(D^{m}a_{s})\circ A^{-j-1}(A^{-j-1})^{\otimes m}$ and, if $Av=\mu v$, we have:
$(v\cdot\partial)^{m}(a_{s}\circ A^{-j-1})=((v\cdot\partial)^{m}a_{s})\circ A^{-j-1}\mu^{m}|v|^{m}.$
So, for high enough $m$, the derivation term by term of (A.3) becomes problematic in general.
To show that indeed the sum (A.3) has only a limited number of derivatives, we recall that for a $\mathcal{C}^{\ell}$ function $b$ we have $|\hat{b}_{k}|\leq C|k|^{-\ell}$, where $\hat{b}_{k}$ are its Fourier coefficients.
If we take $a_{s}$ in (A.3) to be $a_{s}(\theta)=\cos(2\pi k_{0}\cdot\theta)$, then $a_{s}(A^{j-1}\theta)=\cos(2\pi A^{j-1}k_{0}\cdot\theta)$
We have that if $\tilde{k}=(A^{-(j+1)})^{T}k_{0}$, there is only one term in (A.3) with $\tilde{k}$ index.
Therefore, $|\widehat{(b_{s})}_{\tilde{k}}|=\frac{1}{2}\lambda_{+}^{j}$. Since $|\tilde{k}|\leq C\mu_{-}^{j+1}$, we see that it is impossible to have an inequality of the form $|\widehat{(b_{s})}_{k}|\leq C|k|^{-\ell}$ if $\ell$ is large enough that $\lambda_{+}\mu_{-}^{\ell}>1$ and therefore, the $b_{s}$ corresponding to $a_{s}$ as before is not $\mathcal{C}^{\ell}$.
A similar restriction happens for the unstable part.
Remark A.8.
The Example A.7 gives an idea of what are the optimal regularities.
The final optimal regularities depend however upon subtleties such as those in Remark A.3.
In the compact case, this example gives the limit of the regularity that can be obtained using the rates as input. There are proofs that reach this limit.
The example can be modified to yield restrictions on regularity even if the map $f$ is furthermore assumed to be symplectic (take $d$ even, $A$ symplectic on ${\mathbb{T}}^{d}$ and $\lambda_{+}/\lambda_{-}=1$).
The argument above can also be used for fractional regularities and also to conclude that this lack of differentiability is generic (indeed, when $\lambda_{+}\mu_{-}^{\ell}>1$, $b_{s}$ is not $\mathcal{C}^{\ell}$ for all trigonometric polynomials except in a linear space of infinite codimension).
One interesting problem is to study deeper geometric properties (fractal dimension, directional derivatives of these examples).
Once we have the local stable and unstable manifolds, we define the global stable and unstable manifolds.
[…]
This follows from the observation that for any point $y\in W^{s}_{x}$, there exist $N$ such that $f^{N}(y)\in W^{s,\textrm{loc}}_{f^{N}(x)}$. As a consequence, we obtain the point II.(v) in Theorem A.1, giving a characterization of the global stable manifold by rates of convergence.
[…]
$\begin{split}&W^{s}_{\Lambda}=\bigcup_{x\in\Lambda}W^{s}_{x},\qquad W^{u}_{\Lambda}=\bigcup_{x\in\Lambda}W^{u}_{x},\qquad\\ &W^{s}_{x}\cap W^{s}_{x^{\prime}}=\emptyset,\qquad W^{u}_{x}\cap W^{u}_{x^{\prime}}=\emptyset,\textrm{ for }x\neq x^{\prime}.\end{split}$
The above can be described as saying that the decomposition above is a foliation of $W^{s}_{\Lambda}$ with leaves $W^{s}_{x}$. Note that the leaves are in general more regular than $W^{s}_{\Lambda}$. The regularity of $W^{s}_{\Lambda}$ is determined both by the regularity of the leaves and the way that they fit together.
There are many proofs of Theorem A.1, some of them have more assumptions. For the purposes of this paper, an efficient proof of the local stable/unstable manifolds is in [Pesin, p. 33–38]. This proof is based on constructing first the embeddings giving the local stable/unstable manifolds by studying the functional equations satisfied via a fixed point argument. Note that this proof requires uniformity assumptions for the map or the derivative in a neighborhood of $\Lambda$.
Remark A.9.
Large parts of Theorem A.1 require only (U1).
Nevertheless, [Eldering12, Example 3.8] (Illustrated here in Figure 2 ) shows that to get (II).(vi)(II).(vi),$(II).(vii)$, $(III)$ and all the other properties of $W^{s,\textrm{loc}}_{x}$ in Theorem A.1 one needs the uniformity assumption (U2). For the purposes of this paper, the fact that $W^{s}_{\Lambda}$ is foliated by $W^{s}_{x}$ is crucial since it is what we use to define $\Omega_{+}$ and the scattering map.
Remark A.10.
The theory of regularity based on hyperbolicity rates as in Theorem A.1 is not the only possible way to establish regularity of invariant objects.
The paper [Fenichel74] establishes regularity based on the study of the numbers $\alpha$ for which
(A.4) $\sup_{x}\|Df(x)_{\mid E^{s}}\|\|Df^{-1}(x)_{\mid T_{x}\Lambda}\|^{\alpha}<1.$
which is sharper than taking the supremum on all factors. as is done in [HPS].
Conditions similar to (A.4) are used in the invariant cone approach to regularity of invariant manifolds [CapinskiZ15].
For the purposes of this paper, the use of rates is more natural because rates enter in the formulation of vanishing lemmas. We have, however, formulated the standing assumptions (H1)-(H4) and their variants as regularity assumptions that can be verified in concrete examples either using rates or (A.4) or any other method.
Boundedness of symplectic form indeed plays a role in the main results of this paper, starting with the vanishing lemmas (Section 6). When the symplectic form is unbounded some of the main results do not hold (see Example 5.5). On the other hand, in Section 6.8 we give some results that show that the boundedness assumption on the symplectic form can be weakened at the price of other assumptions on rates. This seems like a possible line of research that we hope to pursue in the future.
Appendix B Some basic results on hyperbolicity rates
In this appendix, we collect and prove some basic results on hyperbolicity rates.
B.1. Relations between forward and backwards uniform rates
Lemma B.1.
Let $E$ be an invariant vector bundle over some space $X$. The statement
$\forall\,x\in X,\,v\in E_{x},\,\|Df^{n}(x)v\|\leq C\lambda^{n}\|v\|\textrm{ for all $n\geq 0$}$
is equivalent to the statement
$\forall\,y\in X,\,w\in E_{y},\,\|Df^{-n}(y)w\|\geq\tilde{C}\lambda^{-n}\|w\|\textrm{ for all $n\geq 0$}.$
Proof.
Given $y\in X$ and $w\in E_{y}$, by setting $y=f^{n}(x)$, $w=Df^{n}(x)v$, we have,
$Df^{n}(x)v=w\Leftrightarrow Df^{-n}(y)w=v$
and therefore
$\|w\|=\|Df^{n}(x)v\|\leq C\lambda^{n}\|v\|=C\lambda^{n}\|Df^{-n}(y)w\|$
which gives the result with $\tilde{C}=\frac{1}{C}$. ∎
Lemma B.1 illustrates one of the advantages of the convention of denoting the rates as in (R) and (H), namely, if we change $f$ to $f^{-1}$, it suffices to change $\mu_{+}$ to $\mu_{-}$ and $\lambda_{+}$ to $\lambda_{-}$.
Lemma B.2.
For the rates (R) satisfying (H) we always have:
(B.1) $\lambda_{+}<\frac{1}{\lambda_{-}}\textrm{ and }\mu_{+}\geq\frac{1}{\mu_{-}}.$
Proof.
The reason for the second inequality is that if for $n\geq 0$ we have for all $x\in\Lambda,v\in T_{x}\Lambda$:
$\|Df^{n}(x)(v)\|\leq D_{+}\mu_{+}^{n}\|v\|$
then, by Lemma B.1, we also have for $y\in\Lambda,w\in T_{y}\Lambda$
$\|Df^{-n}(y)(w)\|\geq D_{+}^{-1}\mu_{+}^{-n}\|w\|.$
Thus, using (H), $\mu_{+}^{-1}\leq\mu_{-}$.
The inequality $\lambda_{+}<\lambda_{-}^{-1}$ follows from the previous result about $\mu_{+}$ and $\mu_{-}$ and the normal hyperbolicity assumptions (R). ∎
With the above considerations, we can also write as a characterization of the tangent space to the NHIM.
$v\in T_{x}\Lambda\iff\tilde{D}_{-}\mu_{-}^{-n}\|v\|\leq\|Df^{n}(x)v\|\leq\tilde{D}_{+}\mu_{+}^{n}\|v\|,\quad\forall n\geq 0.$
B.2. Hyperbolicity rates in the stable manifold
In this section we study the hyperbolicity rates for tangent vectors to the stable manifold. The intuition is that for orbits that converge to the manifold $\Lambda$, most of the factors are derivatives near the manifold. Of course, the convergence may be slow, therefore, one needs some precisions on the statements.
Lemma B.3.
Choosing constants $C_{+}$, $D_{+}>0$ larger than those in the manifold $\Lambda$, for all $y\in W^{s,\mathrm{loc}}_{x}\cap\mathcal{O}_{\rho}$ (resp., $y\in\cap W^{s,\mathrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$) we have:
(B.2) $\begin{split}v\in T_{y}W^{s,\mathrm{loc}}_{x}\Leftrightarrow&\|Df^{n}(y)(v)\|\leq C_{+}\lambda_{+}^{n}\|v\|\textrm{ for all }n\geq N,\\ v\in T_{y}W^{s,\mathrm{loc}}_{\Lambda}\Leftrightarrow&\|Df^{n}(y)(v)\|\leq D_{+}\mu_{+}^{n}\|v\|\textrm{ for all }n\geq N.\end{split}$
An analogous property holds for points in $W^{u,\mathrm{loc}}_{\Lambda}\cap\mathcal{O}_{\rho}$.
Remark B.4.
Recalling that if $y\in W^{s,\mathrm{loc}}_{\Lambda}$, there exists $N=N(y)>0$ such that $f^{N}(y)\in\mathcal{O}_{\bar{\rho}}$, we have that (B.2) holds for any $y\in W^{s,\mathrm{loc}}_{\Lambda}$ but with constants $C_{\pm},D_{\pm}$ which depend on $y$.
Remark B.5.
Versions of Lemma B.3 appear in the classical references with the rates of convergence are slightly worse that those in the NHIM. These results imply that the stable/unstable manifolds are NHIMs themselves.
Lemma B.3 is an improvement from previous results in the literature, because the rates claimed in the stable manifold are exactly the same as the rates for the linearization in the NHIM. For most of the results in this paper, the classical results with slightly worse rates are enough, so the proof can be just skimmed.
Lemma B.3 will be a consequence of the following preliminary result (Proposition B.6), the chain rule,
(B.3) $Df^{n}(y)=Df(f^{n-1}(y))Df(f^{n-2}(y))\cdots Df(y),$
and Proposition B.7. We postpone the details of the proof of Lemma B.3 after these two propositions.
The first preliminary result is fairly standard and is indeed enough for many applications. For us, it will be the first step and it will be later bootstrapped to get Lemma B.3.
Proposition B.6.
In the conditions of Lemma B.3, given $\varepsilon>0$ we can find a radius $0<\bar{\rho}=\bar{\rho}(\varepsilon)<\rho$ and a constant $\tilde{C}_{+}=\tilde{C}_{+}(\varepsilon)$ such that if we take any $y\in W^{s,\mathrm{loc}}_{\Lambda}$, $d(y,\Lambda)\leq\bar{\rho}$, and take $x=\Omega^{+}(y)$, so that $y\in W^{s,\mathrm{loc}}_{x}$, we have for all $n>0$, and for all $v\in T_{y}W_{x}^{s,\mathrm{loc}}$
  • (i)
    $\|Df^{n}(y)v\|\leq\tilde{C}_{+}(\lambda_{+}+\varepsilon)^{n}\|v\|$
  • (ii)
    As a consequence,
    $d(f^{n}(x),f^{n}(y))\leq\rho\tilde{C}_{+}(\lambda+\varepsilon)^{n}.$
Proof.
Because of the uniformity of the bounds assumed in Definition 2.16 we have that, for all $x\in\Lambda$ we have, for any $n\geq 0$:
$\|Df^{n}(x)|_{E^{s}_{f^{n}(x)}}\|\leq C_{+}\lambda_{+}^{n}$
Of course, $E^{s}_{f^{n}(x)}=T_{f^{n}(x)}W^{s}_{f^{n}(x)}$ and this will be useful later.
Once we choose $\varepsilon>0$, we can find $N=N(\varepsilon)>0$ so that,
$C_{+}\lambda_{+}^{N}\leq\frac{1}{10}(\lambda_{+}+\varepsilon)^{N}$
Therefore, for any $x\in\Lambda$ we have:
$\|Df^{N}(x)|_{E^{s}_{x}}\|\leq\frac{1}{10}(\lambda_{+}+\varepsilon)^{N}.$
Because of the chain rule (B.3) and the uniformity of the bounds on the derivatives and the uniform continuity of the derivatives, we can find $\bar{\rho}=\bar{\rho}(\varepsilon)>0$ so that for all $y\in W^{s}_{x}$, $d(y,x)\leq\rho$ we have
$\|Df^{N}(y)|_{T_{y}W^{s}_{x}}\|\leq(\lambda_{+}+\varepsilon)^{N}.$
Using (B.3) and that all the derivatives of $f$ are uniformly bounded and uniformly continuous, we have that there exists a (big enough) constant $\tilde{C}$ such that
$\|Df^{j}(y)|_{T_{y}W^{s}_{x}}\|\leq\tilde{C}_{+}(\lambda_{+}+\varepsilon)^{j}\quad\textrm{ for }0<j<N$
Using that any positive number $n$ can be written $n=kN+j$ with $0<j\leq N$ we have, using again (B.3),
$\|Df^{n}(y)|_{T_{y}W^{s}_{x}}\|\leq(\lambda_{+}+\varepsilon)^{kN}\tilde{C}_{+}(\lambda_{+}+\varepsilon)^{j}=\tilde{C}_{+}(\lambda_{+}+\varepsilon)^{n}$
From this, using that $d(f^{n}(x),f^{n}(y))\leq\int_{0}^{1}\|Df^{n}(\gamma(s))\|\gamma^{\prime}(s)|\,ds$ we obtain the last consequence. ∎
B.2.1. A sharp result on perturbations of products of a sequence of operators (cocycles)
The following result assumes rates of growth for a sequence of successive products of a sequence of operators (sometimes called cocycles). and the factors are changed by a summable sequence, then, the new sequence grows at the same rate.
The interesting thing for us is that it shows that the rates of growth of vectors do not need to be modified at all.
Proposition B.7.
Given a sequence of Banach spaces $\{X_{j}\}_{j\in\mathbb{N}}$
Let $\{\alpha_{j}\}_{j\in\mathbb{N}}$ and $\{\beta_{j}\}_{j\in\mathbb{N}}$ be two sequences of operators $\alpha_{j},\beta_{j}:X_{j}\rightarrow X_{j+1}$ Denote by $A_{n,m},B_{n,m}$ the associated cocycles
[…]
Therefore, we have for all $n-1\geq m\geq l$, $A_{n,m}A_{m,l}=A_{n,l}$, $B_{n,m}B_{m,l}=B_{n,l}$.
Assume that
  • (i)
    $\exists C,\lambda>0$ s.t. $\forall n>m\in\mathbb{N}$
    $|A_{n,m}|\leq C\lambda^{(n-m)},$
  • (ii)
    $\sum_{j=0}^{\infty}|\beta_{j}-\alpha_{j}|<\infty.$
Then, there exists $\tilde{C}$ such that for all $n-1\geq m$
$|B_{n,m}|\leq\tilde{C}\lambda^{(n-m)}$
Proof.
If we multiply $\alpha$, $\beta$ by a constant $\sigma$, then both $A_{n,m},B_{n,m}$ multiply by $\sigma^{(n-m)}$, so without loss of generality, we can assume that $\lambda<1$ in the hypothesis of Lemma B.7.
We fix $n>m$ and note that we have adding and subtracting and grouping the terms with different numbers of factors $\beta-\alpha$
$\begin{split}B_{n,m}=&A_{n,m}+\sum_{n-1\geq j_{1}\geq m}A_{n,j_{1}+1}(\beta_{j_{1}}-\alpha_{j_{1}})A_{j_{1},m}\\ &+\sum_{n-1\geq j_{1}>j_{2}\geq m}A_{n,j_{1}+1}(\beta_{j_{1}}-\alpha_{j_{1}})A_{j_{1},j_{2}+1}(\beta_{j_{2}}-\alpha_{j_{2}})A_{j_{2},m}\\ &+\cdots\\ &+(\beta_{n-1}-\alpha_{n-1})\cdots(\beta_{m}-\alpha_{m}).\end{split}$
If we write
$B_{n,m}=(\alpha_{n}+(\beta_{n}-\alpha_{n}))\cdots(\alpha_{m}+(\beta_{m}-\alpha_{m}))$
and expand the product and group by the number of factors $\beta-\alpha$.
The general term in the sum above is obtained by replacing $k$ factors $\alpha$ in the expression for $A_{n,m}$ with $\beta_{j}-\alpha_{j}$ and leaving all the others as $\alpha_{j}$. The products of consecutive factors $\alpha_{j}$ are transformed into the $A$ cocycles.
Using the assumptions, the first term in the sum above is bounded by
$\|A_{n,m}\|\leq C\lambda^{(n-m)}.$
The second term is bounded
$\begin{split}\|\sum_{n-1\geq j_{1}\geq m}A_{n,{j_{1}}+1}(\beta_{j_{1}}-\alpha_{j})A_{{j_{1}},m}\|&\leq C^{2}\sum_{n-1\geq j_{1}\geq m}\lambda^{n-{j_{1}}-1}\|\beta_{j_{1}}-\alpha_{j_{1}}\|\lambda^{{j_{1}}-m}\\ &\leq C^{2}\lambda^{-1}\lambda^{(n-m)}\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\end{split}$
Similarly, we bound the next term as:
$\begin{split}&\|\sum_{n-1\geq j_{1}>j_{2}\geq m}A_{n,j_{1}+1}(\beta_{j_{1}}-\alpha_{j_{1}})A_{j_{1},j_{2}+1}(\beta_{j_{2}}-\alpha_{j_{2}})A_{j_{2},m}\|\\ &\quad\leq C^{3}\sum_{n-1\geq j_{1}>j_{2}\geq m}\lambda^{n-j_{1}-1}\|\beta_{j_{1}}-\alpha_{j_{1}}\|\lambda^{j_{1}-j_{2}-1}\|\beta_{j_{2}}-\alpha_{j_{2}}\|\lambda^{j_{2}-m}\\ &\quad\leq C^{3}\lambda^{-2}\lambda^{n-m}\frac{1}{2}\left(\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\right)^{2}.\end{split}$
Note that both bounds include factors $\lambda^{n-m}$ (remember we are assuming $\lambda<1$). Similarly the general term consists in product of cocycles of $(k+1)$ intervals that cover [m,n1][m,n-1] with $k$ factors that have been changed into $\beta_{j}-\alpha_{j}$.
The last term is bounded (very wastefully) as
$C^{n-m+1}\lambda^{-(n-m)}\lambda^{(n-m)}\frac{1}{(n-m)!}\left(\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\right)^{n-m}.$
Therefore, adding all the bounds we get:
$\lambda^{n-m}C\sum_{k=0}^{n-m}C^{k}\frac{1}{k!}\lambda^{-k}\left(\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\right)^{k}\leq\lambda^{n-m}C\exp\left(C\lambda^{-1}\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\right).$
We conclude:
$B_{n,m}\leq\lambda^{n-m}C\left(1+\exp\left(C\lambda^{-1}\sum_{j=0}^{\infty}\|\beta_{j}-\alpha_{j}\|\right)\right):=\lambda^{n-m}\tilde{C}$.
Proof of Lemma B.3.
If the manifold $M$ is Euclidian and the bundle $E^{s}$ is trivial, we can identify all the tangent spaces with an Euclidean space.
In such a geometrically trivial case, to prove the first inequality in Lemma B.3 it suffices to take:
$\begin{split}&\alpha_{j}=Df(f^{j}(x)):E^{s}_{f^{j}(x)}\rightarrow E^{s}_{f^{j+1}(x)},\\ &\beta_{j}=Df(f^{j}(y)):T_{f^{j}(y)}(f^{j}(W^{s}_{x}))\rightarrow T_{f^{j}(y)}(f^{j}(W^{s}_{x})),\end{split}$
but, in the geometrically trivial case, we identify $E^{s}_{f^{j}(x)}$ with $T_{f^{j}(y)}(f^{j}(W^{s}_{x})$.
Similarly, to prove the second inequality Lemma B.3 it suffices to take:
$\begin{split}&\alpha_{j}=Df(f^{j}(x)):E^{s}_{f^{j}(x)}\oplus T_{f^{j}(x)}\Lambda\rightarrow E^{s}_{f^{j+1}(x)}\oplus T_{f^{j+1}(x)}\Lambda,\\ &\beta_{j}=Df(f^{j}(y)):T_{f^{j}(y)}(W^{s}_{\Lambda})\rightarrow T_{f^{j+1}(y)}(f^{j+1}(W^{s}_{\Lambda})).\end{split}$
In the case that $M$ is a manifold or $E^{s}$ is a non-trivial bundle, use a system of coordinates on $\mathcal{O}_{\rho}$ assumed to exist in (U2). In such a geometric adaptation, we need to include explicitly the connectors (see (2.1)) identifying the neighboring spaces and check that all remains uniform in the number of iterates. Even if it is mostly routine, we include the details.
To prove the first inequality in Lemma B.3 we observe that, geometrically:
$\begin{split}&Df(f^{n}(y)):T_{f^{n}(y)}f^{n}(W^{s}_{x})\rightarrow T_{f^{n+1}(y)}f^{n+1}(W^{s}_{x}),\\ &Df(f^{n}(x)):T_{f^{n}(x)}f^{n}(W^{s}_{x})\equiv E^{s}_{f^{n}(x)}\rightarrow T_{f^{n+1}(x)}f^{n+1}(W^{s}_{x})\equiv E^{s}_{f^{n}(x)}.\end{split}$
The operators $Df(f^{n}(y))$, $Df(f^{n}(x))$ act in different spaces and we have to identify them.
Using the system of coordinates in $\mathcal{O}_{\rho}$ we can use connectors (see (2.1)) to identify $E^{s}_{x}$ with $T_{y}W^{s}_{x}$. A geometrically natural way to identify these spaces is to take $S_{x}^{y}=(D\exp_{x})(\exp^{-1}_{x}(y))$. We note that the norm of these operators is bounded and in $\mathcal{O}_{\rho}$ and the norm becomes close to $1$ if $\rho$ is small.
We apply the Proposition B.7 taking $X_{j}=E^{s}_{f^{n}(x)}$,
$\begin{split}\alpha_{j}=&Df(f^{n}(x)),\\ \beta_{j}=&\left(S_{f^{j+1}(x)}^{f^{j+1}(y)})\right)^{-1}Df(f^{j}(x))S_{f^{j}(x)}^{f^{j}(y)}.\end{split}$
Note that
$B_{n,m}=\left(S_{f^{n+1}(x)}^{f^{n+1}(y)})\right)^{-1}Df^{n-m}(f^{m}(x))S_{f^{j}(x)}^{f^{j}(y)},$
so that the bounds on the cocycle are equivalent to the desired bounds on the derivatives. ∎
Appendix C Rates of convergence of homoclinic channels to a NHIM
The goal of this section is to study quantitatively the convergence of the iterates of channel $\Gamma$ to the invariant manifold $\Lambda$. The explicit values of the rates of convergence enter into several proofs. For example Sections 8.2.5, 8.2.7 as well in other future work.
More precisely, we prove:
Lemma C.1.
With the notations in the previous sections, assume, for simplicity of statements and without loss of generality, that
(C.1) $\mu_{+},\mu_{-}>1.$
  • (i)
    Assume that $r\geq 2$ and the foliation of $W^{s,u,\textrm{loc}}_{\Lambda}$ by $W^{s,u,\textrm{loc}}_{x}$ is of class $\mathcal{C}^{1,1}$.
    • (i.a)
      If $\lambda_{+}\mu_{-}<1$ then:
      (C.2) $d_{\mathcal{C}^{1}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{+}\mu_{-})^{n},\quad n>0.$
    • (i.b)
      If $\lambda_{-}\mu_{+}<1$ then:
      (C.3) $d_{\mathcal{C}^{1}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{-}\mu_{+})^{|n|},\quad n<0.$
  • (ii)
    Assume moreover that $r\geq j+1$ and the foliation of $W^{s,u,\textrm{loc}}_{\Lambda}$ by $W^{s,u,\textrm{loc}}_{x}$ is of class $\mathcal{C}^{j,1}$.
    • (ii.a)
      If $\lambda_{+}\mu_{-}^{j}<1$ then:
      (C.4) $d_{\mathcal{C}^{j}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{+}\mu_{-}^{j})^{n},\quad n>0.$
    • (ii.b)
      If $\lambda_{-}\mu_{+}^{j}<1$ then:
      (C.5) $d_{\mathcal{C}^{j}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{-}\mu_{+}^{j})^{|n|},\quad n<0.$
Remark C.2.
The Lemma C.1 is closely related to results in the literature such as the graph transform, the inclination lemma – a.k.a. $\lambda$-lemma, and even closer, to fiber contraction lemma, or the more general cone conditions, which have many variants. In our case, the result is easier since we already know the fixed point of the contraction. Unfortunately, many of the versions in the literature are only qualitative or involve extra assumptions (e.g. [HPS, p. 35] assumes compactness).
The tangent functor trick [AbrahamR67, HirschP70] – which we explain later – shows that the case $j=1$ implies the results for other $j$ in Lemma C.1. The case $j=1$ we present, basically goes back to [Hadamard98] (translated to English in [Hasselblatt17]).
C.1. Proof of Lemma C.1
We will consider here only the case of $n\rightarrow\infty$. The case $n\rightarrow$-$\infty$ is identical, up to a change in typography.
By Theorem A.1, item (II) (iv), by choosing a suitable constant $C$ we have that
$d_{\mathcal{C}^{0}}(f^{n}(\Gamma),\Lambda)\leq C\lambda_{+}^{n}\textrm{ for }n>0.$
Hence, it suffices to prove the estimates under the extra assumption that $\Gamma$ and its iterates remain in small neighborhood of $\Lambda$. Concretely, we will assume that
$\Gamma\subset\mathcal{O}_{\rho},$
where $\mathcal{O}_{\rho}$ is the neighborhood of $\Lambda$ given in assumption U2. Furthermore, to estimate derivatives, we can work in arbitrary small patches.
In $\mathcal{O}_{\rho}$, and in small enough patch, we can take the system of coordinates of section 2.8 given in (2.17) such that $W^{s,\textrm{loc}}_{x}$ can be identified with
$W^{s,\textrm{loc}}_{x}\simeq\{(x,y)\,|y\in B_{\rho}(0)\}$
with $B_{\rho}(0)\subset E^{s}_{x}$.
All such coordinate systems can be made to have uniform differentiability properties. Remember that the foliation $W^{s,\textrm{loc}}_{x}$ is invariant (the leaves are mapped into leaves by the dynamics). Therefore, in this system of coordinates, the map can be represented as
(C.6) $f(x,y)=(a(x),b_{x}(y))$
The map $b_{x}$ represents in coordinates the motion on the leaf $W^{s,\textrm{loc}}_{x}$, to the leaf $W^{s,\textrm{loc}}_{a(x)}$, where $x$ and $a(x)$ represent the foot-points of those leaves.
Note that the coordinate patches are different for the domain and the range. Even if the domain and the range patches overlap, we do not identify the coordinates. We do not attempt to identify the points belonging to two coordinate patches. Note, however, that the coordinate patches will only enter in the proof to perform some algebraic operations with derivatives and the patches covered by the coordinate systems may be arbitrarily small.
The derivative of (C.6) is, in these coordinates, given by an upper diagonal matrix.
$Df(x,y)=\begin{pmatrix}Da(x)&0\\ \partial_{x}b_{x}(y)&Db_{x}(y)\end{pmatrix}$.
We can write $\Gamma$ as the graph of a section on $E^{s}$ in the foliation by the $W^{s,\textrm{loc}}_{x}$:
$\Gamma=\{(x,\sigma(x)),\ x\in\Lambda\}.$
Given a point $p=(x,\sigma(x))$ in the section, we consider its orbit $f^{n}(p):=p_{n}=(x_{n},y_{n})$.
For future reference, we compute some explicit and elementary expressions for the derivatives of iterations of $f$ (similarly to [DLS08, Proposition 15]).
To simplify notation, we write: $p_{n}=f^{n}(x,y)$ and,
$Df(p_{n})=\begin{pmatrix}A_{n}&0\\ C_{n}&B_{n}\end{pmatrix}$.
In the notation for $A_{n},B_{n},C_{n}$ we omit the dependence on the point $p_{n}$.
We have:
[…]
Following [Hadamard01, Hasselblatt17], we consider $(\Delta,(D\sigma(x))\Delta)$ the representation of a tangent vector of $\Gamma$ at $(x,\sigma(x))$. When $\Delta$ ranges over all the possible values, $(\Delta,(D\sigma(x))\Delta)$ ranges over the graph of $D\sigma(x)$, the tangent space of $\Gamma$.
We have that
$Df(p)\begin{pmatrix}\Delta\\ d\sigma(p)\Delta\end{pmatrix}=\begin{pmatrix}A_{0}\Delta\\ C_{0}\Delta+B_{0}D\sigma(p)\Delta\end{pmatrix}$,
$Df^{2}(p)\begin{pmatrix}\Delta\\ d\sigma(p)\Delta\end{pmatrix}=\begin{pmatrix}A_{1}A_{0}\Delta\\ (C_{1}A_{0}+B_{1}C_{0})\Delta+B_{1}B_{0}D\sigma(p)\Delta\end{pmatrix}$,
and, for $n\geq 3$:
$\begin{split}Df^{n}(p)\begin{pmatrix}\Delta\\ D\sigma(p)\Delta\end{pmatrix}&=\begin{pmatrix}A_{n-1}\cdots A_{0}\Delta\\ E_{n}\Delta\end{pmatrix}\end{split}$
where
$\begin{split}E_{n}=&\sum_{k=1}^{n-2}B_{n-1}\cdots B_{k+1}C_{k}A_{k-1}\cdots A_{0}\\ &+C_{n-1}A_{n-2}\cdots A_{0}+B_{n-1}\cdots B_{1}C_{0}+B_{n-1}\cdots B_{0}D\sigma(p).\end{split}$
As $\Delta$ ranges over the tangent space of $\Gamma$ at $(x,\sigma(x))$, the above iterated tangent space can be described as the graph of the function obtained by writing the second component as a function of the first component.
That is, the graph of the iterated tangent space is the graph of the function:
(C.7) $\begin{split}&\Big[\sum_{k=1}^{n-2}B_{n-1}\cdots B_{k+1}C_{k}(A_{n-1}\cdots A_{k})^{-1}\\ &+C_{n-1}A_{n-1}^{-1}+B_{n-1}\cdots B_{1}C_{0}(A_{n-1}\cdots A_{0})^{-1}\Big]\\ &+B_{n-1}\cdots B_{0}D\sigma(p)(A_{n-1}\cdots A_{0})^{-1}.\end{split}$
Now, we proceed to estimate the terms in (C.7).
By Theorem A.1, items II (iv), (v), the last term in (C.7) is straightforwardly estimated by
(C.8) $\|B_{n-1}\cdots B_{0}\,D\sigma(p)\,(A_{n-1}\cdots A_{0})^{-1}\|\leq C(\lambda_{+})^{n}\|D\sigma\|_{\mathcal{C}^{0}}(\mu_{-})^{n},$
where we combine all constants into a new constant which we still denote by $C$ (as we will do in subsequent estimates).
The first term of (C.7) require a bit more care. We observe that, since $b_{x}(0)=0$ for all $x$ we have $\partial_{x}b_{x}(0)=0$. Taking into account that $C_{k}=\partial_{x}b(y_{k})$, and that $\|y_{k}\|\leq C\lambda_{+}^{k}$, since the foliation $\{W^{s,\textrm{loc}}_{x}\}_{x}$ is $\mathcal{C}^{1,1}$ and using Schwarz’ theorem on mixed partials, we have $\|C_{k}\|\leq C\lambda_{+}^{k}$.
Hence the first term of (C.7) can be estimated by:
(C.9) $\begin{split}&\sum_{k=1}^{n-2}C(\lambda_{+})^{n-k-1}(\lambda_{+})^{k}(\mu_{-})^{n-k}+C(\lambda_{+})^{n-1}\mu_{-}+C(\lambda_{+}\mu_{-})^{n-1}\\ &\leq C(\lambda_{+}\mu_{-})^{n}\left(\sum_{k=1}^{n-2}(\lambda_{+})^{-1}(\mu_{-})^{-k}+(\lambda_{+})^{-1}(\mu_{-})^{1-n}+(\lambda_{+}\mu_{-})^{-1}\right)\\ &\leq C(\lambda_{+}\mu_{-})^{n},\end{split}$
where in the last inequality we have used (C.1). Combining (C.8) and (C.9) gives the desired result and finishes the proof of the case $j=1$ of Lemma C.1.
Once we have established the case $j=1$ of Lemma C.1, the other cases are a corollary.
For this, we use the ‘tangent functor trick’ [AbrahamR67, HirschP70]. Note that given $f:M\rightarrow N$, $g:N\rightarrow P$, differentiable maps among manifolds, defining $Tf:TM\rightarrow TN$, $Tg:TN\rightarrow TP$ by $Tf(x,v)=(f(x),Df(x)v)$, we have:
$T(g\circ f)=Tg\circ Tf.$
Also, if $\Lambda$ is an invariant manifold for $f$, $T\Lambda$ is an invariant manifold for $Tf$.
Applying the $j=1$ result for the $\mathcal{C}^{1}$-convergence of $T\Gamma$ to $T\Lambda$ under $Tf$ iteration, with $\lambda_{+}$ replaced by $\lambda_{+}\mu_{-}$, we obtain
$d_{\mathcal{C}^{1}}(Tf^{n}(T\Gamma),T\Lambda)\leq C(\lambda_{+}\mu_{-}^{2})^{n},\quad n>0,$
which implies the $\mathcal{C}^{2}$ convergence of $\Gamma$ under $f$ iteration
$d_{\mathcal{C}^{2}}(f^{n}(\Gamma),\Lambda)\leq C(\lambda_{+}\mu_{-}^{2})^{n},\quad n>0.$
Repeating the tangent functor trick yields the desired conclusion.∎
Remark C.3.
Assume that $\mu_{+},\mu_{-}$ satisfy (N) rather than (C.1). Then, we can obtain the same results as in (C.2) and (C.3), for $|n|$ sufficiently large, provided that either $\lambda_{\pm}$ or $\mu_{\pm}$ are chosen not to be the optimal rates, that is,
(C.10) $\lambda^{*}_{\pm}<\lambda_{\pm}\textrm{ or }\mu_{\pm}^{*}<\mu_{\pm}.$
This is because we can choose $\lambda^{*}_{\pm}<\tilde{\lambda}_{\pm}<\lambda_{\pm}$ or $\mu_{\pm}^{*}<\tilde{\mu}_{\pm}<\mu_{\pm}$ and run the argument in the proof of Lemma C.1 for the rates $\tilde{\lambda}_{\pm}$, $\tilde{\mu}_{\pm}$, except for the last inequality in (C.9), when we use $n(\tilde{\lambda}_{\pm}\tilde{\mu}_{\mp})^{n}\leq C(\lambda_{\pm}\mu_{\mp})^{n}$ for $|n|$ sufficiently large. Without the extra condition (C.10), instead of (C.2) and (C.3) we obtain we obtain $d_{\mathcal{C}^{1}}(f^{n}(\Gamma),\Lambda)\leq Cn(\lambda_{+}\mu_{-})^{n},\quad n>0$, and $d_{\mathcal{C}^{1}}(f^{n}(\Gamma),\Lambda)\leq Cn(\lambda_{-}\mu_{+})^{n},\quad n<0$, respectively.
Similar statements hold for (C.4) and (C.5).
Appendix D Extensions to other models
The machinery developed in this paper is rather robust and produces similar results for other models.
In this Appendix, we discuss two models (see sections D.1D.2) of physical interest that have appeared in the literature. We show in Section D.3 that the vanishing lemmas apply to these models, and in Section D.4 that the scattering map also preserves the corresponding forms.
The methods of this paper apply to these cases, and we hope they could lead to results that complement the ones presented here.
D.1. Partially conformally symplectic systems
We consider products of symplectic manifolds
$(M,\omega)=(M_{1}\times M_{2}\times\cdots M_{L},\omega_{1}\oplus\omega_{2}\oplus\cdots\oplus\omega_{L})$
where $\omega_{i}$ is a symplectic form on $M_{i}$, $i=1,\ldots,L$.
We consider maps $f$ on $M$ such that
(D.1) $f^{*}\omega=\eta_{1}\omega_{1}\oplus\eta_{2}\omega_{2}\oplus\cdots\oplus\eta_{L}\omega_{L},$
for $\eta_{i}>0$, $i=1,\ldots,L$. Such a map is not conformally symplectic (see Definition 2.10).
Systems of the form (D.1) appear in mechanics when we consider models of $L$ particles interacting by a Hamiltonian, where each particle is subject to a friction proportional to its velocity. See [CallejaCL13, Remark 3].
Such models have been explored in the literature for various applications. Below, we provide only a few examples.
  • The spin-spin model [^19] The spin-spin model is a time dependent flow. The time advance maps are of the form (D.1). describing two rotating bodies, each with its own tidal friction [Misquero21, eq. (42)].
  • Networks of oscillators with Hamiltonian coupling among nearest neighbors, where one node experiences dissipative effects [EckmannW20, EckmanW18, CuneoEW17, DolgopyatFKY2025]. These are models of the form (D.1) in which $\eta_{i}=1\textrm{ for }i\neq 1$, $\eta_{1}<1$. There are invariant manifolds that affect the transfer of energy from the conservative modes to the dissipative modes.
  • Models of planets with a two-layer structure subject to viscous and tidal friction, each of them with different friction coefficients. See [PinzariSV24].
Formally, several Hamiltonian PDEs subject to dissipation are of the form (D.1). For example, consider the telegraph equation
$u_{tt}+u_{t}-u_{xx}=0$
with periodic boundary conditions. Writing $u(t,x)=\sum_{k}\hat{u}_{k}(t)e^{2\pi\imath kx}$, we obtain formally
$\hat{u}_{k}^{\prime\prime}(t)+\hat{u}_{k}^{\prime}(t)+k^{2}\hat{u}_{k}(t)=0,$
which is a system of uncoupled conformally symplectic oscillators with uniform dissipation. Other models with strong dissipation $\frac{d}{dt}u_{xx}$ instead of $u_{t}$ have also been considered. Similar considerations apply to other Hamiltonian Partial Differential equations and infinite dimensional coupled systems. For example, [EckmannW20, EckmanW18, CuneoEW17] explore analogies with locally dissipative variants of the nonlinear Schrödinger equation.
Making rigorous sense of geometric properties of NHIMs in PDEs seems an interesting problem, but it is tractable for finite NHIMs [BatesLZ08].
One should also note that in infinite dimensional models, there may be dissipation of energy even in Hamiltonian systems [CorsiG22, ComechKK23].
D.2. The Gaussian thermostat
Several models of non-equilibrium thermodynamics are based on introducing some forcing as well as some dissipation to keep energy constant. See [wojtkowski1998conformally].
These systems have several extra structures, but they lead to (see [wojtkowski1998conformally, equation (2.4)]) maps that satisfy:
(D.2) $(f^{*}\omega)(x)=\eta(x)\omega(x)$
where $\omega$ is a non-closed form (which allows for $\eta$ to be not constant), and $\eta(x)$ is a bounded function, whose inverse is also bounded.
The form $\omega$ and the factor $\eta(x)$ used in [wojtkowski1998conformally] satisfy several other properties that lead to other consequences, but they will not be used below.
Condition (D.2) appears in this paper in the consideration of presymplectic maps (see (2.10)) with rank smaller than $4$.
D.3. Vanishing lemmas for the generalized models
Both the models in (D.1) and models of the form (D.2) are particular cases of maps that satisfy
(D.3) $\begin{split}&|\omega(x)(u,v)|\leq C\eta_{-}^{-n}\|\omega(f^{n}(x))\|\|Df^{n}(x)u\|\|Df^{n}(x)v\|,\quad n\geq 0,\\ &|\omega(x)(u,v)|\leq C\eta_{+}^{-n}\|\omega(f^{n}(x))\|\|Df^{n}(x)u\|,\|Df^{n}(x)v\|\quad n\leq 0.\end{split}$
Instead of (6.1). Observe that these inequalities are similar to the ones obtained in the presymplectic case (10.1).
In the case of the models (D.1) we take $\eta_{-}=\min(\eta_{1},\ldots,\eta_{L})$, $\eta_{+}=\max(\eta_{1},\ldots,\eta_{L})$.
In the case of the model (D.2), which corresponds to the presymplectic case studied in 2.10, we can take [^20] Since $\eta_{\pm}$ are just bounds, one can use sharper bounds to improve the conditions of theorems. Some simple improved bounds used in [wojtkowski1998conformally] are, for any $K$: $\eta_{-}=\inf_{x}\left(|\eta(x)\eta(f(x))\cdots\eta(f^{K}(x))|\right)^{1/(K+1)}$, $\eta_{+}=\sup_{x}\left(|\eta(x)\eta(f^{-1}(x))\cdots\eta(f^{-L}(x))|\right)^{1/(K+1)}$. $\eta_{+}=\sup_{x}|\eta(x)|$, $\eta_{-}=\inf_{x}|\eta(x)|$.
For these models, the proofs of the vanishing lemmas remain valid under hypotheses that the rates of vectors and the numbers $\eta_{\pm}$ are appropriately related.
For example, if for $x\in\Lambda$, $u\in T_{x}M$ and $v\in E^{s}_{x}$, and for all $n\geq 0$ we have
$\begin{split}&\|Df^{n}(x)u\|\leq C\mu_{+}^{n}\|u\|,\\ &\|Df^{n}(x)v\|\leq C\lambda_{+}^{n}\|v\|,\textrm{ with }\\ &\lambda_{+}\mu_{+}\eta_{-}^{-1}<1,\end{split}$
then we conclude , by the vanishing Lemma 6.8, that $\omega(u,v)=0$.
Observe that this condition on rates also appears in the statements of Theorem 3.9. See (S’).
In conclusion, the proofs of the vanishing lemmas can be adapted without change. Unfortunately, the fact that it could happen that $\eta_{+}\neq\eta_{-}$ prevents a proof of the pairing rules.
D.4. Geometric properties of the scattering map for the generalized models
Many of the proofs of part (B) generalize to these cases (as noted before, the statement that $S$ is symplectic, when $\omega$ is not a symplectic form should be understood to mean that $S$ preserves $\omega$).
For both models in Section D.1 and  D.2 we can use the proofs of simplecticity of scattering map in Sections 8.2.5, 8.2.7, using the appropriate assumptions on rates. As remarked in the text, these proofs do not require that $\omega$ is closed or non-degenerate. So, they can be used modulo changing the assumptions [^21] As it turns out, the proof in Section 8.2.6 can also be adapted. Taking $d$ of (D.2), we obtain that $f^{*}d\omega=\eta d\omega+d\eta\wedge\omega$. The geometric setup of [wojtkowski1998conformally] implies that $d\eta\wedge\omega=0$. on the rates to include $\eta_{\pm}$ instead of $\eta$.
For the models in Section  D.1 $\omega$ is a symplectic form. Hence we can use without change (except in the assumptions on the rates) the proofs in Sections 8.2.1, 8.2.3, which use that $\omega$ is closed.